A Complete Bibliography of Publications in IEEE Software: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.01

Title word cross-reference

3 [Ano14z, Ano15y]. 4 [FDE^{+}17]. N [MHM19].

-Gram [MHM19].

/Should [HR11].

10 [TS11]. 10-Minute [Whi12]. 15 [Ano15-35]. 178C [MLD^{+}13].

2.0 [GAM^{+}17, ST10]. 2010 [Shu11f]. 2013 [ZK14]. 2015 [Ano15-32]. 2016 [Ano15b]. 2019 [Ano18g, Ano18a, Ano18b, Ano19r, Ano19-54, CM19]. 21st [Ano14-32, MPD14, PRRT14].

3.0 [ABB^{+}18].

40 [Gre17a]. 40th [Ano17b, Ano18c]. 41st [Ano18d].

5 [FSM14].

9 [Bla18a].

Abstractions [Zam17]. Abundance [Spi17b]. Academia [CP18, SPA11, WAA^{+}12]. Academic [CH15c, GF17, MSSMDC14]. Access [ASB^{+}12, GP13, KSW10, TT11]. Accessible [Ano17c]. Accomplishments [TD19]. Accuracy [LCYL14]. Accurate [FS12]. Achieve [Ano16c, Ano17e, Ano17d]. Achieving [FSM14, HBCH14, ZXT^{+}15].

ACM [Ano17f, Ano19c]. ACM/IEEE [Ano19c]. Acquisition [BKB12]. across [PMCE12]. Act [LTS12]. Action [RCP^{+}12, SPA11]. Actionable [GAM^{+}17, PT18, YFMH18]. Actor [Til16b]. Ad [RNA^{+}14, RNA^{+}16a]. AdaCore [Bro19]. Adages [PHA^{+}17]. ADAPT [VSB^{+}16]. Adaptable [Ano17m, LX17]. Adaptation [AGK^{+}19, BCLP15, PWF^{+}19, RRL^{+}17, WG10]. Adapting [GFT15]. Adaptive [BBB^{+}19, CGE19, HLS17, NPJ19]. Adding [KC16]. Adopt [MFS15]. Adopters [ES16]. Adopting [SF15]. Adoption [ACHC11, SSS15a]. Advancement [Sca19]. Advances [SWB^{+}17, EHS15]. Advancing [ABH^{+}11]. Adversaries [CHDK^{+}16]. Advertisement [Ano14d, Ano14j, Ano14l, Ano14n, Ano14x, Ano14-31, Ano14-32, Ano15e, Ano15f, Ano15n, Ano15o, Ano15q, Ano15r, Ano15b, Ano15z, Ano15-27, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15-35, Ano15-36, Ano16c, Ano16g, Ano16n, Ano16t, Ano16u, Ano16r, Ano16y, Ano16-37, Ano16-39, Ano16-41, Ano16-35, Ano16-44, Ano16-43, Ano16-50, Ano17e, Ano17l, Ano17q, Ano17x, Ano17-38, Ano17-41, Ano17-44, Ano14c, Ano14o, Ano14k, Ano14m, Ano14u, Ano14v, Ano14w, Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14z, Ano15g, Ano15p, Ano15y]. Advice [GJO^{+}18]. Aesthetics [Mai11b]. Affects [LL16, VGD19]. After [TD19, RvGH17]. Again [BDH^{+}19, Vol11]. against [BFKZ15]. Age [Ano16-42, Bre14, HOZZ16, Ozk19c]. Agent [CCS^{+}11]. Agents [Pec13]. Agile [ADH14, BHN14, BH19, Bav12, BGK15, BMS^{+}14, CE19, CdAC^{+}16, CPP10, CH17, CCWP11, CC19, DFP19, DFD^{+}16, EP17a, ES16, EP16, FCS^{+}10, FRM15, FP13, GGGM19, HBS19, HSG18, HM13, Jor19, Kna19, Mad10, MJ10, MCL12, Mey18, MRS10, PL19, Poo14, PLTC16, PLC18, PLC19, Ric11, RH13, SPA11, SMN19, Shu13a, Sma18, SMLF19, Stu18, SS18, VSB^{+}16, Wat18a, Wat18b, Woo15a, ZP11, SC18]. Agile-Methodology [SS18]. Agility [ABK10, BH13, PLC19, RBE19, Sma18, Spi11a, Wat18a, Wat18b]. Agriculture [SJ15]. Ahead [CHDK^{+}16, JPM^{+}18, KAC^{+}18, LMM^{+}15, Shu11e]. Air [CFF^{+}13, Lou10, VOGB18]. Airbus [BHH13]. Ajax [DS13]. Ajax-Based [DS13]. Alexander [Blu16]. Algorithm [RTV12]. Algorithms [PZTF12]. Aligning [BB17, Woo15a]. Alive [CH13b]. ALM [Gat16]. Alternatives [MLD^{+}13]. Ambient [GVSP^{+}17]. Ambiguity [Mai12a]. Amisoft [RVB13]. Analogy [Ano17o]. Analysis [CPS18, CMS11, DS13, GTS10, Jor19, KCT12, KTH11, KSTS17, MW11, MTM^{+}19, MHLCL16, RPP19, ZKB^{+}16]. Analyst [Mai10b, Mai11a]. Analytics [Ano14-27, Ano16-40, AHMS18, BHG13, CNSM13, HHR^{+}13, IGBKR19, Joh13, LE13, MZ13a, MZ13b, MZ18, Men18, MCBT13, PT18, RVB13, SC18, SCB15, YFMH18, ZHD^{+}13]. Analyzing [PCP14, RNA^{+}16a, VMVS19]. Anarchy [Boo15b]. Android [RNA^{+}14, RNA^{+}16a, ZKB^{+}16]. Anecdotal [HBCH14]. Angular [Far17]. AngularJS [RVT18]. Ann [Ano11b]. Annals [Ano19d, Ano19e]. Annotation [FS12]. Annual [vGH12]. Answering [SB16]. Anticipation [Poo16]. Antipatterns [KUG19]. Anxiety [Gau13]. Apache [Mey16]. Apart [GF17]. API [LXB^{+}15, UR15]. APIs [Spi12a]. App [BCN16, CM19, KSNH15, KNH16, PSO12, RNA^{+}16b, SS17, VMvG12, Sad14, GAM^{+}17]. Appealing [NHM19]. Applicability [Scr18]. Application [BKB12, CSS12, GKD^{+}15, LL16, NS17, SSE12, VAC^{+}17, ZML^{+}12]. Application-Screen [GKD^{+}15]. Applications [ABB^{+}15b, BW17, DSM^{+}18, DS13, GRD11, GE11, GD12, Gho10, GP13, Gun10, KAC^{+}18, LXC15, Lou10, Lou16, RPP19, Ano19f]. Apply [TD12]. Applying [LMMO18]. Appreciating [BHN^{+}10]. Approach [BW14, CFF^{+}13, CH13c, Dig11, FSB^{+}12, GKD^{+}15, HJP16, KSW10, LNH^{+}11, MZLD12, SCB15, TEvdHP12, WLD19]. Approaches [KDM^{+}19]. Approximating [CM12]. Apps [AFT^{+}15, MHJ17, MSAH17, MAN^{+}14, RNA^{+}14, RNA^{+}16a, SHG13, ZCLL19]. Archaeology [UG18]. Architect [Ano16-42, Bas18, Boo11a, Boo12d, Bus11c, Bus12b, CH14a, EP16, HOZZ16, Kee15a, Kle16, Kri17, Poo16, Ric11, Spi16b, TKF16, WG16, Woo14b, WF18]. Architecting [MGMS11, Poo14, RL16, Woo15b]. Architects [AMK16, BSD16, Cle10, Cor13, EP17b, Fab10, FCN16, Ris12a, Woo17, Wri11]. Architectural [AACF13, BW17, Boo10a, Fai14, KHYT18, LMM^{+}15, Lat12, Roz15, SM10a, Woo14a, Woo16a, Woo16b, ZCTZ13, Zim11, Zim15]. Architecturally [CBN13, CH13c, CH13d]. Architecture [ABK10, ASB^{+}12, ASH13, BHJ16, BGK15, BWC10, Boo10c, Boo10b, Boo10e, Boo10f, Boo11b, BMSF15, Bus10c, BH10a, BH10b, Bus11a, Bus11b, BAA^{+}12, BH13, CM10, CNZC17, CVEK13, CHHSM13, Cra14, CMC^{+}12, DSM^{+}18, FCS^{+}10, GK15a, HA11, Joh15, Kee15a, Kee15b, KSW10, Ker17, KDGV13, LAH10, LDO^{+}14, Mad10, MH14, NOS12, PvdHM15, PTV^{+}10, PCRPE19, Pra17, SR16, SHG14, Shu13a, Spi15a, TVS10, TMB17, Vol10, Wat18a, Wat18b, WL12, Woo15a, Woo15b, Woo16c, vHEA^{+}14, Sun19]. Architecture-Conformance [PTV^{+}10]. Architecture-Driven [CM10]. Architectures [BvD10, CCC^{+}19, LR12, SA19, Spi18d, Stu18, TM18, WE16, WCG16]. Areas [GF17]. Arguing [GHH^{+}12]. Art [MIC19, Spi17f]. Articles [Ano19o, ZK14].

Artifacts [ZZJJ15]. Artificial [BDH^{+}19]. Aspects [Boo14b, Haz10]. Assembly [SV16]. Assertive [Hol15a]. Assessing [APPV19, CSS13]. Assessment [FN14, HBS19, ILM11, ZF13]. Assets [Zim11]. Assignment [dSAPW^{+}14]. Assisted [dSAPW^{+}14, GVSP^{+}17]. Assurance [CPP16, RLBR12]. Assuring [Shu11a, Shu12c]. Asymmetry [Hol19a]. Asynchronous [Dig15]. ATAM [BGK15]. Athletic [HJP16]. Attacks [BFKZ15]. Attribute [KHYT18]. Attributes [HL11]. Authors [RBE19, ZPHW16]. Autobiography [Boo15d]. Automated [AFT^{+}15, AIT15, BH17, CS18, GF16, GTS^{+}19, LCYL14, MHM19, SAO^{+}17, ZML^{+}12]. Automatic [YLK^{+}17]. Automation [GE17, MHJ17, Ozk19a, Pec13, PRPE13]. Automotive [AS17, dSAPW^{+}14, EF17, FLC17, HOBK17, LWCP17, LDBF19, Mos10, PST^{+}17, SSE12, TMB17, Tsa11]. Autonomic [SEK^{+}19]. Autonomous [EW19]. Availability [Ron16]. Aviation [GTS^{+}19]. Award [Ano16e, Ano16p, Ano16r, Ano16q, Ano16b, Ano17i, Ano18f, Ano19c, Ano15b, Ano17f, Ano17x, Ano18a, Ano18b, Ano19c]. Awards [Ano16d, Ano17h, Ano18-37, Ano18-38]. Aware [Ano17m, BW17, CPB^{+}17, FKL19, MCHK17, MTT^{+}15, SWB^{+}17, WLD19]. Awareness [CYM^{+}15, LCE13]. Away [Bro19].

Babbage [Ano18a, Ano18b]. Back

[SDSP17, Vol11]. Backouts [SCB15]. Bad [BZd18, TL18, vGH19]. Badges [NHM19]. Balancing [LTS12]. Balls [Hol18b]. Bank [LTv19]. Banking [BDD^{+}18, CPP10]. Barriers [ZSMP17]. Barry [EM18, Tho15a]. Baseball [Val17]. Based [AGK^{+}19, ABB^{+}15b, AFT^{+}15, Ano16-41, ADH14, ABC^{+}11, BBB^{+}11, BKB12, BHS^{+}14, BDD^{+}14, BO17, BW14, DHB15, DS13, DPCL14, FSB^{+}12, FFGM13, GD11, GVSP^{+}17, HA11, Hof10, KCT12, KZW15, KCD12, LX17, LNH^{+}11, MFS15, MHF17, NS17, PL19, PvdHM15, PS14, RRL^{+}17, Sch12, Til15, TEvdHP12, ZZJJ15]. Baton [Shu14e]. Bayesian [FN14]. Be [Ano16t, Ano16u, Ano16s, Ano18t, Ano18u, BWD^{+}17, Erd10a, HR19b, HR11, BS15, PPMD17]. Beach [Shu11b]. Become [Ano17g]. Before [Bus12a]. Behavior [Ano16-32, RRL^{+}17, dSMdSC16]. Behavior-Driven [Ano16-32]. behind [Rou12, WK15]. Being [SvGH15, Spi13c, Spi16a, Spi18a]. Belief [DZB18]. Believe [Ano12b, Shu12d]. Bells [DeM13]. Ben [Blu19, Mey16]. Benefit [HBS17a, HBS19, Ris10a]. Benefits [Che15, HTE16]. Bespoke [Spi14a]. Best [BCD^{+}17, Bus11c, DFFP18, GLA14, HBS17a, LCA17, MGS^{+}18a, PPMD17, Shu14d, Ano18v, Ano19-35]. BestBuy.com [Cra14]. Better [Ano16-33, Ano16-34, Ano17-38, Ano17-35, Ano17-36, Ano17-37, EHMFT19, Fai19a, IML16, Jor16, Ano17j, Ano17-34]. Between [Jor19, MAGT19, EJ12, GV11, KNH16]. Beyond [BBM^{+}15, HBCH14, Men13, CSS11]. Big [Ano14-27, Ano16x, Ano16-37, Ano16-36, Ano19k, Ano19l, Ano19p, Ano19q, Boo14a, Boo14b, CKH16, GK15a, GBM16, LE13, MHLCL16, Shu13c, SPB16, WZX^{+}16, HTE16]. BIP [BBB^{+}11]. Birth [GO15]. Bit [Boo15e]. Black [MHJ17]. Black-Box [MHJ17]. Blind [Boo10d, Car19]. Bloat [MSS10]. Blockchain [KV18, LX17, Scr18]. Blockchain-Based [LX17]. Blockchain-Enabled [KV18]. Blocks [KV11]. BlueGriffon [DeM14]. Blueprint [RMM^{+}12]. Board [Tsa11, Ano19-31, Ano19-32, Ano19-33]. Bob [Ano10b]. Boccara [Bel19]. Boehm [EM18]. Boilers [Boo15e]. Boldly [Bus12a]. Books [SCC11]. Boring [Boo13d]. Boson [Rou12]. Bots [Boo15e, LSZ18]. Bottlenecks [LDBF19]. Bought [LC11]. Bound [RTV12]. Bounds [Hol15d]. Box [MHJ17]. BPMN [Ano16-32, LMG16]. Brace [Hol16a]. Bracha [TW14a]. Brand [Mai10a]. Brave [Shu12a]. Brazil [FWK^{+}15, dOMFP16]. Break [RK13]. Breezing [Kri17]. Brian [WCFG10]. Bridge [Ano14-41]. Bridging [BBS13a, GJO^{+}18]. Bringing [Cap14, Obr13]. Brock [Ano10b]. Broken [Jun18]. Budzier [Blu16]. Bug [NHM19, RPP19, ZCLL19]. Build [BMDK15, MM18b]. Building [CGE19, CNSM13, KV11, NPJ19, SCW^{+}18, WZX^{+}16]. Bullet [Ozk19a]. Bullets [Mar12]. Burning [FS10]. Burnout [MAGT19]. Business [ABC^{+}11, BKB12, Bro19, CB10, EJ12, FP11, HBS17b, LRP11, LKR16, MCH15, MZLD12, ORM^{+}17, OT11, Pop11, Rie11, Ris12a, RBE19, RSH12, SR16, Spi16d]. Bütow [Sal18]. Buzz [PB16]. Bytecode [SEK^{+}19].

C [HR19a, YLK^{+}17]. Cabin [BHH13].

Calculating [CM10]. Call

[Ano14b, Ano14-34, Ano15e, Ano15c, Ano15d, Ano15-32, Ano16d, Ano16n, Ano16p, Ano16r, Ano16q, Ano17h, Ano17i, Ano18f, Ano19c, Ano19o]. Calls [Ano14i]. Cambrian [Ker18a]. Can [ABK10, Ano17j, APPV19, Erd10a, HBB^{+}11, Ric11, WLD19, vGMRW14]. Candidates [CSS13]. Can’t [Ris12c]. Capabilities [EBS16]. Capability [vWHZ^{+}10]. Capability/Maturity [vWHZ^{+}10]. Capital [Sma18]. CAPTCHA [KC16]. Capture [CMS11]. Capturing [TEvdHP12]. Car [HCF^{+}15, SA11]. Career [Ano15q, Ano15r, Ano16w, Ano17y, Ano16c, Ano17e, Ano17d]. Careful [Par18]. Cares [DHH^{+}19]. Caring [LLV16]. Case [AtHR11, BCLP15, BBN^{+}17, BSD16, BMSF15, ESZ15, Kee15a, LX17, McG11, NHM19, NCK11, OT11, RVB13, RSH12, SJS12, WKS^{+}14, ZF13]. Caseload [Lob11]. Cases [Ano16-32, CdAC^{+}16]. Categorization [SS17]. Causal [KCT12]. Cause [MTM^{+}19]. Cause-and-Effect [MTM^{+}19]. Causes [MTM^{+}19]. Center [Ano16t, Ano16u, Ano16s, Ano18u, BWN16, Ano18t]. Central [BBP^{+}16]. Centric [EP16, SM10a, WHR10, Zim15, vHEA^{+}14, GMB^{+}14]. Century [MGS18b, Ano14-32, MPD14, PRRT14]. Certification [Kru10]. Certified [Cle10]. CG&A [Ano14i]. Chaiken [Fav12]. Challenges [Che15, CCWP11, CC19, GSA^{+}17, JPM^{+}18, LvdH16, MGS^{+}18a, MG17, MHLCL16, RSS18, SA19, SCL18, SSS15a, Shu11b, Spi18b, SWA17, TM17]. Change [EEBB11, Gre19, ILM11, Shu13e, WHR10]. Change-Centric [WHR10]. Changes [Spi17b]. Changing [AS17, Spi16b, Woo16c]. Chaos [BBdR^{+}16, EV10, Sal18]. Characteristics [VB16]. Characterizes [FWS19]. Characterizing [CBN13]. Charles [Ano18a, Ano18b]. Chart [BCLP15]. Chatbots [PSGd18]. Cheap [Spi12e]. Check [BvD10, PZA^{+}17a]. Checking [BMSF15, KCK^{+}13, PTV^{+}10]. Checks [Gim14]. Cherishing [Mai12a]. Chilean [ORM^{+}17]. Chinese [TYXL16]. Choice [VB16, ZHG14]. Choose [Ano16w, Ano17y]. Choosing [Eng10, Spi11b]. Choreographies [AIT15]. Chrome [RR15]. Chu [Hen19]. CiSE [Ano17k]. City [CPB^{+}17]. Clarity [BJ11, Hol16b]. Class [PCP14]. Classification [BCLP15, MHM19]. Classify [FC19]. Clear [BJ11]. Clearing [JN10]. Click [BBC^{+}19]. Client [KR10, TM18]. Clients [Rie11]. Climate [BJ11, CR11, CMC^{+}12, EEBB11, HR11, ILM11, RDMA11, Shu11a]. Climate-Modeling [CMC^{+}12]. Clock [Shu14a]. Clone [MOMM11]. Cloud [Ano14j, Ano14k, Ano14-30, Ano14-31, Ano16n, BHJ16, Cra14, DSM^{+}18, EP16, HCF^{+}15, Lou10, LXB^{+}15, MKD14, MOK^{+}18, OT11, RMM^{+}12, SGH15, Spi14b, TS12, Til15, ZXT^{+}15, ASB^{+}12, GKKL12, RKTS12, ZML^{+}12]. Cloud-Based [Til15]. Cloud-Centric [EP16]. Cloud-Native [BHJ16]. Cloudberry [TS12]. Clusters [RK13]. CMMI [FSM14, MSSMDC14]. CMU [MGS18b]. Code [APPV19, BH17, BBdH^{+}11, BJ11, Bel19, Boo14f, BE11a, Bro19, BB12, CCH^{+}15, CCCM17, EM15, Fai19a, Fai19b, FFZ^{+}18, GKF19, Hol14, Hol15a, Hol15b, Hol15f, Hol15c, Hol16b, Hol17a, Hol18a, Hol19b, Hol19c, Joh17, MGS^{+}18a, Mor10, Rie11, Spi10a, Spi12a, Woo17]. CodeBender [CT11]. CODEMINE [CNSM13]. Coderetreats [PMSL14]. Coexist [ABK10]. Coexistence [FCS^{+}10]. Cold [Boo16c]. Collaborate [Ozk19e]. Collaboration [CE19, CP18, DvGvS14, LEPV10, PMCE12, RI12, SPA11, WAA^{+}12]. Collaborations [CCH^{+}15]. Collaborative [BBM18, DBRV18, DDIP15, DD12, JWCR18, MBv18, PBE14, PSGd18, SPA11]. Collaborative-Design [BBM18]. Collective [GMB^{+}14]. Columns [vGH19]. Combination [Ano17k]. Combining [LK10]. Come [DFFP18]. Comet [BBP^{+}16, Bal18]. Coming [Ker18b]. Commits [CCMT16]. Common [Gre18a, Gre18b, RO15]. Common-Enough [Gre18b]. Communicating [SR16]. Communication [BMS^{+}14, LA19, WSE14]. Communications [NH15]. Communications-Critical [NH15]. Communities [BBS13a, GBICMR13, TLvV13]. Community [CMC^{+}12, MIC19, PL19, SB16, TKF16, ZWGY12]. Community-Based [PL19]. Companie [EBS16]. Companies [GBICMR13, ORM^{+}17, RVB13, RBE19]. Compare [Jor13]. Comparing [FWK^{+}15, LA19]. Comparison [SJ13, SS17]. Compass [FRM15]. Competence [KHD19]. Competing [LDBF19]. Competitor [LC11]. Compilers [HAE11]. Complain [KSNH15]. Complex [MTM^{+}19, SCW^{+}18, VMVS19]. Complexity [DS13, EC16, Hin19, RDMA11]. Compliance [FSB^{+}12, GD12, Shu12c, TEvdHP12, ZBOC12]. Complicated [Hin19]. Component [BBB^{+}11, Lou16]. Component-Based [BBB^{+}11]. Components [CSS11, GWLY11, HKR11, Hol18c, SJ13, Spi11b]. Composing [KR10]. Compositional [Bos10]. Compound [vGH12]. Comprehensive [RO15, ZKB^{+}16]. COMPSAC [Ano18g, Ano19r]. Computational [Boo16a, Hv19]. Compute [Hol19d]. Computer [Ano14l, Ano16e, Ano16t, Ano16u, Ano16x, Ano16p, Ano16v, Ano16r, Ano16q, Ano16s, Ano16w, Ano16o, Ano17y, Ano17x, Ano18p, Ano18q, Ano18t, Ano18r, Ano18s, Ano18u, Ano19f, Ano19y, Ano19z, Ano19-27, Ano19-28, Ano19a, Ano19b, Ano19-29, Ano19-30, BBP^{+}16, Boo15c, GP13, Gla12, Sca17]. Computer-Managed [GP13]. Computing [ASB^{+}12, Ano14k, Ano18-37, Ano18-38, Ano19g, BBM^{+}15, Boo15b, EAG^{+}14, FTP11, Gla12, LAAN10, LRB^{+}15, LNH^{+}11, MOK^{+}18, PSO12, Ano14j, Ano16n, Ano16y, Ano18e, Ano19d, Ano19e, Ano19m, Ano19n]. Concentration [KE10]. Concept [Che14]. Concern [Ozk19b]. Concerns [Che14, GHV10, LAH10]. Conclusions [VD15]. Conference [Ano17b, Ano17z, Ano18c, Ano18d]. Conferences [Ano15f, Ano17l, Ano18h, Kee15b]. Conflict [SM10a]. Conflict-Centric [SM10a]. Conflicts [BBM18]. Conformance [GHH^{+}12, PTV^{+}10]. Congress [Ano19-54].

Connect [Ano17-44, Ano17-42, Ano17-43, Ano18i, Ano18-36]. Connected [FCN16, HCF^{+}15, Ano16-44]. Connection [MAGT19]. Connector [PZ18]. Conservation [CdAC^{+}16]. Considerations [BH10a, BH10b]. Consistency [FFZ^{+}18]. Constant [Shu13e]. Constraint [Hof10]. Constraint-Based [Hof10]. Construction [ZZJJ15]. Consumption [BFKZ15, CM19, KHYT18, PHAH16, PCP14, ZHG14]. Contemporary [RCP^{+}12]. Content [Cab18, RMM^{+}12]. Contents [Ano14-36, Ano14-37, Ano14-38, Ano16-46, Ano16-47, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano18-43, Ano18-44, Ano19-48, Ano19-49, Ano19-50, Ano19-51, Ano19-52, Ano19-53, Ano14-39, Ano14-40, Ano15-33, Ano15-34, Ano16-45, Ano16-48, Ano17-45, Ano18-39, Ano18-40, Ano18-41, Ano18-42]. Context [Ano17m, BBN^{+}17, CPB^{+}17, MCHK17, MTT^{+}15, SWB^{+}17, Tor15, Woo14a]. Context-Aware [Ano17m, CPB^{+}17, MCHK17, MTT^{+}15, SWB^{+}17]. Context-Driven [BBN^{+}17]. Contextual [TAJB17]. Contextual-Variability [TAJB17]. Contextualizing [Dyb13]. Continuous [Che15, LMP^{+}15, Mey14a, NBF^{+}18, PHA^{+}17, SCL18, SCW^{+}18]. Continuously [HSEG10]. Contract [DHB15, Fai19a]. Contract-Based [DHB15]. Contribute [MSSMDC14]. Control [ASB^{+}12, CFF^{+}13, Fai19d, GP13, Jep13, PGG^{+}18, TT11, TI11]. Controllers [Tsa11]. Convergence [ED19, GK15a]. Conversation [EM18, RBE19, ZPHW16]. Conversations [Coo10]. Conway [KCD12]. Coordinating [MOK^{+}18]. Copier [TS11]. Coping [DHB15]. Copyrights [GD11]. Core [Ano15-36, Ano16-50, Ano17-52]. Corporate [PTNF^{+}19]. Correct [MM18b]. Corresponding [LRP11]. Cost [Poo14, Shu14f, SCW15, SvS16, UM10, WLD19]. Cost-Aware [WLD19]. Costs [BBM18].

COTS [GRDL^{+}12, Lob11]. Could [MSSMDC14]. Count [CM12]. Count-Min [CM12]. Counterpoint [GE12, GWP^{+}11, HR11, KvdL10, VE12, WK11, WES10]. Country [LMP^{+}15]. Coupled [DIP12, RDMA11]. Coupling [KE10]. Cover [Ano14e, Ano14g, Ano15i, Ano15k, Ano15l, Ano15m, Ano16h, Ano16i, Ano16k, Ano16l, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano14f, Ano14h, Ano15h, Ano15j, Ano16j, Ano16m, Ano17r]. Coverage [Als13, ADSS18]. Covers [Spi18f]. CPUs [SRI14]. Craft [Hol17a]. Crash [CCMT16]. Crash-Inducing [CCMT16]. Cray [Ano18-37, Ano18-38]. Creating [GFT15, Gor13, ZB16, vWHZ^{+}10]. Creative [Coo10]. Creativity [HM13]. Credit [MH14, SJG13]. Crisis [Gre17a]. Critical [CCP^{+}18, CCD^{+}17, KSTS17, LCF13, MJZCH13, MG17, NH15, PT18, Sca19]. Crosscutting [Che14]. Crosses [HvG16]. Crowd [ASR17, GVSP^{+}17, GSA^{+}17, LBL17, MRF^{+}13]. Crowd-Based [GVSP^{+}17]. Crowdsourced [GAM^{+}17]. Crowdsourcing [BH17, LvdH16, PBE14, SLB17, ZMP^{+}16, ZSMP17]. CrowdSummarizer [BH17]. CS [Ano17f, Ano18a, Ano18b, Ano19c]. CSI [KB12]. Culture [PDNdSB12]. Current [SCL18]. Curricula [Haz10]. Curriculum [FP13]. Curve [Hol18b]. Customer [CPP16, HBS17b, WLD19]. Customers [HMPS16]. Customization [MS14]. Customs [Fre10]. Cyber [Ano16-38, BMSS18, Gor13, RCLR18]. Cyber-physical [BMSS18, RCLR18]. Cybersecurity [Ano14-28, Ano14-29, Ano15z, Ano15-27, Ano16-39]. Cycle [Fav10, LIL^{+}13, SJG13]. Cyclomatic [EC16].

D [Ano17x, Ano14z, Ano15y, FDE^{+}17]. Dafny [Ano17c]. Dark [LML^{+}15]. Darkitecture [Pra17]. Dashboards [BHG13]. Data [Ano14-27, Ano16x, Ano16-37, Ano16-36, Ano18g, BWN16, BGK15, Boo14b, Bos16, CPP16, CKH16, CM12, CNSM13, EHMFT19, FFZ^{+}18, Gho10, GK15a, GBM16, Li14, LKR16, LE13, MNJR16, MCH15, Men13, MHLCL16, Shu13c, Shu14b, SRI14, SPB16, VD15, Via19, WZX^{+}16, HTE16, Ano19k, Ano19l, Ano19p, Ano19q]. Data-Driven [MNJR16]. Database [Hen19, Via15, Sun19]. Databases [Via18]. Dataflow [Che14]. Dave [Joh16]. David [Ano13b, Fav12]. Dead [Hol17b, Spi15c]. Dealing [HR19a]. Dear [LC11]. Death [DLR10]. Debate [Ozk19e]. Debt [Bav12, Bus11d, CCCM17, Con12, CSS12, KNO12, LI12, LTS12, RKW15, SWA17, WJ15]. Debug [Spi16c]. Debugging [Ano13c, ESZ15, Spi13a]. Decade [ZPHW16]. Decades [Ker18b]. Decision [AACF13, FN14, PL19, PSGd18, WG16, vHEA^{+}14]. Decision-Centric [vHEA^{+}14]. Decisions [CH15a, HvG12, ZCTZ13, Zim11]. Declarative [Spi13c]. Deep [Car19, FLC17, ZXL^{+}16]. Deep-Intelligence [ZXL^{+}16]. Deepening [HBCH14]. Defect [FB15, KCT12, LMP16, LSCK12, ZF13]. Defect-Density [ZF13]. Defects [ZK10]. Defense [Boo13d, Sca19]. Defined [PGG^{+}18]. Defining [LLC12]. Definition [TL18]. Deja [Erd10b]. Deliver [HBS17b]. Delivering [HJ17, TD19]. Delivery [CR14, Che15, SCW^{+}18]. Demand [BWN16]. Democracy [CPP16]. Democratize [Cab18]. Density [ZF13]. Department [Sca19]. Dependability [HPM16]. Dependency [KE10]. Depends [Gre18b]. Deploying [AHMS18, BMR14]. Deployment [GK15a, LMP^{+}15, NBF^{+}18, PHA^{+}17].

Depth [MCBT13]. Description [Woo14a]. Deserve [Ano16v]. Design [BvdHOP12, BBM18, BBB^{+}11, Boo11c, Bud13, Bus10b, CVEK13, DD12, Fai19a, HvG12, JWCR18, Kee18, LLC12, MBLN10, Mai10a, MT19, NYMS12, Ozk19b, PZA^{+}17a, RSS18, RI12, Sha12, SSS15b, TvV12, VAC^{+}17, WKV14, WGC^{+}13, Woo16a, ZCTZ13, Zim11]. Designer [Fai19d]. Designing [Gue14, PTNF^{+}19, RLBR12, Shu12b, Wat18a]. Desktop [GS11, KV11, PSK11, ZCLL19]. Desktops [LNH^{+}11]. Detection [RCLR18, WMM18]. Determine [MOMM11]. Determining [Scr18]. detisiveR [Ano11a]. Deus [Boo13a]. Developer [BHG13, CMY17, FCS^{+}10, MCD19, NHM19, Ozk19f, SSWM16, Spi15a, Spi16a, Spi16c, Spi18a]. Developers [ASR17, GWA14, Joh15, LA19, MAGT19]. Developing [CKH16, GF16, GP13, HBB^{+}11, OT11, RTV12, See12, SHR12, Spi14b]. Development [ASP13, Ano16-32, ADH14, AKN^{+}16, BHN14, BZd18, BBB^{+}19, BBS13b, BA10, BSW19, BNW10, CMM16, CdAC^{+}16, CMP17, CH17, CR11, CCWP11, CNSM13, DFP19, DFD^{+}16, DD10, EAO12, EBS16, ES16, FFGM13, FRM15, FSRdA11, GE11, GMP14, GUP^{+}14, GFPK10, Gue14, HR11, HSG18, HHLJ10, ILM11, IGBKR19, Jor13, Jor14, Jor19, Jun18, KT18, Kna19, KLM11, Kri17, KDM^{+}19, MCD19, MM13, MRS10, MS12, MIC19, NOS12, NS17, Obr13, OSBA14, PL19, PTNF^{+}19, PBE14, PC12, PAS10, PPMD17, Rie11, RCP^{+}12, SMT^{+}10, Shu11b, SV16, Som16, Spi13b, SAT14, Spi17c, SF15, TLvV13, TYXL16, VSB^{+}16, VB16, WKS^{+}14, WHR10, XYCW18, ZP11, ZF13, dSMdSC16, Sad14]. Developments [Car17b]. Devices [Dig15, RMM^{+}12, TS12]. DevOps [BHJ16, Bas18, CS16, Dör18, EGHS16, Ker18a, Ozk19a, SC18, Spi16a, TJC^{+}19, ZBCS16]. DevSecOps [Car17a]. Diagnose [Tou11]. Diagnosing [LXC15].

Did [Ano17n]. Difference

[RTV12, ZCLL19]. Difference-Bound [RTV12]. Different [Via18]. Differential [Ano13c, Spi13a]. Differentiates [ORM^{+}17]. Digest [PLC19]. Digital [Ano14n, Ano16-42, ED18, FSRdA11, HOZZ16, LL16]. Digital-Rights-Management-Sharing [LL16]. Dimension [KSW10, Poo16]. Dimensions [KTH11]. Disbanding [Shu12c]. Discovery [And16, Rou12]. Discussion [MF13, WCFG10]. Diseases [Bus10a]. Disentangle [MTM^{+}19]. Disks [Spi10b]. Dispersed [Ter11]. Disrupters [Boo16b]. Disrupting [Boo16b]. Disseminating [LCA17]. Distance [BH19, JWCR18]. Distinguished [Ano15b, Ano16b]. Distributed [ASB^{+}12, Bav12, Blu19, BSD16, CE19, CMY17, HSEG10, JWCR18, LDO^{+}14, PM10, PAS10, REV10, RH13, Sal19, VSB^{+}16]. Distribution [GK15a]. Diversification [ABB^{+}15b]. Diversity [BSW19, IHSR19]. Divided [BCD^{+}17]. Diving [Gor13]. Divorced [BH13]. Do [ASR17, CS16, CH14b, GUP^{+}14, Jor14, KT18, KSNH15, PHAH16, She15, SMT^{+}10, Sma18, Spi14c, VB16, MLD^{+}13]. DO-178C [MLD^{+}13]. Docker [And15]. Document [MHM19]. Documentation [Joh10, Kee15a, Spi10a, UR15]. Does [ASP13, Hol19d, JWCR18, Ozk19d]. Domain [BDD^{+}18, CH15b, Fre10, RSS18]. Domain-Driven [RSS18]. Domain-Specific [Fre10]. Dominant [Boo11c]. Dominating [Hol19a]. Donation [NHM19]. Don’t [CH14a, Jor14, SR16, SCC11, Sma18, Spi12b]. Doody [Bla17]. Doran [Jun18]. Doubt [Hol17e]. Draw [Boo11d]. Drawing [VD15]. Drive [FCN16]. Driven [Ano16-32, Ano18g, BWC10, BBN^{+}17, CM10, CFF^{+}13, CCD^{+}17, CH13c, DDIP15, DIP12, EHS15, EBS16, Gue14, KT18, LDO^{+}14, MNJR16, MBLN10, RSS18, SMT^{+}10, WHR14, ZP11].

Drivers [Spi11a]. Driving [HV16, Poo14]. Drupal [WKS^{+}14]. DSL [Gun10]. DSLs [EFO14]. Dumpster [Gor13]. Duration [Sym12]. during [SC18]. Dynamic [GFT15, ZKB^{+}16].

E-Voting [KV18]. Early [BCMN13, DD12, HvG12, KCK^{+}13, MM19]. Early-Stage [DD12]. Earned [BS15, Erd10f, HBS17b]. EARS [Mav12, MW19]. Ease [SJ13]. Easy [CS16]. Eccles [Mai11c]. Eckert [Ano17f, Ano19c]. Eclipse [BMSF15, KR10, NHM19]. Economic [CR14]. Economy [dMHSC16]. Ecosystem [BCN16, LDBF19, LKR16, SA19]. Ecosystems [AVF19, Bos16, BSS^{+}17, FWS19, GAM^{+}17, HKB16, JCP19, KHD19, MTM^{+}19, VMVS19]. Ed [Shu11b]. Edge [MOK^{+}18, Ano19g]. Editor [Fav10, Ozk19c, Shu13b, Shu14c, Spi18c]. Editors’ [FCAJ11, ASH13, BvdHOP12, BBM^{+}15, BCH15, EEBB11, FP11, GK15b, GKKL12, LAH10, LCF13, MMYJ10, MGFRD10, MCD19, MHRSO15, PSK11, PZTF12, SKK14, VK11, WC10, ZBOC12]. edoC [Ano11a, Spi11c]. Education [Ano16d, Ano17h, ABH^{+}11, BW14, HJP16, MGS18b, MSSMDC14, Off13]. Educators [BCD^{+}17]. Effect [MTM^{+}19, MZLD12, PCP14, SJS12]. Effective [PS14, PAS10, TBvRB12]. Effectiveness [APPV19, Men18, SMAU16, TvV12]. Efficiency [LXC15]. Efficient [FMMP16, MM18b]. Effort [FC19, Jor13, Jor14, Sym12]. Efforts [RVB13]. Elaine [CH13c]. Elderly [GVSP^{+}17]. Election [Ano19a, Ano19b]. Element [Shu13d]. Elephant [Boo10d, Hv19]. Elusiveness [Spi17a]. elytS [Ano11a, Spi11c]. Emails [HJ17, LA19]. Embedded [AMK16, ADH14, BO17, ED19, GFKY18, HHLJ10, LIL^{+}13, LDO^{+}14, LCYL14, LE13]. Embedded-Software [AMK16]. Embedding [BHN14]. Embracing [Bri12]. Emergence [LW11]. Emerging [Kee15b, LKR16, dOMFP16]. Emotion [NS19]. Emotional [MM19]. Emotions [CCP^{+}18, GWA14, MM19]. Empirical [CSC^{+}18, Dyb13, GGGM19, KZW15, LMP16, MAN^{+}14, RKW15, SJ13]. Employment [Kru15]. Empower [PL19]. Enabled [KV18]. Enables [BHJ16]. Enabling [BSS^{+}17, ILM11]. End [Ano17o, BCMN13, BDD^{+}14, Gat16, Nic19]. End-to-End [Gat16]. End-User [BDD^{+}14]. Ends [Hol18d]. Energy [BWN16, BW17, CYM^{+}15, CdAC^{+}16, CM19, FKL19, KHYT18, LXC15, PHAH16, ZHG14]. Energy-Aware [BW17, FKL19]. Engagement [SMN19, Spi15c]. Engine [PSO12]. Engineer [DMG14, Gre17b]. Engineering [ABB^{+}15a, ABB^{+}18, Ano14-32, Ano15n, Ano16-43, Ano17b, Ano17z, ABH^{+}11, BBB^{+}18, BBdR^{+}16, BCD^{+}17, BBM^{+}15, Blu19, Boo18, Bos16, Bri12, BBN^{+}17, Bro18, BW14, Cap14, CFF^{+}13, Car17b, CSC^{+}18, CP18, CS18, CS19, CCD^{+}17, DFFP18, DHH^{+}19, DDIP15, DIP12, EKP16, Ebe18, Erd10b, Erd10e, EM18, EMP18, FWK^{+}15, Fer18, FP13, GBM16, GCH13, Gre18a, Gre19, GO15, GSA^{+}17, GKKL12, HOBK17, Haz10, HSF^{+}16, HJP16, HV16, JEJ12, KAC^{+}18, Kim19, KUG18, KUG19, LvdH16, LEPV10, LCE13, LMP16, LMMO18, LBWX17, LCFT17, LK10, LRB^{+}15, MNJR16, MG17, MM18a, MGS18b, dOMFP16, MRS10, MSSMDC14, MHF17, MBv18, NBF^{+}18, NS19, NS17, Off13, Ozk19c, PDV^{+}18, RWZ10, RSH12, Sal18, Sca17, SdA13, SSE12, Shu13a, SCC^{+}16, SW11b, SCW15, Spi15f, Spi17b, Spi17d, Spi18b, SLB17, SPB16, TD19]. Engineering [VD10, VK11, VGD19, WHR14, WMM18, YFMH18, Zam17, ZBOC12, dGNA^{+}11, And15, Ano18c, Ano18d]. Engineers [ABB^{+}15a, BCB^{+}17, Bla17, DZB18]. Enhance [Als13]. Enhanced [NCK11]. Enhancing [LSCK12, Pel12]. Enough [BS15, Gre18a, Gre18b, Poo16]. Enterprise [Boo10e, ED19, FFZ^{+}18, Gho10, GP13, Ker18c, Lou16, PCRPE19, ZPHW16]. Enterprises [Tho15a]. Entities [LMMO18]. Entrepreneur [Ano16e]. Environment [BCB^{+}17, Spi13b, dMHSC16]. Environmental [ZML^{+}12]. Environments [GLA14, WZX^{+}16, ZXT^{+}15]. Equivalent [FC19]. Era [MT19, SMN19, TM17]. Ericsson [BSD16, PL19]. Errors [Ham18]. Escalations [WLD19]. Establishing [TBvRB12]. Estimating [CSS12, NCK11]. Estimation [BE11b, ES14, Jor13, Jor14]. Ethical [Boo14b, GM10]. Ethics [Ozk19b]. European [KDM^{+}19]. Evaluating [Ano17p, HSEG10, PSO12, dOdA16]. Evaluation [KZW15, OCFLI14, RO15]. Evasion [Hol15b]. Everyone [VGD19]. Everywhere [Shu14b, Spi10d]. Evidence [BHS^{+}14, BO17, DZB18, DS12, Dyb13, Erd10c, KCT12, KCD12, NS17, PM18, RKW15, SDSP17, SMP15, SW11b, SCW15]. Evidence-Based [BHS^{+}14, BO17, KCT12, NS17]. Evolution [Ave11, CPS18, CS18, DIP12, EFO14, HSG18, MGFRD10, MIC19, ZK10, Zim15, dOdA16]. Evolutionary [ZF13]. Evolves [WF18]. Evolving [DeM14, PAB11, Spi18d, TI11]. ex [Boo13a]. Exactly [Mai12b]. Examining [KNH16, RNA^{+}16b]. Excellence [Ano18-37, Ano18-38, Fav12]. Exchange [BBS13b]. Execution [Dig15, GE17, WZX^{+}16, YLK^{+}17]. Executions [MOK^{+}18]. Exemplar [MGS18b]. Exercising [AVF19]. Experience [Boo12c, Bro19, BDD^{+}18, GRDL^{+}12, GTS^{+}19, MLD^{+}13, PT18]. Experiment [KLM11, Obr17]. Experimentation [SCL18]. Expertise [LL16, RL16]. Experts [Ano14-33, Ano15-30, Ano15-36, Ano16-50, SPB16, Ano17-52]. Exploiting [HTE16]. Exploratory [AGIF17]. Exploring [CH13c, IML16, NBS^{+}16, RKTS12, Sym12]. Explosion [Ker18a]. Exposing [SM10a]. Expressiveness [Fai19e]. Extending [CNZC17, HAE11, HM13, Spi15b]. Extensible [EFO14]. Extension [LLC12]. EYE [VD15]. Eyes [SSWM16].

Faced [ZSMP17]. Faces [MZ13a]. Facilitate [LSCK12]. Facilitating [ASK14, HKR11, SW11a]. Facilitators [Wri11]. Facing [Boo12b]. Factor [Cap14, CPSY17]. Factors [BA10, Sca19, WAA^{+}12]. Facts [ACHC11]. Fad [VE12]. Fail [Blu16]. Failing [SB16]. Fails [UR15]. Failure [Bus10a, Bus10b, Sal19]. Failures [AGIF17]. Fain [Far17]. Fair [BDH^{+}19]. Faking [Spi11d]. Fall [EV10]. Falling [Bus10b]. Far [JPM^{+}18]. Farewell [Spi10b]. Farming [BCN16]. Fast [FS12]. Faster [BMDK15]. Fault [HBB^{+}11, Hol14]. Fault-Prediction [HBB^{+}11]. Favorite [Ano16-35]. Feature [RSH12]. Feature-Oriented [RSH12]. Featuritis [Bus10a]. Feedback [GAM^{+}17]. Feminine [RL16]. Fernbach [Ano18-37, Ano18-38]. Fewer [Spi15c]. Field [Spi15b]. Fields [Til16a]. Fifteen [MH14]. Fighting [CH13a]. Figures [EV10]. Final [Ris14]. Financial [BA10]. Find [PDV^{+}18]. FindBugs [KNH16]. Finding [KB12, RPP19]. Fingerprinting [KZW15]. Fingertips [Shu12b]. Finite [Shu11c]. Fire [CH14a]. Firm [MS14]. First [CCSW10, Spi14c]. Fit [Ano16w, Ano16-33, Ano16-34, Ano17y, Ano17-38, Ano17-35, Ano17-36, Ano17-37, Ano16c, Ano17e, Ano17d, Ano17-34]. Five [ACHC11, BMSF15, BH10a, BH10b, Ker18b, Roz15]. Fixing [Jun18]. Flexible [ABC^{+}11, Kee15b]. Flight [Ave11, Jep13]. Flow [NOS12]. Flows [Ker18d, SEM14].

Floyd [RTV12]. Focus [Ano14d, Ano14c, Ano15g, Ano16g, Ano16f, Ano17q, GF17]. Focused [Bus10d]. Fog [MOK^{+}18]. Forces [FTP11]. Ford [WCFG10]. Forgotten [Woo16b]. Form [DZB18]. Formal [BBC^{+}19, FFGM13, MLD^{+}13]. Format [EJ12]. Formative [RI12]. Formatting [CCCM17]. Forward [Ano15q, Ano15r, Shu12e]. Foster [BBB^{+}18]. Four [KTH11, MSS10, MONF16]. Framework [ABC^{+}11, BBB^{+}11, GP13, Gue14, MGMS11, NH15, Pel12, Scr18, SEK^{+}19, VSB^{+}16, VMVS19, ZXL^{+}16]. Frameworks [CVEK13, CC19, SS17]. Framing [LAH10, Mai12c]. Francois [Car17a]. Free [GBICMR13, MSAH17, RvGH17]. Frequency [MHM19, ZXT^{+}15]. Frequently [Hol16c]. Friction [AKN^{+}16]. Frictionless [Spi13b]. Friend [Bus11c]. Friends [BH13]. Front [Ano14e, Ano14f, Ano14g, Ano14h, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x]. Frontier [DFP19]. Fun [Shu11b, VK11]. Functional [ES14, Ebe15a, AACF13]. Future [ABB^{+}15a, Ano15n, Boo12b, Bos16, Ebe15b, HSF^{+}16, HJP16, LRB^{+}15, MM18b, dOMFP16, MCD19, PTNF^{+}19, Shu11a, SCC^{+}16, TYXL16, TMB17, VWK15].

Game [DK11, FSRdA11, Sca17, PMSL14]. Gameplay [CMS11]. Games [BO17, LW11, Shu13a]. Gamification [Bas17]. Gaming [OT11]. Gap [Gla12, GJO^{+}18, Ris12b]. Gaps [Woo15b]. Gardening [Bus11a, Bus11b]. Gender [BSW19, CCP^{+}18, CS19, Gla12, IHSR19, Ris12b]. Generalizability [BBN^{+}17]. Generation [BH17, BMS^{+}14, EAG^{+}14, MW11, Ris12a, YLK^{+}17]. Generational [RSH13]. Genuine [HJ17]. Geographically [RH13, Ter11]. Germany [FWK^{+}15]. Gestures [HLS17]. Get [Ano16v, BWN16, SR16, Sym10]. Getting [Osb11, Shu13c]. Ghosts [CT13]. Gilad [TW14a]. Git [Spi12c]. GitHub [BBS13b, CCCM17, ZCLL19]. Giving [Bro19]. Glass [Ano10b]. Global [BCD^{+}17, BCB^{+}17, BNW10, CHL14, EKP16, LEPV10, LCE13, MS14, SW11b, SCW15]. Go [Bus12a, Mey14b]. Goal [BB17]. Goals [CB10, HBCH14, Ano16c, Ano17e, Ano17d]. God [PCP14]. Goetz [WCFG10]. Going [vMM12]. Golden [Ozk19c]. Gone [Bus12a]. Good [BZd18, BH13, DHH^{+}19, Erd10e, vGH19]. Goodies [Spi13d]. Google [MSAH17, PSO12, Shu12a]. GORE [Ale11]. Got [BWN16]. GOTO [Nag18]. Governance [BH19, CR14]. Governing [KHD19]. Government [Lob11, SCW^{+}18]. GPU [FTP11]. Grace [Boo15a]. Graceful [Gre17a]. Grail [GMP14]. Gram [MHM19]. Graphics [Ano19f]. Gratae [CH14b]. Great [Boo13c]. Green [ASK14, BMR14, BMM14, SMS^{+}14]. Greener [ASK14]. Greening [SMS^{+}14]. Ground [Ker18c]. Group [LCE13, PSGd18, RRL^{+}17, TBvRB12, dSMdSC16]. Group-Based [RRL^{+}17]. Grow [Erd10a]. Grownups [LMMO18]. Growth [Rut16, vGH12]. Guarantee [Shu13b]. Guest [ASH13, BvdHOP12, BBM^{+}15, BCH15, EEBB11, FCAJ11, Fav10, FP11, GK15b, GKKL12, LAH10, LCF13, MMYJ10, MGFRD10, MCD19, MHRSO15, PSK11, PZTF12, SKK14, VK11, WC10, ZBOC12]. GUI [AGIF17, Als13, HLS17, LCYL14]. Guide [SC18]. Guidelines [HL11, KCT12, SAT14, STG19, TSPB17]. Guilds [SMLF19]. Guilty [MOMM11].

Hackathons [KPR^{+}15, PTNF^{+}19].

Hadoop [RK13]. Haiku [Kee15a]. Hair [Spi13d]. Half [MGS18b]. Hallucination [Boo10b]. Hammers [Bus10b]. Hand [Ano15f, Ano17l, Ano18h, Spi12b]. Handling [Shu11b]. Handshaking [FGBS10]. hardware [Sun19]. Harlan [Ano17x]. Harm [Spi14c]. Harmful [Nag18, PCP14]. Harnessing [IML16, Woo14a, Woo16a]. Harsh [Ris12a, Rei18]. Haystack [KB12]. Headed [MONF16]. Health [GVSP^{+}17]. Healthcare [LLV16, Li14, NS17, SWB^{+}17, Spi17a]. Healthy [CPB^{+}17, Fai19b]. Heavenly [Spi18e]. Heights [Ano15-36, Ano16-50, Ano17-52]. Helland [Sal19]. Helped [MRF^{+}13]. Here [BWN16, UG18]. Heterogeneous [BMS^{+}14, WZX^{+}16]. Heuristics [Gre18b]. Hey [TJC^{+}19]. Hi [Hol16d]. Hierarchy [WGC^{+}13]. Higgs [Rou12]. High [Ano18-37, Ano18-38, GF16, GO15, PSO12, Ron16, ZXT^{+}15]. High-Availability [Ron16]. High-Frequency [ZXT^{+}15]. High-Impact [GO15]. High-Performance [PSO12]. High-Quality [GF16]. Higher [KE10]. Highlights [CS17, CM19, ZK14]. Highways [LMP^{+}15]. Hilton [Her17]. Hindman [Mey16]. Hiring [Kaa14]. History [Ano19d, Ano19e, Boo18, Obr17, Obr18]. Hoisting [Fai14]. Holy [GMP14]. Hope [Shu14d]. Horizontally [Fai19f]. Hospital [RLBR12]. Hourly [SvS16]. House [Ano14d, Ano14x, Ano14-31, Ano14-32, Ano15e, Ano15f, Ano15n, Ano15o, Ano15q, Ano15r, Ano15b, Ano15-27, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15-35, Ano15-36, Ano16g, Ano16n, Ano16t, Ano16u, Ano16r, Ano16y, Ano16-37, Ano16-39, Ano16-41, Ano16-35, Ano16-44, Ano16-43, Ano16-50, Ano17e, Ano17l, Ano17q, Ano17x, Ano17-41, Ano17-44, Ano17-38]. Howard [Hen19]. HTML5 [TS12]. Huge [Che15].

Human

[Boo12c, Boo14b, Boo16a, Cap14, CPSY17, CMP17, Haz10, Hol15f, IML16, Shu13d]. Humphrey [Shu11f]. Hurricane [Ric11]. Hybrid [GKD^{+}15, KDM^{+}19, NKR^{+}18, Rie11, SMN19]. Hygge [VBP17].

ICSE [CS17, CM19]. ICT [CPS18]. IDE [SSWM16]. Ideas [Erd10b, Erd10e]. Identifying [APPV19, FRM15]. IEEE [Ano17f, Ano18a, Ano18b, Ano14j, Ano14k, Ano14i, Ano14l, Ano14n, Ano14m, Ano15o, Ano16n, Ano16t, Ano16u, Ano16x, Ano16p, Ano16v, Ano16r, Ano16q, Ano16s, Ano16w, Ano16o, Ano16y, Ano17y, Ano17x, Ano18p, Ano18q, Ano18t, Ano18r, Ano18s, Ano18u, Ano19c, Ano19y, Ano19z, Ano19-27, Ano19-28, Ano19a, Ano19b, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano19-34, Obr17, Som16, Ano18e, Ano19d, Ano19e, Ano19g, Ano19f, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Obr18, Spi18f]. IEEE-CS [Ano18a, Ano18b]. If [Ozk19d]. Ignore [Fai19c]. III [Bus10b]. ility [LLV16]. Illustrative [PTV^{+}10]. I’m [Mai10b]. Imaging [ZF13]. Impact [BBB^{+}18, BSW19, Gim14, GO15, HWBYZ13, MCL12, RNA^{+}14, RSH13, SvGH15, SJ15, TMB17, ZHG14, vGH13, vGH15, vGH19]. Impacts [Wat18b, dMHSC16]. Implementation [FGBS10, MPD14, SCL18]. Implementations [RO15]. Implemented [BvD10]. Implementing [CC19, Ebe15a, RLBR12]. Implicit [SEM14]. Importance [Spi13c, Spi15f]. Important [Erd10c]. Impressions [CCSW10]. Improve [HL11, MRF^{+}13, NOS12]. Improvement [GFH18, LOCPL16, LSCK12, MSSMDC14, SC18, UG18]. Improving [FSRdA11, HMPS16, HOBK17, LCYL14]. In-Depth [MCBT13]. Incident [BBG19, PGG^{+}18]. Incorporating [DD10].

Increased [Bre14]. Incredible [Boo14c]. Incremental [Erd10a]. India [See12]. Indian [Sad14]. Individual [Nic19]. Inducing [CCMT16]. Industrial [CH15c, GF17, MLD^{+}13, Pec13, PGG^{+}18, SSS15a]. Industries [EHS15]. Industry [BBB^{+}18, CP18, EKP16, ES17, Fer18, FFGM13, GTS^{+}19, Gre17b, HBB^{+}11, Kaa14, LMG16, LDBF19, MONF16, Pop11, PT18, SPA11, SSE12, Sym10, WCFG10, WAA^{+}12]. Industry-Academia [SPA11]. Inescapable [Mai12d]. Inference [BFKZ15, SSE12]. Inflation [Hol15c]. Influence [ASP13, dSMdSC16]. Influencer [FWS19]. Information [Ano19z, Ano19-27, Ano19-28, Ano19-29, Ano19-30, CH13a, Fre10, GAM^{+}17, Hol19a, LDO^{+}14, SAT14, VMVS19, vdBB17, Ano19y]. Infotainment [CdAC^{+}16]. Infrastructure [HGS15, Joh17, SGH15]. Infrastructure-Monitoring [HGS15]. Infrastructures [Spi14a]. Infusing [Lat12]. Ingredients [Mar12]. Inhibited [Mai11a]. Initiative [Fer18]. Initiatives [PDNdSB12]. Injecting [CH15a]. Inland [vdBB17]. Inner [SF15]. Innovating [Joh16]. Innovation [BBG19, Bus13, EBS16, GFPK10, Obr13]. Innovation-Driven [EBS16]. Innovative [dSMdSC16]. Ins [KR10]. Insight [BBG19]. Insights [BGK15, Obr17, PZ15, RKW15, RBE19]. Install [Spi12b]. Integrate [BCMN13]. Integrating [LC11, MRS10]. Integration [GRDL^{+}12, GHV10, Mey14a, PZA^{+}17b, PZ18, ZPHW16]. Intel [KV11]. Intelect [Ano14o, Ano15p]. Intellectual [Fai19d]. Intelligence [Ano18g, BDH^{+}19, MCH15, ZWGY12, ZXL^{+}16]. Intelligent [Ano19h]. Intelligently [HKB16]. Intensive [GLA14, GFPK10, KUG18]. Interact [GBICMR13, Ozk19e]. Interaction [LLC12]. Interactions [Mad10, MBLN10]. Interactive [NCK11]. Interactive-Software [NCK11].

Interchange [EJ12]. Interests [Con12]. Interface [Ano18i, MM13]. Interfaces [BDD^{+}14, Shu12b, Tsa11]. International [Ano17b, Ano17z, Ano18c, Ano18d]. Internet [And12, Ano16-43, BW17, BBM^{+}15, BO17, CMM16, ESZ15, FMMP16, LCFT17, LRB^{+}15, Spi17d, WE16]. Internet-Scale [BW17]. Internetware [BBM^{+}15, LXC15, ZZJJ15]. Interoperability [BSS^{+}17]. Interpreters [Ano17m]. Interview [Fav12, Shu11b, Shu12a, TW14a]. Interviews [MCBT13]. Intolerance [Hol14]. Introducing [SJG13]. Introduction [BvdHOP12, EEBB11, FCAJ11, Fav10, FP11, Gre17b, GKKL12, LAH10, MMYJ10, MGFRD10, MCD19, PSK11, PZTF12, VK11, WC10, ZBOC12, ASH13, BBM^{+}15, BCH15, GK15b, LCF13, MHRSO15, SKK14]. Introductions [Spi15d]. Intuition [Shu13c]. invent [Ano17j]. Inverse [MHM19]. Investments [SNMF13]. invokedynamic [OCFLI14]. Involvement [Ano14-35, BZd18]. Involving [ASP13]. IoT [BSS^{+}17, CCD^{+}17, MHF17, TM17, TM18, VAC^{+}17, Zam17]. IoT-Oriented [Zam17]. Issues [LMG16, TJC^{+}19]. Iteration [WGC^{+}13]. Iterative [Erd10a, SJG13]. iUCP [NCK11].

James [And16, Shu12a]. Java

[BH17, BMDK15, Bla18a, GD12, GS11, MSS10, OCFLI14, ZKB^{+}16]. JavaScript [Kie10, RPP19, TT11]. Jay [Til16a]. Job [Ano14d, Ano14c, Ano16g, Ano16f, Ano17q, Ano18v, Ano19-35, SS18, Ano15g]. Jobs [Ano19-31, Ano19-32, Ano19-33, LTv19, LAAN10]. John [Jun18]. Johnathan [Kim19]. Joint [FTP11]. Jonathan [Bel19]. Josh [Bla17]. Journey [Boo11a, Gre17b, Gue14, JPM^{+}18, DK11]. July [Shu11f]. Jungle [MHF17]. Jürgen [Blu16]. Just [BH13, GE17, GTS10, Poo16, SR16].

Kanban [SJS12]. Katie [Sal17]. Keep [Ano15q, Ano15r]. Keeping [CHDK^{+}16]. Ken [Ano18-37, Ano18-38]. Kennedy [Ano18-37, Ano18-38]. Kernel [KB12]. Key [BBC^{+}16, CCWP11, Zam17]. Kicking [CH13b]. Kief [Joh17]. King [Tor15]. Kirks [Mai12e]. KLOVER [YLK^{+}17]. Know [CH14b, GFH18, GFKY18, GKF19, GUP^{+}14, Jor14, KT18, PHAH16, Ris10c, She15, SMT^{+}10]. Knowledge [CNZC17, CH15b, IML16, ORM^{+}17, PvdHM15, SMLF19, WG16, Ano17-41, Ano17-40, Ano18-34, Ano18-35]. Knowledge-Based [PvdHM15]. Kudos [Ano10b].

Laartz [Blu16]. Labels [NKR^{+}18]. Lander [BBP^{+}16]. Landing [Bal18, Hol13]. Landscape [AS17, CMP17, GSA^{+}17]. Language [BBB^{+}19, CCMT16, DFFP18, FDE^{+}17, MPD14, Mey14b, PB16, Vol10, WKV14]. Languages [Ano14-34, EFO14, Fre10, HAE11, LMM^{+}15, TW14b]. Large [BH19, Blu16, BSD16, CC19, DHB15, Kna19, MAN^{+}14, NH15, OSBA14, PL19, SCW^{+}18, SMLF19]. Large-Scale [BH19, BSD16, CC19, Kna19, MAN^{+}14, NH15, OSBA14, PL19, SCW^{+}18, SMLF19]. Late [Ano12a, DeM11]. Latent [TLvV13]. Lattice [Che14]. Laughing [Boo16d]. Law [KCD12, WK15]. Laws [dOdA16]. Layer [GRDL^{+}12]. Leaders [HSF^{+}16, MONF16, dSMdSC16]. Leading [Ano15-36, Ano16-50, MSS10, RH13, Ano17-52]. Lean [DS12, EAO12, Kee15a, MNS12, NOS12, PC12, Tho15a, VE12]. Learn [Ric11, WLD19]. Learned [BHN^{+}10, Bal18, HR19b, Jep13, LW11, MS12, NS17, RH13, VD10, Wri11]. Learning [BHN14, BNW10, Bus10a, Bus10b, Car19, ES16, FLC17, HBN13, KAC^{+}18, Kru15, LMG16, LE16, Mai12c, MHM19, Sal17, SMP15, TAJB17]. Least [Fai19e]. Led [UG18]. Legacy [BJ11, Bel19, FFZ^{+}18, Joh16, KH18, SHG14]. Legal [CCMT16]. Lehman [dOdA16]. Less [Gor13]. Lessons [BHN^{+}10, Bal18, BO17, CMM16, Jep13, KLM11, MM19, MS12, NS17, RCP^{+}12, RH13, SHR12, SS11, VD10, Via15, Via19, Wri11, HR19b]. Let [Hin19, STG19]. Letters [Ano13c]. Level [FSM14, SAO^{+}17]. Leverage [PDV^{+}18]. Leveraging [DSTH13, EBS16, HAE11, LBL17, MRF^{+}13, PGG^{+}18]. Libraries [RNA^{+}14, Spi12a]. Library [OCFLI14, RNA^{+}16a]. License [Eng10, GD12]. Licenses [LRP11]. Licensing [MOMM11]. Lichty [Bla18b]. Life [Bus12b, Erd10b, Fav10, LIL^{+}13, SJG13, PMSL14]. Life-Cycle [LIL^{+}13, SJG13]. Lifelong [Kru15]. Lifetime [Shu13b]. Lightness [Boo14c]. Lightning [Hen19]. Lightweight [BvD10, GFPK10, Kee15b]. Like [Val17]. Line [Ano17o, HvG16, Joh10, JN10, LK10, MMYJ10, MRS10, Mor10, SNMF13, SdA13]. Line-of-Code [Mor10]. Lines [BCS10, Bos10, CdAC^{+}16, CSC^{+}18, FSRdA11, GFT15, dMSNRdCM^{+}11, dOdA16]. Linguistic [GHV10]. Link [Hol16f]. Linked [PZ18, VD15]. lint [Gim14]. Linux [RR15]. LISA [Ano15-35]. Listen [Mav12]. Listing [Ano19-42, Ano19-43, Ano19-44, Ano19-45]. Lists [LA19]. Lived [CGE19]. Living [GVSP^{+}17]. Log [MHLCL16]. Lonely [Boo16c]. Long [CGE19, SNMF13, Stu18]. Long-Lived [CGE19]. Long-Term [SNMF13]. Longevity [CNZC17]. Look [Par18, Rut16, Shu11a, Shu12b, SDSP17]. Looking [Ano18v, Ano19-35, Ebe15b, GMP14, Shu12e]. Loom [Boo12f]. Love [Bus10b, LW11]. Loving [Boo15a]. Luddite [Rut16]. Lunch [RvGH17]. Lync [MRF^{+}13].

Machina [Boo13a]. Machine [DPCL14, KAC^{+}18, LE16, MHM19, Sal17, WLD19, WSE14]. Machine-Learning [KAC^{+}18]. Machine-to-Machine [WSE14]. Machines [Boo15a]. Made [Cor13, Jep13]. Magazine [Ano14n, Ano19g, Ano19h, Ano19j, Ano19o, Ano19-46, Ano19-47]. Magic [Bud13, DD10]. Mailing [LA19]. Mainframe [VD10]. Mainstream [RSAT19]. Maintain [APPV19]. Maintainers [APPV19]. Maintaining [FSM14, GF16]. Maintenance [CS18, ES14, Hol16d, PT18]. Major [Ano16p, Ano16r, Ano16q]. Make [Hin19, Stu18]. Maker [WG16]. Makes [Kle16]. Making [AACF13, CS16, Mey18, NOS12, PL19, PSGd18, RBE19, Shu11b, SGSM10, WL12]. Malone [Sal17]. Managed [GP13, LDO^{+}14]. Management [AtHR11, Ave11, Cab18, Ebe14, Fav10, GVE11, Kim19, LL16, LWCP17, Lob11, Lou11, MNS12, NOS12, PS14, RCLR18, Rei18, SMP15, Spi12d, TD12]. Manager [WG16]. Managers [Ano17p, KHD19]. Managing [BCS10, BH19, Bla18b, CCSW10, GD11, GGGM19, JCP19, KHYT18, LI12, MCHK17, PAB11, Pel12, PM10, RKW15, RDMA11, Shu11b, Spi16d, TSPB17, Tsa11, VMVS19, vGH15]. Manifesto [FKL19, Pen15]. Manufacturing [Ano17o, PvdHM15]. Manufacturing-Line [Ano17o]. Many [MZ13a]. Mapped [Hen19]. Mapping [GV11, Via19]. Market [DeM14, Sad14]. Marketplace [McG11]. Marriage [Spi18e]. Married [BH13]. Mars [Ano13d, Hol13]. Masking [GKD^{+}15]. Mass [MS14]. Massively [DK11]. Masthead [Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-30, Ano16-31, Ano17-27, Ano17-28, Ano17-29, Ano17-30, Ano17-31, Ano17-32, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-28, Ano19-36, Ano19-37, Ano19-38, Ano19-39, Ano19-40, Ano19-41]. Matching [LRP11]. Material [LCA17]. Matlab [GRDL^{+}12]. Matrices [RTV12]. Matter [Boo16d, JWCR18, VB16, WL12]. Matters [Ano14u, Ano14v, Ano14w, Ano14x, BB12, Jor13, Ron16]. Maturity [FSM14, FRM15, GFH18, Gre17a, vWHZ^{+}10]. Mauchly [Ano17f, Ano19c]. maximize [Sun19]. Me [Boo11d, Mai10b, Spi12f, STG19]. Meaningful [CH15c, Gre19]. Means [Hol18d]. Measure [NKR^{+}18, Sym10]. Measurement [HPM16]. Measuring [CCC^{+}19, KDGV13, LA19]. Mechanisms [RRL^{+}17]. Media [BFKZ15]. Medical [RCLR18, ZF13]. Meet [CH13c]. Meets [BBS13b, BFKZ15]. Member [dSMdSC16, Ano17g]. Membership [Ano14l, Ano14u, Ano14v, Ano14w, Ano14x, Ano16-33, Ano16-34, Ano17-38, Ano17-35, Ano17-36, Ano17-37, Ano17-34, Ano17-41, Ano17-40, Ano18-34, Ano18-35]. Memory [Boo14d, Hen19]. Memory-Mapped [Hen19]. Merwin [Ano15b, Ano16b]. Mesos [Mey16]. Metaphor [KNO12, Woo15b]. Metaprogramming [DD10]. Meteorologist [Ric11]. Method [GD12, LI12]. Methodologies [ILM11, SSE12, VB16]. Methodology [SMN19, SS18]. Methods [BBC^{+}19, FFGM13, Mey18]. Metric [Mor10]. Metrics [CM10, CMS11, Gre19, HPM16, MOMM11, vGH13]. Microblogging [WKS^{+}14]. Microservice [RSS18, TL18]. Microservices [BHJ16, BDD^{+}18, FFZ^{+}18, JPM^{+}18, KH18, LSCPE18, PZA^{+}17a, PZA^{+}17b, Tho15b]. Microsoft [CNSM13]. Middleware [BMS^{+}14, DK11]. Midlife [Gre17a]. Might [PPMD17]. Migrating [FFZ^{+}18]. Migration [BHJ16]. Mileage [vGH11]. Miller [Ano11b]. Mills [Ano17x]. Min [CM12]. Mind [TJC^{+}19]. Minds [Boo13b].

Minecraft [Boo13b]. Mining [APPV19, CH15b, Hol19b, Ker18c, LAAN10, Men13, vGMRW14]. Minute [Whi12]. Miracle [Boo15f]. Missing [Kna19]. Mission [CCD^{+}17]. Mission-Critical [CCD^{+}17]. Mistakes [Jep13, SMP15]. Mitigation [BMSS18]. MLOC [TS11]. MobiGUITAR [AFT^{+}15]. Mobile [AFT^{+}15, Ano14-30, Ano14-31, BMS^{+}14, Bou10, BCN16, CM19, CPB^{+}17, Dig15, EAG^{+}14, GFT15, GE11, GAM^{+}17, GMB^{+}14, KSNH15, LXC15, MHJ17, MAN^{+}14, RMM^{+}12, RNA^{+}14, RNA^{+}16b, Sad14, See12, SJ15, SHG13, Shu12b, TS12, VMvG12, ZCLL19]. Mobile-App [BCN16, RNA^{+}16b]. Mobile-centric [GMB^{+}14]. Mobility [Shu14f]. Model [AGK^{+}19, AFT^{+}15, CFF^{+}13, CCD^{+}17, CR11, DDIP15, DIP12, FSB^{+}12, FFGM13, GCH13, GMB^{+}14, HR11, KCK^{+}13, LDO^{+}14, MHF17, RSS18, Rie11, SJG13, Sch12, Shu11a, Til16b, WHR14, ZP11]. Model-Based [AGK^{+}19, AFT^{+}15, FSB^{+}12, FFGM13, MHF17, Sch12]. Model-Driven [CFF^{+}13, CCD^{+}17, DDIP15, DIP12, LDO^{+}14, RSS18, WHR14, ZP11]. Modelica [HAE11]. Modeling [Ano16-32, BDD^{+}14, Che14, CMC^{+}12, DBRV18, ILM11, MCHK17, MBv18, PSGd18, Vol11, ZML^{+}12]. Models [CSC^{+}18, HBB^{+}11, HR11, KSTS17, LvdH16, LMG16, LRP11, OT11, Pop11, RDMA11, TAJB17, vWHZ^{+}10]. Modern [Til15]. Modernization [CM10, KH18, Sdd^{+}18]. Modular [Stu18, VWK15]. Modules [Bla18a, KE10]. Money [SJ15]. Monitor [GVSP^{+}17, Tou11]. Monitoring [HGS15, Mai13a, VRG16, ZK10]. Monolithic [BDD^{+}18]. Montezuma [Shu11b]. Mood [BSW19]. Moods [GWA14]. Moore [WK15]. Moral [GD11]. Morris [Joh17]. Most [Osb11]. Motivations [LvdH16]. Moving [Ano15q, Ano15r, Lou10]. MRI [HvdL10].

MSDN [BBS13b]. Much [SMS^{+}14]. Multi [Ano15-36, Ano16-50, Gun10, Ano17-52]. Multi-Core [Ano15-36, Ano16-50, Ano17-52]. Multi-DSL [Gun10]. Multiagent [WG10]. Multicore [KV11, LNH^{+}11]. Multidimensional [Ano17p]. Multimedia [Bou10]. Multinational [AHMS18]. Multiparadigm [Gho10, WC10, WCFG10]. Multiplatform [ZKB^{+}16]. Multiplayer [DK11]. Multiple [HHLJ10, SRI14]. Multitenancy [FFZ^{+}18]. Multiterabyte [KB12]. Multithreaded [BBdH^{+}11, FTP11]. Multitier [ABB^{+}15b]. Must [BWD^{+}17, BS15, HR19b]. Mutation [Als13, RPFA^{+}14, UM10]. My [Boo15d, Kri17, WGC^{+}13]. myCS [Ano17-33, Ano18-29, Ano18-30, Ano18-31, Ano18-32, Ano18-33]. Myth [Bud13, LR12, Nic19]. Mythical [ADSS18]. Myths [HO15].

Nails [Bus10b]. Naivete [GV11]. Naming [FWK^{+}15, Her17]. NaPiRE [Fer18]. Nate [Bha18]. Native [BHJ16]. Natural [DFFP18, FDE^{+}17]. Nature [GO15]. Navigation [SA11]. Neal [WCFG10]. Nearby [NBS^{+}16]. Need [BHG13]. Needle [KB12]. Needs [Fer18]. Negative [LA19]. Negotiating [FGBS10]. Negotiation [Bla17]. Neo [Rut16]. Neo-Luddite [Rut16]. Network [MOK^{+}18]. Networked [HHLJ10]. Networking [BBS13a, BBS13b, PGG^{+}18, Shu13d, Ano19-34]. Networks [BMS^{+}14, ESZ15, FN14, Li14, PvdHM15, PSGd18]. Neutral [RMM^{+}12]. Newcomers [STG19, ZSMP17]. Next [BMS^{+}14, DFP19, EHS15, EAG^{+}14, GCH13, HHR^{+}13, MW11, MZ18, Ozk19a, PST^{+}17, Ris12a, UG18]. Next-Generation [BMS^{+}14, EAG^{+}14, Ris12a]. Nice [Ano13e, Ris12c]. Niche [ORM^{+}17]. Nicolai [Bla18a]. Nightingale [Kim19]. Nimble [RBE19]. NLP [FC19]. No [Boo16d, Bus12a, Mar12, RvGH17, Spi14c]. Node.DPWS [FMMP16]. Nominations [Ano16d, Ano16p, Ano16r, Ano16q, Ano17h, Ano17i, Ano18f, Ano18-37, Ano18-38, Ano19c]. Nominees [Ano16d, Ano17h]. Non [AACF13, CH14b]. Non-functional [AACF13]. Nonfunctional [HL11, PRRT14, SHR12]. Nontechnical [Car19]. Norms [TBvRB12]. Notkin [Ano13b]. Nutshell [SAO^{+}17].

Object [GTS10, GS11, Hof10, RRL^{+}17, Via19]. Object-Oriented [GTS10, GS11, Hof10, RRL^{+}17]. Object-Relational [Via19]. Objects [SJ13]. Observations [CMC^{+}12]. Obstanovka [NBS^{+}16]. Occurs [Boo15f]. Ocean [HJ17]. October [Shu11f]. Office [TS11]. Offs [SM10a, Sym12]. Offshoring [SvS16]. Often [BCMN13, MM19]. Oh [LC11]. On-Board [Tsa11]. On-Premise [BKB12]. Onboard [BBP^{+}16]. Onboarding [FGBM14, STG19]. Once [Boo16e]. One [Ano17-41, Ano17-40, Ano18-34, Ano18-35, BBC^{+}19, Boo15c, Bus12a]. One-Click [BBC^{+}19]. Online [BO17, DK11, Hol19a, ZXL^{+}16]. Only [AMK16, Shu13e]. Open [AtHR11, ACHC11, CCSW10, Eng10, FGBM14, GD12, HR11, ILM11, LRP11, LKR16, MTM^{+}19, MFS15, MOMM11, RCP^{+}12, RSAT19, Sca19, SvGH15, Spi11b, STG19, SF15, WKS^{+}14, dBLMT11]. Open-Data [LKR16]. OpenStack [IHSR19]. Operating [Tor15]. Operational [BNW10, MHLCL16, PLC19, Woo16b]. Operational-Log [MHLCL16]. Opinion [SAT14]. Opinions [DZB18]. Opportunistic [KLM11, MT19]. Opportunities [SA19, TD19]. Optimica [HAE11]. optimizes [Sun19]. Options [Ano16-33, Ano16-34, Ano17-38, Ano17-35, Ano17-36, Ano17-37, Ano17-34]. Orchestration [Spi14d]. Order [Boo15b]. O’Reilly [Tho15a]. Organization [AHMS18, FSM14, PL19, SCW^{+}18]. Organizational [PDNdSB12, VB16]. Organizations [BO17, LOCPL16, Lat12, SMLF19, SF15]. Orientation [LK10, LMSS11]. Oriented [GTS10, GS11, Hof10, LR12, MBLN10, MGMS11, MH14, RRL^{+}17, RSH12, SHG14, Zam17]. Original [HR19a]. Orthogonal [CT11]. Other [Bus10a]. Our [Boo15c, CHDK^{+}16, HBCH14, LC11, LTv19, Mai13a, Ozk19a, Ris10b, Shu14d, SMP15, Spi15b]. Ourselves [Boo14e]. Outdoors [GVSP^{+}17]. Over-the-Air [VOGB18]. Overcome [Hol19a]. Overflow [ASR17, SB16]. Overload [Hol19c]. Overlords [Boo15c]. Overly [LCH17]. Overrated [BBN^{+}17]. Oversight [Roz15]. Overview [MG17, PTV^{+}10]. Ownership [BH19, EM15]. Oxymoron [Boo10a]. Oz [Boo13c].

Package [Spi12d]. Pages [FS12]. Pain [FWK^{+}15]. Pair [Ano10c, SMAU16, Wra10]. Palm [Ano15f, Ano17l, Ano18h]. Panel [MF13]. Papers [Ano14b, Ano14i, Ano14-34, Ano15e, Ano15d, Ano16n, Ano15c]. Paradigm [KC16, VGD19]. Paradigms [Bus10c, DD10, HHLJ10]. Parallelisation [GS11]. Parallelism [Dig11, PSK11]. Parallelizing [VM12]. Parlog [Bla18a]. Part [BH10a, BH10b, Bus10a, Bus11a, Bus11b, Bus11e, Bus11f, HBS17a, PZA^{+}17a, PZA^{+}17b, Wat18a, Wat18b, Bus10b]. Participation [CMY17]. Participatory [CPP16]. Partners [EJ12]. Partnership [KHD19]. Partnerships [CH15c]. Passing [Erd10d, Shu14e]. Past [Boo16f, Obr17]. Pat [Sal19]. Patch [SCB15]. Path [Dör18]. Pattern [HA11, HWBYZ13, TEvdHP12]. Pattern-Based [HA11, TEvdHP12]. Patterns [Bud13, LLC12, MM19, PAS10, Ris10a, ZPHW16]. Pay [Bus11d]. PC [Gim14]. PC-lint [Gim14]. Peaceful [FCS^{+}10]. Peaks [CHHSM13, MCH13]. Peer [Gau13, RCP^{+}12]. People [CCWP11, EP17b, GMB^{+}14, Hod13]. Perceive [HMPS16]. Perceptions [BSW19, SMAU16, SS18]. Perfect [Ano17k]. Perfectionists [Shu11c]. Performance [Ano18-37, Ano18-38, DS13, DFD^{+}16, GWA14, HKR11, LXC15, MRF^{+}13, OCFLI14, PSO12, RVT18, Sym10, TJC^{+}19, Sun19]. Performitis [Bus10a]. Perils [Ebe18]. Persistence [Boo14d]. Persisting [Via19]. Persona [CH13c]. Persona-Driven [CH13c]. Personae [CH14b]. Personality [MAGT19]. Personalized [BFKZ15]. Perspective [Ano17p, CCC^{+}19, EKP16, Kna19, KDM^{+}19, KCD12, MSSMDC14, RSS18, SSS15a, Spi15a, TVS10]. Perspectives [BBS13b, BB10, BKCF14, EM15, FCS^{+}10, PLTC16, PLC18, SWA17]. Pervasive [Ano16-40]. Peter [Her17]. Philae [BBP^{+}16]. Philaes [Bal18]. Phillips [And16]. Phone [TS12]. Phones [Bou10]. physical [BMSS18, RCLR18]. Pick [BE11a]. Pick-Up [BE11a]. Picture [Boo11d]. Piloting [BCN16]. Pipeline [GWLY11]. Pipelines [WZX^{+}16]. Pitfalls [GLA14]. Pivot [BWD^{+}17]. Plan [Whi12]. Planned [SNMF13]. Planning [FSB^{+}12, LIL^{+}13]. Platform [BSS^{+}17, CCS^{+}11, CNSM13, HCF^{+}15, KR10, LDO^{+}14, SHR12, TS12, Til15]. Platforms [GE11, JCP19, PBE14, ZMP^{+}16]. Play [Ris12c, MSAH17]. Playing [Ano13e]. Plug [KR10]. Plug-Ins [KR10]. Point [GE12, GWP^{+}11, HR11, KvdL10, VE12, WK11, WES10]. Point/Counterpoint [GWP^{+}11, HR11, KvdL10, VE12, WK11, WES10]. Points [HBS17a, HBS19, Hol15e, NCK11, PDV^{+}18]. Police [Shu12c]. Policy [CCMT16].

Politics [Mai12d]. Portability [Spi13d]. Positive [LA19]. Possibility [Boo13e]. Potion [DD10]. Power [AVF19, DeM13, Hod13, LBL17, PCP14, Woo16a]. Powering [WAA^{+}12]. Practical [GJO^{+}18]. Practice [ABB^{+}15a, BW14, DMG14, Gre18a, Gre18b, GGGM19, KDL17, KNO12, KDM^{+}19, PMSL14, PZA^{+}17a, PZA^{+}17b, RKTS12, WG16, WHR14, ZP11, ZHD^{+}13]. Practices [FWK^{+}15, GLA14, Jor19, MGS^{+}18a, MMYJ10, MCL12, MSSMDC14, Sca17, Spi18b, SF15, ZBCS16]. Practicing [Gha14]. Practitioner [PZTF12, PLC19]. Practitioners [FS10, LCA17, LTS12, MCBT13, PLTC16, PLC18, ST10, SS18]. Pragmatic [Fai19d, Kee15a, Woo14b, WF18]. Preach [Gha14]. Predict [WLD19]. Predictable [BBdH^{+}11]. Prediction [HBB^{+}11, LMP16]. Predictions [HKR11, Ker18b]. Predictive [Ano16-40, CSC^{+}18]. Premise [BKB12]. Prepare [Ano17-44, Ano17-42, Ano17-43, Ano18-36]. Preparing [BCB^{+}17]. Presentations [Ano15-32]. Preserving [SB16]. Pressure [Ano16-49]. Prevention [WMM18]. Pricing [BA10]. Prince [LMP16]. Principal [CSS12]. Principle [Fai19e]. Principled [CVEK13]. Principles [BW17, DFD^{+}16, SPA11, Woo16a]. Printing [Ano14z, Ano15y]. Prioritization [CH15a, SAO^{+}17]. Privacy [BFKZ15, Bre14, GK15b, HPM16, Pel12, Ano14m, Ano15o, Ano19i, Ano19j, Ano19-46, Ano19-47]. Privacy-Enhancing [Pel12]. PriView [BFKZ15]. Probabilistic [FN14, RCLR18]. Probing [CPP16]. Problem [Fai19b, Gla12]. Problems [MNS12, MTM^{+}19]. Process [CPP10, CCWP11, DeM13, GVE11, GFH18, GFPK10, Jun18, KC16, LOCPL16, MS14, MSSMDC14, PM10, Shu12c, SSS15b, UG18, vWHZ^{+}10, vGMRW14]. Processes [ABC^{+}11, BMDK15, Erd10a, HM13].

Processing [DFFP18, FDE^{+}17, WZX^{+}16, ZXL^{+}16]. Product [BCS10, BH19, Bos10, CdAC^{+}16, CSC^{+}18, Ebe14, EBS16, FSRdA11, GFT15, GFPK10, Joh10, JN10, KUG18, LK10, LX17, MNS12, MMYJ10, MRS10, dMSNRdCM^{+}11, Rei18, SNMF13, SdA13, SW11a, ZF13, dOdA16]. Product-Development [EBS16]. Productivity [Nic19]. Professional [ABH^{+}11, BvdHOP12, Boo12d, NYMS12, Som16, SMAU16]. Professionals [Spi16f, Uit13]. Profiles [GMB^{+}14, KHD19]. Program [RLBR12, SGSM10, ZKB^{+}16, ZK14]. Programmable [TM17]. Programmer [Nic19]. Programmers [Bla18b, Boo10d, Car19, PHAH16, Ris10c]. Programming [Ano10c, Ano14-34, CT13, CSS11, FTP11, Hof10, KV11, Mey14b, PB16, SEM14, SMAU16, Til16b, TW14b, Vol11, WC10, WCFG10, Wra10]. Programs [BH17, GS11, Hol17b, Hol17d, YLK^{+}17]. Progress [Amo18, Ebe18, Erd10e, Erd10f]. Progression [Shu14c]. Progressive [BBdH^{+}11]. Project [Ano17p, Fav10, Jor19, MCH15, MCL12, MS12, NCK11, SMP15, Sym12, VB16]. Projecting [VWK15]. Projects [Ano12a, BH19, Blu16, BSD16, CCSW10, DHB15, DeM11, FGBM14, MJZCH13, MFS15, MCBT13, STG19, ZSMP17]. Promised [TD19]. Properties [Roz15]. Proposals [FGBS10]. Protection [CT11, FCAJ11, Li14, Shu11d]. Proteus [BBB^{+}19]. Prototyping [CKH16, HCF^{+}15]. Providers [Fab10, Jor16]. Providing [GMB^{+}14]. Provisioning [BMR14]. Prowl [CH16b]. Psychology [TBvRB12]. Publications [Ano16-35]. Publishing [Cab18]. Purpose [PTNF^{+}19]. Pursuit [VBP17]. Putting [Haz10, Off13]. Python [RO15].

Qualitative [BHG13]. Qualities [HSEG10]. Quality [BKCF14, Bus10d, BAA^{+}12, CYM^{+}15, CPP16, Car19, DHB15, FV19, GF16, HMPS16, HBCH14, HL11, KHYT18, LMG16, LSCK12, PS14, RLBR12, SM10a, Spi16e, SSS15b, ZK10]. Quantifying [SJS12]. Quantitative [TD12]. Quest [Gre19]. Question [Boo14f, SB16]. Questions [Boo14a, CPP16, FS10, Hol16c]. Quoins [CH13d]. Quotes [Obr18].

R [LA19]. Race [KC16]. Racing [CMS11]. Radio [Blu19]. Railway [FFGM13]. Randomly [Hol17c]. Range [Tor15]. Rapid [SCB15]. Rare [AGIF17]. Rate [vGH12]. Rating [RNA^{+}16b]. Ratings [KNH16, RNA^{+}14]. Rationales [TSPB17]. Raw [MCH15]. Raynaud [Car17a]. RE [DHH^{+}19, Gre17a]. Reach [Spi15b]. Reactive [DSM^{+}18, Til16b]. Read [Ano16-35, SCC11]. Ready [BWD^{+}17]. Real [LDO^{+}14]. Real-Time [LDO^{+}14]. Reality [LR12, PZA^{+}17a, Pra17]. Really [ASP13, Ano10c, Erd10c, KT18, Wra10]. Reasoner [VD15]. Rebecca [Ano10b]. Recognition [Ano16v, DPCL14]. Recognizing [Ano18-37, Ano18-38]. Recollections [EM18]. Recommendation [RWZ10]. Recommendations [CC19]. Recommender [CPB^{+}17]. Reconsidered [Bus13]. Reconsidering [Nag18]. Recruiting [Spi15e]. Redirections [PM18]. Reducing [AKN^{+}16]. Reduction [UM10]. Redux [Spi17e]. Reengineering [Bus11b, CPS18, PCdGPE11]. Refactor [Fai19c]. Refactoring [BS15, Bus11a, Dig11, Dig15, GRD11, GO15, HO15, LML^{+}15, MHRSO15, PCP14, Sdd^{+}18, SSS15a, SGSM10, Spi12e, Via15, Zim15]. Reference [PvdHM15, WE16]. Reflect [CH16a]. Reflecting [BHS^{+}14, Spi16e]. Reflection [BHN14, NYMS12]. Reflective [BW14, DMG14, PMSL14]. Regress [Erd10e]. Regression [CCMT16, Shu14c, SAO^{+}17]. Reins [Erd10d]. Relational [Via19]. Relationship [BBB^{+}18, KNH16]. Relationships [Jor19]. Relative [Jor13, KE10]. Release [ABB^{+}15a, MM19, RR15, Spi15f, Spi18b, ABB^{+}18]. Releases [MFS15, SCB15, ZXT^{+}15]. Reliability [KCK^{+}13, LBWX17, Spi17e]. Reliable [Hol15a, ZXT^{+}15, Hol14]. Remembering [Ano11b]. Remembrance [Boo16f]. Remote [CT11]. Renewing [Fav10]. Repair [SEK^{+}19]. Repairing [LW11]. Replacement [CT11]. Replicating [AGIF17]. Report [BDD^{+}18, EV10, GTS^{+}19, IHSR19, PT18]. Reports [NHM19, ZCLL19]. Repositories [DDIP15]. Repository [APPV19]. Repository-Mining [APPV19]. Reproducibility [Hv19]. Reputation [ORM^{+}17]. ReqIF [EJ12]. Requirement [PRRT14]. Requirements [AACF13, dSAPW^{+}14, BBC^{+}16, BB17, Bre14, Car17b, CMP17, CBN13, CH13c, CH13b, CH13d, CHHSM13, CH14a, CHL14, CH15b, CH16a, Coo10, DFFP18, DHB15, DeM13, DHH^{+}19, EJ12, FC19, FB15, FV19, FWK^{+}15, Fer18, FDE^{+}17, FGBS10, GM11, GCH13, Gre17b, Gre18a, Gre18b, Gre19, GSA^{+}17, Kna19, LCH17, MNJR16, MBLN10, MJ10, Mai11b, MW11, Mai12b, Mai12c, Mai12e, Mai13a, Mai13b, MG17, NBF^{+}18, Pen15, RSH12, SM10a, Shu13a, SHR12, SM10b, TD19, TSPB17, TEvdHP12, VRG16, VGD19, WCG16, dGNA^{+}11]. Requirements-Driven [MBLN10]. Requirements-Engineering [Fer18, Gre18a, TD19]. Research [BBB^{+}18, BBN^{+}17, DFD^{+}16, Fer18, FS10, GJO^{+}18, GO15, HBB^{+}11, LMP16, MF13, SPA11, SCL18, She15, Shu12f, TD19, VD10]. Researchers [MM18a]. Resilience [Sal19]. Resource [KSW10, NPJ19]. Resources [Ano16w, Ano17y, PZ18, Shu11c]. Responders [BBG19]. Responding [MSAH17]. Response [PGG^{+}18, WMM18]. Responses [Ano10c, LA19, NHM19]. Responsibility [BWC10, CH16a, Spi17c]. Responsibility-Driven [BWC10]. Responsiveness [OSBA14]. Resting [PZ18]. Result [She15]. Results [DFD^{+}16, Jor19]. Retrofit [Ron16]. Retrospective [GO15, Kri17, MM13, MCBT13, Spi18c]. Retrospectives [Ter11]. Return [Woo14b]. Reusable [LWCP17, Zim11]. Reuse [KLM11, MT19, MAN^{+}14, SJ13]. Reveals [Fai19b]. Reverse [SSE12, VD10]. Reverse-Engineering [SSE12]. Review [Gau13, RCP^{+}12, TD19]. Reviewers [Ano10a, Ano13a, Ano14a, Ano15a, Ano16a, Ano17a, Tay18]. Reviewing [MGS^{+}18a]. Reviews [Boo10c, CCH^{+}15, Fai19a, HA11, MSAH17, vHEA^{+}14]. Revisited [Bal18, BAA^{+}12, KCD12]. Revisiting [JWCR18]. Rework [Mor10]. Rewrite [Fai19c]. Rewriting [BJ11, Bus11b]. Rich [KR10]. Richard [Ano15b, Ano16b]. Right [CS16, Hol17c, Ano16c, Ano17e, Ano17d]. Rights [GD11, LL16]. Rigorous [BBB^{+}11]. Rise [EV10, HSG18]. Risk [Ano16-41, BA10, BMSS18, FN14, LCH17, Poo14, PS14, RCLR18, Wat18a, Wat18b]. Risk-Based [Ano16-41, PS14]. Risks [ES16, GM10, PM10]. Road [KAC^{+}18, LMM^{+}15, SA11, Shu11e]. Roadmap [TM17]. Roads [LMP^{+}15]. RoboPlatform [LTv19]. Robot [PB16]. Robotic [MHJ17, VOGB18]. Robotics [AGK^{+}19]. Robots [LTv19]. Robust [XYCW18]. Rock [Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano14z, Ano15z, Ano15-27, Ano15y, Ano15-28, Ano15-29, Ano15-31, Ano16x, Ano16-37, Ano16-36, Ano16-39, Ano16-38, Ano16-40, Ano16-41]. Role [Ano16-42, EP16, HOZZ16, Sha12, Spi16b, TKF16, WG16]. Roles [LKR16]. Ron [Bla18b]. Room [Hv19]. Roundtable [ABB^{+}15a, HHR^{+}13, HSF^{+}16, LRB^{+}15]. Routes [CPB^{+}17]. Routines [CSS11]. Rover [Ano13d]. Ruby [Gun10]. Rules [MZLD12]. Run [MTT^{+}15, Stu18]. Run-Time [MTT^{+}15]. Rundeck [Spi14d]. Runtime [BBB^{+}19, LW11, MSS10, OCFLI14].

SaaS [BKB12]. SaaS-Based [BKB12]. Safe [MOK^{+}18, SRI14]. Safer [SGSM10]. Safety [dSAPW^{+}14, BMSS18, CCP^{+}18, CH17, Ebe15a, FSB^{+}12, KSTS17, LCF13, MJZCH13, MG17, PRRT14]. Safety-Critical [CCP^{+}18, KSTS17, LCF13, MJZCH13, MG17]. Salary [Bla17]. Same [Ano12a, BB17, DeM11]. Sanity [BvD10]. SASSY [MGMS11]. Satellite [CMC^{+}12]. SATURN [ZK14, Kee15b]. Savings [FC19, SCW15]. Say [LTS12]. Scalability [DLR10]. Scalable [DSM^{+}18, VAC^{+}17]. Scalable-Application [VAC^{+}17]. Scale [BW17, BH19, BSD16, CC19, DFP19, Fai19f, Kna19, MAN^{+}14, NH15, OSBA14, PL19, SCW^{+}18, SMLF19, Ozk19d]. Scaling [EP17a, Kim19]. Scanners [HvdL10]. SCEPYLT [CH13a]. Scheduler [RTV12]. Science [BJ11, EEBB11, EHMFT19, Gla12, MIC19]. Scientific [KTH11, MS12]. Scoping [Joh10]. Scorecard [Spi13b]. Scrapheap [KLM11]. Screen [BDD^{+}14, GKD^{+}15]. Screen-Based [BDD^{+}14]. Scripting [SRI14]. Scripts [GF16]. Scrum [SJS12]. ScyllaDB [Sun19]. Seamless [EJ12]. Search [Ano14d, Ano14c, Ano16g, Ano16f, Ano17q, Fav12, GV11, Ano15g]. Searching [And12, Joh13]. Secret [Mar12]. Secure [Ano17-39, DBRV18, LDO^{+}14, MOK^{+}18, PST^{+}17, WCG16]. Securing [SA19]. Security [Amo18, Ano14m, Ano15o, Ano16-38, Ano16-41, Ano19i, BMSS18, CM19, GP13, GK15b, HPM16, McG11, PRRT14, SS17, WMM18, Ano19-46, Ano19-47]. See [HBS17b]. Seeing [Erd10e]. Seeking [PZ15]. SEI [ZK14]. Selection [Jor16]. Self [BBB^{+}19, GFT15, GGGM19, MGMS11, Spi18d, WG10]. Self-Adaptation [WG10]. Self-Adapting [GFT15]. Self-Adaptive [BBB^{+}19]. Self-Architecting [MGMS11]. Self-Evolving [Spi18d]. Self-Managing [GGGM19]. Semi [SEK^{+}19]. Semi-Autonomic [SEK^{+}19]. Sense [Mey18]. Sensor [ESZ15, MKD14]. Sensors [MKD14]. Sentiment [CCP^{+}18, MHM19, NS19, WLD19]. Separation [GHV10]. Sequential [Erd10a]. Seriously [Kie10]. Serverless [Bha18]. Service [And16, Ano15b, ABC^{+}11, AIT15, Fab10, GD11, GMB^{+}14, LK10, LMSS11, LR12, MBLN10, Mai10a, MGMS11, MH14, PZA^{+}17a, PZA^{+}17b, RMM^{+}12, See12, SHG14, SGH15, Spi14d, Ano16b]. Service-Based [ABC^{+}11, GD11]. Service-Orientation [LK10]. Service-Oriented [LR12, MBLN10, MGMS11, MH14, SHG14]. Services [Ano19-54, CSS11, FMMP16, HCF^{+}15, PAB11]. Serving [Spi16f]. Seymour [Ano18-37, Ano18-38]. Shared [Boo10b]. Shareholder [Con12]. Sharing [Bre14, CMC^{+}12, LL16, Rie11, Shu13f, SRI14, SMLF19]. Shepherding [TKF16]. Shift [KC16, VGD19]. Shipping [vdBB17]. Shirt [Spi13d]. Short [FS12, SA11]. Shot [LML^{+}15]. Should [HR11, MFS15, Ris10c, Woo17]. Show [HBB^{+}11]. Side [Bri12]. Sidney [Ano18-37, Ano18-38]. Sigelman [Blu19]. Sigil [DeM14]. Significant [CBN13, CH13c, CH13d]. Silver [Mar12, Ozk19a]. Similar [LC11]. Simple [MPD14]. Simply [Gla12]. Simulation [ADH14, GTS^{+}19]. Simulation-Based [ADH14]. Sin [HR19a]. Single [Gat16]. Sinha [Rei18]. Size [ES14, Jor19, NCK11]. Skepticism [RSAT19]. Sketch [CM12]. Sketchifying [Obr13]. Skill [Kri17]. Skills [Car19, IML16]. Skirted [Pra17]. Small [Boo11b, EBS16, FSM14, LMMO18, LOCPL16, RVB13]. Smaller [KE10]. Smart [BCN16, CPB^{+}17, LLV16, MTT^{+}15, PvdHM15, SWB^{+}17, Spi17a]. Smarter [Ano18g]. Smartphones [HLS17]. Smells [GKF19, TL18]. SOA [GLA14]. Social [Ano19-34, BBS13a, BBS13b, CH16a, DHH^{+}19, Li14, MTM^{+}19, MCD19, PSGd18, Shu13d, Spi17c, TLvV13, TBvRB12, ZWGY12, CSS13]. Society [Ano17x, Ano14l, Ano16t, Ano16u, Ano16x, Ano16p, Ano16v, Ano16r, Ano16q, Ano16s, Ano16w, Ano16o, Ano17y, Ano18p, Ano18q, Ano18t, Ano18r, Ano18s, Ano18u, Ano19y, Ano19z, Ano19-27, Ano19-28, Ano19a, Ano19b, Ano19-29, Ano19-30]. Sociological [GMB^{+}14]. Software [ASP13, AtHR11, AS17, AGK^{+}19, AGIF17, ABB^{+}15b, AVF19, Amo18, And15, Ano14-32, Ano14-33, Ano14-34, Ano15n, Ano15-30, Ano16-42, Ano16-43, Ano17c, Ano17m, Ano17p, Ano17b, Ano17z, Ano18c, Ano18d, Ano19j, AMK16, ABH^{+}11, AHMS18, APPV19, AKN^{+}16, ACHC11, BCS10, BHN14, BvdHOP12, BBP^{+}16, BZd18, BBB^{+}19, BCLP15, BJ11, BBB^{+}18, Bas18, BCD^{+}17, BCB^{+}17, BBS13a, BBS13b, BDH^{+}19, BA10, BMR14, BMM14, BBM^{+}15, BBG19, BMSS18, Bla17, BSW19, Blu19, BNW10, Boo14c, Boo18, Bos10, BCH15, Bos16, Bou10, BKCF14, Bri12, BBN^{+}17, BSD16, Bro19, Bro18, BMSF15, BW14, BHH13, BH10a, BH10b, Bus11e, Bus11f, CE19, CR14, CMM16, Cap14, CYM^{+}15, CSC^{+}18, CP18, CPS18, CS18, CS19, CM19, CT11, CCSW10, Cle10, CR11, CMC^{+}12, CSS11, CNSM13, DeM14, DZB18, DDIP15, DFD^{+}16, DvGvS14]. Software [DMG14, EEBB11, EAO12, ES14, Ebe14, EHS15, EKP16, EF17, EC17, Ebe18, EHMFT19, EBS16, ES16, EP17b, Erd10b, Erd10e, EM18, EMP18, EFO14, FCAJ11, FLC17, FCS^{+}10, FWS19, Fav10, FP11, FN14, FKL19, FRM15, FP13, FCN16, FSRdA11, GFT15, GD11, GRDL^{+}12, GF17, GFH18, GFKY18, GTS^{+}19, GKF19, GMP14, GUP^{+}14, Gim14, GBICMR13, GFPK10, Gor13, GK15a, GBM16, GM10, GWA14, GO15, GKKL12, HMPS16, HOBK17, HKR11, HHR^{+}13, HvG16, HPM16, Haz10, HKB16, HSF^{+}16, HJP16, HL11, HSG18, HvdL10, HOZZ16, Hol19a, Hol18c, HV16, IGBKR19, JCP19, Jep13, Joh15, JEJ12, Joh13, JWCR18, JN10, Jor13, Jor14, Jor16, Jor19, Kaa14, KDL17, Kee15b, KTH11, Ker18b, Ker18d, KAC^{+}18, KCK^{+}13, KUG18, KUG19, KH18, KLM11, KHD19, KDM^{+}19, LvdH16, LAH10, LC11]. Software [LEPV10, LCE13, LMP16, LMMO18, LCF13, LOCPL16, LCFT17, LWCP17, LSZ18, LIL^{+}13, LBL17, LSCK12, LDBF19, LTS12, LRP11, LRB^{+}15, MNS12, MF13, MTM^{+}19, MM18a, MS14, McG11, MMYJ10, MCL12, MGS18b, MAGT19, dOMFP16, MGFRD10, MCD19, MZ13a, MZ13b, MZ18, Men18, MZLD12, MDOS11, MT19, MCBT13, MRS10, MAN^{+}14, MSSMDC14, MHF17, Mor10, MS12, Mos10, MBv18, NYMS12, NPJ19, dMSNRdCM^{+}11, NOS12, NS19, NS17, NCK11, Obr13, ORM^{+}17, Off13, OSBA14, Ozk19b, Ozk19c, PHAH16, Par18, PZ18, PTNF^{+}19, Pec13, PBE14, PDV^{+}18, PWF^{+}19, PGG^{+}18, PST^{+}17, Pop11, PC12, PT18, PS14, PAS10, PPMD17, RVB13, RWZ10, RSAT19, RPE19, Ron16, RBE19, Rou12, RDMA11, RSH13, Rut16, RvGH17, Sdd^{+}18, SM10a, Sca17, SA19, SA11, See12, Shu11a, Shu11b, SCC^{+}16, SMS^{+}14, SMP15, SW11a, SW11b]. Software [SCW15, SV16, SWB^{+}17, SHR12, SCB15, Spi10c, Spi12b, Spi13e, SAT14, Spi16b, Spi16d, Spi17b, Spi17f, Spi17d, Spi17e, Spi17c, Spi18a, Spi18d, SLB17, SMAU16, SSS15b, Sym10, Sym12, TM17, TLvV13, TvV12, TYXL16, TBvRB12, TVS10, Ter11, TI11, Tor15, Tsa11, Uit13, Val17, VSB^{+}16, VM12, VBP17, VB16, VMvG12, VMVS19, WKS^{+}14, WK15, Whi19, Woo16c, YFMH18, Zam17, ZMP^{+}16, ZSMP17, ZBOC12, ZK10, ZHD^{+}13, ZCLL19, Zim15, dSMdSC16, dMHSC16, dOdA16, dBLMT11, vGH10, vGH11, vGH12, vGH15, vWHZ^{+}10, Ano14n, Obr17, Som16, Ano19o, Obr18, Spi18f]. Software-Crowdsourcing [ZSMP17]. Software-Defined [PGG^{+}18]. Software-Development [FRM15]. Software-Driven [EHS15]. Software-Engineering [Spi17d]. Software-Intensive [GFPK10, KUG18]. Solicited [Ano18-37, Ano18-38]. Solution [Fai19b, Gat16, Kri17]. Solutions [BBM18, BKB12, MNS12, MHLCL16, SSS15a, SWB^{+}17]. Solving [MTM^{+}19]. SORE [Ale11]. Sorrow [Boo12f]. Soul [Boo11e]. Source [AtHR11, ACHC11, Bro19, CCSW10, Eng10, FGBM14, FFZ^{+}18, GD12, HR11, LRP11, MTM^{+}19, MFS15, MOMM11, Rie11, RCP^{+}12, RSAT19, Sca19, Spi11b, STG19, SF15, WKS^{+}14, dBLMT11]. Space [KSTS17, NBS^{+}16, SS11]. Spacecraft [Hol13]. Spaces [DvGvS14, Sha12]. Spam [HJ17]. Specific [Fre10]. Specification [SSE12]. Specifications [ASK14]. Speech [DPCL14]. Speed [Bos16]. SPI [PDNdSB12]. SPLASH [MF13]. Spocks [Mai12e]. Spoken [CCMT16]. Spoofax [WKV14]. Spot [Ano14-35]. Spotify [SMLF19]. SQALE [LI12]. Squared [GV11]. Stabilization [RR15]. Stack [ASR17, BBS13b, SB16]. Stacks [Lou16]. Staff [Ano19-42, Ano19-43, Ano19-44, Ano19-45]. Stage [DD12]. Staged [SNMF13]. Stakeholder [LAH10, TVS10]. Stakeholders [CH16b, SR16]. Standards [Ano17i, Ano18f, FSB^{+}12, LMMO18, LDBF19]. Star [Spi15e]. Stars [Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano14z, Ano15z, Ano15-27, Ano15y, Ano15-28, Ano15-29, Ano15-31, Ano16x, Ano16-37, Ano16-36, Ano16-39, Ano16-38, Ano16-40, Ano16-41]. Start [BWD^{+}17, KUG18, KUG19]. Start-Ups [BWD^{+}17, KUG18, KUG19]. Startup [Ano15-31]. Startups [GUP^{+}14, LMMO18]. Stasis [Shu14c]. State [BCLP15, HOBK17, KDL17, KSTS17, Lob11, Spi17f, WHR14]. State-of-the-Art [Spi17f]. State-Space [KSTS17]. Statefulness [FFZ^{+}18]. Static [PTV^{+}10]. Statistical [KCK^{+}13]. Statistics [LE13]. Status [GE11]. Stay [Ano16-44]. STC [Ano15-32]. Steering [CMM16]. Steps [PST^{+}17]. Sticks [BE11a]. Stigmergy [ZZJJ15]. Stigmergy-Based [ZZJJ15]. Still [JWCR18]. Stop [LW11]. Storage [Gho10]. Store [GAM^{+}17]. Stores [RNA^{+}16b]. Stories [Boo12e, Boo13e, Boo14e, CH17, Ris10b, Ris14]. Story [HBS17a, HBS19, Shu13f]. Storyboarding [SM10b]. Storytelling [Uit13]. Strategic [BNW10, CKH16, CH13d, MJZCH13, PT18, PLC19, Spi15f]. Strategies [Ano15-36, Ano16-50, DD12, LDBF19, SW11a, Ano17-52]. Strategy [TvV12]. Stream [Ker17, Ker18d]. Streamlining [HHLJ10]. Streams [NYMS12]. Streetlight [Joh13]. Strict [LCH17]. Strokes [Via18]. Structures [PvdHM15, Par18]. Studio [BW14]. Studio-Based [BW14]. Study [ASR17, BCLP15, BSD16, BMSF15, GGGM19, Kee15a, LX17, McG11, MCBT13, MAN^{+}14, NHM19, OT11, RVT18, SJS12, ZF13]. Studying [BvdHOP12, MSAH17]. Styles [Bus10c]. Subscribe [Ano19g, Ano19j, Ano19-46, Ano19-47]. Substitution [Rut16]. Succeed [Bro19, SMLF19]. Success [ASP13, JN10, Lob11, PDNdSB12, Spi18e, WAA^{+}12]. Successful [Jor19, Kle16, MMYJ10, Roz15, STG19]. Successfully [KHD19]. Suisse [SJG13, MH14]. Summaries [BH17]. Summit [Ano14-33, Ano15-30]. Super [LLV16]. Super-ility [LLV16]. Support [BBB^{+}19, FN14, Gat16, PMCE12, WLD19, WHR10]. Supported [FSB^{+}12]. Supporting [BW14, Fer18, LWCP17]. Surprise [Ano17n]. Survey [RVT18, UM10, dBLMT11]. Sustainability [ASH13, BBC^{+}16, CCC^{+}19, CNZC17, CPS18, KDGV13, PZA^{+}17b, PRRT14, Pen15, PDV^{+}18, SNMF13, SB16]. Sustainable [Ano16y, Ano18e, Ano19m, Ano19n, LIL^{+}13, ZCTZ13]. Swarms [PB16, VOGB18]. Sweet [Ano14-35]. Switzerland [SJG13]. Symbiotic [BBB^{+}18]. Symbolic [YLK^{+}17]. Synergetic [LNH^{+}11]. Synthesis [AIT15]. System [ASP13, BBB^{+}11, Bus10d, Cab18, CH13a, CNZC17, Fre10, HSEG10, Kna19, Lob11, MCHK17, RNA^{+}16b, Sdd^{+}18, SAO^{+}17, TI11, WGC^{+}13, vdBB17]. System-Level [SAO^{+}17]. Systematic [SSE12]. Systems [Ave11, BMSS18, Boo10f, BCH15, BO17, CCC^{+}19, CGE19, CFF^{+}13, CCP^{+}18, CKH16, CCD^{+}17, ED19, EW19, EFO14, GFT15, Gor13, GK15a, GBM16, Hol19a, HHLJ10, Joh16, KR10, KSTS17, LC11, LDO^{+}14, LMSS11, LSCK12, LCYL14, LX17, MG17, MM18b, MGMS11, MDOS11, MHLCL16, Mos10, NH15, NPJ19, Pec13, PGG^{+}18, RCLR18, RRL^{+}17, RWZ10, Ron16, Sal19, SA11, SHG14, Spi12d, Spi13e, Tou11, VRG16, VMVS19, WG10, Ano19h].

T [Sma18]. Table [Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano14-40, Ano15-33, Ano15-34, Ano16-45, Ano16-46, Ano16-47, Ano16-48, Ano17-45, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano18-39, Ano18-40, Ano18-41, Ano18-42, Ano18-43, Ano18-44, Ano19-48, Ano19-49, Ano19-50, Ano19-51, Ano19-52, Ano19-53]. Taggart [Bha18]. Tail [LXB^{+}15]. Tail-Tolerant [LXB^{+}15]. Tailored [MM18b]. Tailoring [BH19]. Take [Ano15-36, Ano16-50, Kie10, Ano17-52]. Takeover [VMvG12]. Taking [LTv19]. Tale [Hol17d]. Tame [MHF17]. Tammy [Sal18]. Tangible [BDD^{+}14]. Tanzanian [SJ15]. Target [Ano14y]. Task [KCD12, TBvRB12, Zim15]. Task-Based [KCD12]. Task-Centric [Zim15]. Task-Effective [TBvRB12]. Taxes [DLR10]. Taxonomies [FB15, GV11]. Taxonomy [KUG18, TM18]. TCSE [Ano17g]. Teach [BCD^{+}17]. Team [AHMS18, DFD^{+}16, Fai19f, Kee18, Spi15e, VB16, XYCW18, dSMdSC16]. Teams [BSW19, CE19, CMY17, DvGvS14, FRM15, GGGM19, HBN13, MCL12, PM10, PPMD17, RL16, REV10, RH13, SKK14, TBvRB12, Ter11, Woo15a]. Tech [Ano18v, Ano19-35]. TechIgnite [Ano17-51]. Technical [Bav12, Boo10e, Bus11d, CSS13, CCCM17, Con12, CSS12, KNO12, LI12, LTS12, MTM^{+}19, RKW15, SWA17, WJ15]. Techniques [APPV19, BE11b, HM13, MPD14, SCL18, UM10, VM12]. Technologies [ES14, EHMFT19, PCdGPE11, PMCE12, REV10, RPE19, Sca17, SGH15, ZWGY12, Ano17j]. Technology [Bus10b, CE19, CMP17, DDIP15, EC17, McG11, MM13, Pel12]. Telematics [HCF^{+}15]. Tell [Boo14e, Ham18, vGMRW14]. Telling [Ris10b]. Temporary [PPMD17]. Ten [MW19, vGH19]. Term [SNMF13]. Terrible [Boo13c]. Terrorism [CH13a]. Test [Ano16-32, ADSS18, GF16, GE17, GFH18, GKF19, Gue14, IML16, KT18, Lou11, NH15, PRPE13, SMT^{+}10, SAO^{+}17, Whi12, WHR10, YLK^{+}17]. Test-Driven [Gue14, KT18, SMT^{+}10]. Testing [AGIF17, Als13, AFT^{+}15, BBdH^{+}11, Bav12, BB17, CCMT16, CR11, DS13, FB15, GF17, GFKY18, GTS^{+}19, Hol15a, HLS17, KDL17, KTH11, LBL17, LCYL14, MHJ17, MZLD12, MLD^{+}13, dMSNRdCM^{+}11, RPFA^{+}14, RKTS12, Sch12, Shu12a, Spi17f, SS17, UM10, ZML^{+}12]. Tests [Bus11c, Til16a]. Texts [FS12]. Their [DZB18]. Them [GV11]. Theory [GTS10, JEJ12, KE10, KNO12]. There [BWN16, MPD14]. Thing [CS16]. Things [Boo11b, Boo12a, Boo16f, Ano16-43, BO17, ESZ15, FMMP16, LCFT17, Spi17d, WE16]. Think [BDH^{+}19]. Thinking [CH13d, CH15a, HBCH14, Lat12, Shu11d]. Thomas [Joh16]. Thought [MONF16, NYMS12]. Threading [KV11]. Threat [RCLR18]. Three [ABB^{+}15a, Boo12e, Hol17d, SPB16]. Throughput [LNH^{+}11]. Thyself [Spi16c]. Ticket [WLD19]. Ticking [Shu14a]. Time [Ano16-49, Boo16e, Gre19, Kie10, LDO^{+}14, Mai11c, MFS15, MTT^{+}15, Poo16, PMCE12]. Time-Based [MFS15]. Timelines [BHS^{+}14]. Tiny [Hol16e]. Today [Bro18, PZTF12, Ano17g]. Together [BL10, BB17, RI12]. Tolerant [LXB^{+}15]. Tomorrow [BCB^{+}17, Bro18, HSF^{+}16]. Tone [LA19]. Too [Che15, Wri11]. Tool [CM10, FSB^{+}12, FRM15, Gat16, WHR10]. Tool-Supported [FSB^{+}12]. Toolchains [Ker18c]. Tooling [SAT14]. Tools [BS15, BE11b, GVE11, HGS15, HSEG10, Hol16e, Ker18a, LEPV10, MF13, Mey14a, REV10, RPE19, SAT14, TVS10, VM12, dGNA^{+}11]. Top [FS10, GCH13, MM18a, MSAH17, PHA^{+}17]. TopCoder [BBS13b]. Topics [MM18a]. Traceability [CH13d, LX17, MJZCH13]. Tracery [GM11]. Tracing [Blu19, Tou11]. Tracking [Erd10f, LSCK12, SSWM16, vdBB17]. Tracks [Spi10c]. Trade [SM10a, Sym12]. Trade-Offs [SM10a, Sym12]. Traffic [CFF^{+}13, TI11]. Tragedy [LMP16]. Train [TI11]. Transactions [Ano16y, Ano18e, Ano19k, Ano19l, Ano19m, Ano19n, Ano19p, Ano19q]. Transfer [McG11]. Transfers [SW11a]. Transform [Sma18]. Transformation [ED18, Sma18, SC18]. Translation [DPCL14]. Transparency [Boo11f, DSTH13]. Transparent [HKB16]. Traversing [MCH13]. Trees [Spi15c]. Trenches [MGS^{+}18a, Via15]. Trends [BCH15, ES17, GE11, Kee15b, MM18a, MSS10, PLTC16, PLC18, PLC19, TMB17, ZWGY12]. Trialsourcing [Jor16]. True [Shu14f, SvS16]. Truly [MHJ17]. Trust [BS15, Mai10b, MCL12, She15, SCW^{+}18]. Trusted [CCS^{+}11]. Trustworthiness [dBLMT11]. Trustworthy [BS15, MDOS11]. Truth [Hol15e, Ker18c]. Tug [SSS15b]. Tutorial [PC12, SF15]. Twenty [HWBYZ13, RSAT19]. Twin [CHHSM13, MCH13]. Twitter [WKS^{+}14]. Two [GRDL^{+}12, GV11, LA19]. Two-Layer [GRDL^{+}12]. Type [EP17b]. Types [MAGT19].

U.S. [Sca19]. Ubiquitous [VGD19]. Ugly [BZd18, vGH19]. UML [CCMT16, LLC12, Spi10d, Woo14a]. Unanswered [Hol16c]. Unbalanced [Boo11f]. Uncertainty [HBS19, KC16, TJC^{+}19, Wat18a, Wat18b]. Uncovering [TLvV13]. Underground [vMM12]. Undersea [PWF^{+}19]. Understand [Mor10, VMVS19]. Understanding [FGBS10, GBICMR13, HBCH14, TVS10]. Undiscovered [LLV16]. Unified [CPP10]. Unintentional [Boo11f]. Unit [ADSS18, Til16a]. Universe [Mai12e, ZB16]. University [Haz10]. Unlimited [Ano17-41, Ano17-40, Ano18-34, Ano18-35]. Unmanned [PWF^{+}19]. Unmasking [GM10]. Unplanned [Gre17b]. Unreasonable [Men18]. Unseen [Boo12a]. Unusable [Bus11e, Bus11f]. Unweaving [NYMS12]. Updated [PLC18]. Updates [RNA^{+}16a, VOGB18]. Upgrade [vGMRW14]. upon [Boo16e]. Ups [BWD^{+}17, KUG18, KUG19]. Uptime [Ron16]. Usability [GRD11, LL16]. Use [ASR17, CdAC^{+}16, CVEK13, FV19, Kri17, Mav12, NCK11, SS18]. Use-Case [NCK11]. Used [RNA^{+}16b]. Useful [Joh13]. Useless [Bus11e, Bus11f]. USENIX [Ano15-35]. User [Ano14-35, BZd18, BDD^{+}14, MM19, MM13, SMN19, Shu12b, ZHG14]. Users [ASP13, KSNH15, LA19, MRF^{+}13]. Using [ASR17, Als13, BBB^{+}11, CT11, CPP10, FC19, FB15, FN14, HKR11, HL11, HHLJ10, Joh10, KH18, KSTS17, MHM19, MTM^{+}19, MOMM11, Mor10, PSGd18, SSE12, SJS12, SC18, Spi11b, Tou11, Via19, WG10, YLK^{+}17]. Usual [Shu11b].

Validating [KCK^{+}13, Shu11a]. Validation [EW19, MDOS11, RPE19]. Value [Bus10d, CH15a, Erd10f, Gat16, HBS17b, Hol17e, Ker17, Ker18d, PS14, VE12, VMVS19]. Value- [PS14]. Value-Focused [Bus10d]. Value-Thinking [CH15a]. Valueless [Whi19]. Values [CMP17, Shu13a]. Variability [Ano17m, BCS10, BCH15, GLA14, MCHK17, MTT^{+}15, RDMA11, TAJB17]. Variability-Intensive [GLA14]. Variants [LLC12]. Vaughn [Til16b]. Vault [Hol18a]. Vehicle [PWF^{+}19]. Vehicles [FCN16, WCG16]. Vendor [RMM^{+}12]. Vendor-Neutral [RMM^{+}12]. Verification [Ano17c, CR11, MDOS11, MLD^{+}13, RPE19]. Verifying [GF16]. Vernon [Til16b]. versus [DFD^{+}16, SJS12, SSS15b, Sym12]. Very [LMMO18, LOCPL16]. Video [ZXL^{+}16]. View [Ano14-41, Woo14a, Woo16b, Zim15]. Viewpoint [CB10]. Views [DBRV18, SM10a]. Violations [MOMM11]. Virtual [DvGvS14, HLS17, MKD14, SKK14, Ter11, ZB16]. Virtualize [Spi12f]. Visible [NOS12]. Vision [FDE^{+}17]. Visions [Shu12c]. Visual [AGIF17, GP13, TVS10]. Vivo [VD10]. Voice [DPCL14, Ozk19f, PM18, SDSP17, SMP15]. Voice-Based [DPCL14]. Voting [KV18]. Vroom [BMDK15]. vs [Spi13d, VE12]. Vu [Erd10b]. Vulnerability [WMM18].

War [SSS15b]. Warnings [KNH16].

Warshall [RTV12]. Watch

[Ano15-36, Ano16-50, Ano17-52, MM19]. Watched [Boo15a]. Waterfall [Ano17n]. Watson [Boo11e]. Watts [Shu11f]. Way [Ano16-35, CVEK13, JN10, Kri17]. Weakest [Hol16f]. Wearables [Ano15-28, Ano15-29]. Web [ABB^{+}15b, CSS13, DS13, FMMP16, GRD11, GK15b, KZW15, Osb11, PZ18, RPP19, ST10, SHG13, SS17, VD15]. Web-Based [ABB^{+}15b, KZW15]. WebDSL [GHV10]. Week [Bus12b]. Welcome [Boo15c]. Well [CH14b]. We’re [BL10, CH14a]. Where [Ano16w, Ano17y, Bus12a, Con12, JEJ12, MONF16, SCW15, CH14a]. Whether [Nag18, She15]. While [Bro19]. Whisper [SW11b]. Whistles [DeM13]. Whiteboard [RI12]. Whittaker [Shu12a]. Who [APPV19]. Wikipedia [FS12]. Wild [RPP19]. Winding [SA11]. Wireless [ESZ15]. Wirfs [Ano10b]. Wirfs-Brock [Ano10b]. Wishful [Shu11d]. Wonder [Boo13f]. Won’t [BWN16]. Word [Ris14]. WordPress [Cab18]. Work [BCB^{+}17, Mai12c, Mai13b, Rut16, Woo15a, Woo15b, Ozk19d]. Workbench [WKV14]. Workflow [AtHR11, BDD^{+}14, CMM16]. Workflows [MTT^{+}15]. Working [BB17, RI12, Til16a]. Works [Ano10c, Wra10]. Workshops [RH13]. World [Ano15-36, Ano16-50, Ano18g, CHL14, EP16, Ris12a, Shu11c, Shu12a, Shu12b, TM17, Woo16c, Ano17j, Ano17-52, Ano19-54]. Worlds [GF17]. Worrying [LW11]. Worth [HR11, MSAH17, vGMRW14]. Worthwhile [RVB13]. Woven [Boo12f]. Wrapper [LXB^{+}15]. Wrapping [GRDL^{+}12]. Written [Mai12b].

x86 [LNH^{+}11]. x86-Based [LNH^{+}11]. Xen [IGBKR19]. XML [BCLP15].

Yakov [Far17]. YAWL [AtHR11]. Years [BGK15, Boo13f, Bro18, BMSF15, Ebe18, EM18, EMP18, HWBYZ13, MW19, MH14, RSAT19, TD19, vGH19]. Yesterday [Bro18]. Yourself [Hol16a].

Zones [PMCE12].

References

Ameller:2013:NFR

[AACF13] David Ameller, Claudia Ayala, Jordi Cabot, and Xavier Franch. Nonfunctional requirements in architectural decision making. IEEE Software, 30(2):61–67, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Adams:2015:PFR

[ABB^{+}15a] Bram Adams, Stephany Bellomo, Christian Bird, Tamara Marshall-Keim, Foutse Khomh, and Kim Moir. The practice and future of release engineering: A roundtable with three release engineers. IEEE Software, 32(2):42– 49, March/April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020042.html.

Allier:2015:MDW

[ABB^{+}15b] Simon Allier, Olivier Barais, Benoit Baudry, Johann Bourcier, Erwan Daubert, Franck Fleurey, Martin Monperrus, Hui Song, and Maxime Tricoire. Multitier diversification in Web-based software applications. IEEE Software, 32(1):83–90, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010083-abs.
html.

Adams:2018:RE

[ABB^{+}18] B. Adams, S. Bellomo, C. Bird, B. Debić, F. Khomh, K. Moir, and J. O’Duinn. Release Engineering 3.0. IEEE Software, 35(2):22– 25, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ardagna:2011:SBF

[ABC^{+}11] Danilo Ardagna, Luciano Baresi, Sara Comai, Marco Comuzzi, and Barbara Pernici. A servicebased framework for flexible business processes. IEEE Software, 28(2):61– 67, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ardis:2011:ASE

[ABH^{+}11] Mark Ardis, Pierre Bourque, Thomas Hilburn, Kahina Lasfer, Scott Lucero, James McDonald, Art Pyster, and Mary Shaw. Advancing software engineering professional education. IEEE Software, 28(4):58–63, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Abrahamsson:2010:AAC

[ABK10] Pekka Abrahamsson, Muhammad Ali Babar, and Philippe Kruchten. Agility and architecture: Can they coexist? IEEE Software, 27(2):16–22, March/ April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ayala:2011:FFA

[ACHC11] Claudia P. Ayala, Daniela Cruzes, Oyvind Hauge, and Reidar Conradi. Five facts on the adoption of open source software. IEEE Software, 28(2):95– 99, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ard:2014:SBE

[ADH14] Jason Ard, Kristine Davidsen, and Terril Hurst. Simulation-based embedded agile development. IEEE Software, 31(2):97– 101, March/April 2014.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Antinyan:2018:MUT

[ADSS18] V. Antinyan, J. Derehag, A. Sandberg, and M. Staron. Mythical unit test coverage. IEEE Software, 35(3):73–79, May/ June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Amalfitano:2015:MAM

[AFT^{+}15]  Domenico  Amalfitano,

Anna Rita Fasolino, Porfirio Tramontana, Bryan Dzung Ta, and Atif M. Memon. MobiGUITAR: Automated model-based testing of mobile apps. IEEE Software, 32(5):53–59, September/ October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        05/mso2015050053-abs.
                                        html.

Alegroth:2017:RRS

[AGIF17] Emil Alégroth, Johan Gustafsson, Henrik Ivarsson, and Robert Feldt. Replicating rare software failures with exploratory visual GUI testing. IEEE Software, 34(5):53–59, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050053-abs.
html.

Aldrich:2019:MBA

[AGK^{+}19] J. Aldrich, D. Garlan, C. Kaestner, C. Le Goues, A. Mohseni-Kabir, I. Ruchkin, S. Samuel, B. Schmerl, C. S. Timperley, M. Veloso, I. Voysey, J. Biswas, A. Guha, J. Holtz, J. Camara, and P. Jamshidi. Model-based adaptation for robotics software. IEEE Software, 36(2):83–90, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Augustine:2018:DST

[AHMS18] Vinay Augustine, John Hudepohl, Przemyslaw Marcinczak, and Will Snipes. Deploying software team analytics in a multinational organization. IEEE Software, 35(1):72–76, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010072-abs.
html.

Autili:2015:ASS

[AIT15] Marco Autili, Paola Inverardi, and Massimo Tivoli. Automated synthesis of service choreographies. IEEE Software, 32(1):50–57, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        01/mso2015010050-abs.
                                        html.

Avgeriou:2016:RFS

[AKN^{+}16] Paris Avgeriou, Philippe Kruchten, Robert L. Nord, Ipek Ozkaya, and Carolyn Seaman. Reducing friction in software development. IEEE Software, 33(1):66–73, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2016/
                                        01/mso2016010066.html.

Alexander:2011:GSW

[Ale11] Ian Alexander. GORE, SORE, or what? IEEE Software, 28(1):8–10, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Alsmadi:2013:UME

[Als13] Izzat Mahmoud Alsmadi. Using mutation to enhance GUI testing coverage. IEEE Software, 30(1):

67–73, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Antonino:2016:ESA

[AMK16] Pablo Oliveira Antonino, Andreas Morgenstern, and Thomas Kuhn. Embeddedsoftware architects: It’s not only about the software. IEEE Software, 33(6):56–62, November/ December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060056-abs.
html.

Amoroso:2018:RPS

[Amo18]  E. Amoroso.  Recent

progress in software security. IEEE Software, 35(2): 11–13, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Andrews:2012:SI

[And12] Mike Andrews. Searching the Internet. IEEE Software, 29(2):13–16, March/ April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anderson:2015:DSE

[And15] Charles Anderson. Docker [software engineering]. IEEE Software, 32(3):102–c3, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        03/mso2015030102.html.

Anderson:2016:JPS

[And16] Charles Anderson. James Phillips on service discovery. IEEE Software, 33 (6):117–120, November/ December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        06/mso2016060117.html.

Anonymous:2010:R

[Ano10a] Anonymous. 2009 reviewers. IEEE Software, 27(1): 92–94, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2010:KBG

[Ano10b] Anonymous. Kudos to Bob Glass and Rebecca Wirfs-Brock. IEEE Software, 27(1):7–9, January/ February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2010:RHP

[Ano10c] Anonymous. Responses to “How Pair Programming Really Works”. IEEE Software, 27(2):8–9, March/ April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2011:EED

[Ano11a] Anonymous. elytS edoC detisiveR. IEEE Software, 28(4):7–8, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2011:RAM

[Ano11b] Anonymous. Remembering Ann Miller. IEEE Software, 28(5):8, September/ October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2012:ALP

[Ano12a] Anonymous. All late projects are the same. IEEE Software, 29(1):8–11, January/February 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2012:B

[Ano12b] Anonymous. I believe.

IEEE Software, 29(3):7, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2013:R

[Ano13a] Anonymous. 2012 reviewers. IEEE Software, 30(1): 92–93, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2013:DN

[Ano13b]  Anonymous.  David

Notkin: 1955–2013. IEEE Software, 30(4):7, July/ August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2013:LDD

[Ano13c]  Anonymous.  Letters:

Differential debugging. IEEE Software, 30(6):9– 10, November/December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2013:MR

[Ano13d] Anonymous. Mars Rover. IEEE Software, 30(4):8, July/August 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2013:PN

[Ano13e] Anonymous. Playing nice. IEEE Software, 30(1):7– 8, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:R

[Ano14a] Anonymous. 2013 reviewers. IEEE Software, 31(1): 85–87, January/February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:CP

[Ano14b]  Anonymous.  Call for

papers. IEEE Software, 31(6):c3, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso20140600c3.pdf.

Anonymous:2014:FYJb

[Ano14c] Anonymous. Focus on your job search [advertisement]. IEEE Software, 31 (4):c2, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:FYJa

[Ano14d] Anonymous. Focus on your job search [house advertisement]. IEEE Software, 31(3):55, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:FCa

[Ano14e] Anonymous. Front cover.

IEEE Software, 31(1):c1, January/February 2014.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:FCb

[Ano14f] Anonymous. [front cover]. IEEE Software, 31(2):c1, March/April 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:FCc

[Ano14g] Anonymous. Front cover. IEEE Software, 31(3):c1, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:FCd

[Ano14h] Anonymous. [front cover]. IEEE Software, 31(6): c1, November/December 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

                                        URL http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        06/mso20140600c1.pdf.

Anonymous:2014:ICCc

[Ano14i] Anonymous. IEEE CG&A calls for papers. IEEE Software, 31(3):25, May/ June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:ICCa

[Ano14j] Anonymous. IEEE Cloud Computing advertisement.

IEEE Software, 31(1):c4, January/February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:ICCb

[Ano14k] Anonymous. IEEE cloud computing [advertisement]. IEEE Software, 31(2):c3, March/April 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:ICS

[Ano14l] Anonymous. IEEE Computer Society membership advertisement. IEEE Software, 31(1):c2, January/ February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:ISP

[Ano14m] Anonymous. IEEE Security & Privacy [advertisement]. IEEE Software, 31 (4):81, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:ISD

[Ano14n] Anonymous. IEEE Software digital magazine advertisement. IEEE Software, 31(1):c3, January/ February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:IA

[Ano14o] Anonymous. Intelect [advertisement]. IEEE Software, 31(6):c2, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        06/mso20140600c2.pdf.

Anonymous:2014:Ma

[Ano14p] Anonymous. Masthead. IEEE Software, 31(2):4, March/April 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:Mb

[Ano14q] Anonymous. Masthead. IEEE Software, 31(3):1, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:Mc

[Ano14r] Anonymous. [masthead]. IEEE Software, 31(4):1, July/August 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:Md

[Ano14s] Anonymous. [masthead]. IEEE Software, 31(5):1, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050001.pdf.

Anonymous:2014:Me

[Ano14t] Anonymous. Masthead.

IEEE Software, 31(6): 1, November/December 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060001.pdf.

Anonymous:2014:MMAa

[Ano14u]  Anonymous.  Member-

ship matters [advertisement]. IEEE Software, 31 (2):c2, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:MMAb

[Ano14v]  Anonymous.  Member-

ship matters [advertisement]. IEEE Software, 31 (4):c3, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:MMAc

[Ano14w] Anonymous. Membership matters [advertisement]. IEEE Software, 31(5):c3, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        05/mso20140500c3.pdf.

Anonymous:2014:MMH

[Ano14x] Anonymous. Membership matters [house advertisement]. IEEE Software, 31(3):c3, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:T

[Ano14y] Anonymous. On target. IEEE Software, 31(5):8, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        05/mso2014050008.html.

Anonymous:2014:RSP

[Ano14z] Anonymous. Rock stars of 3D printing [advertisement]. IEEE Software, 31(6):c4, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        06/mso20140600c4.pdf.

Anonymous:2014:RSB

[Ano14-27] Anonymous. Rock stars of big data analytics [advertisement]. IEEE Software, 31(5):c4, September/October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso20140500c4.pdf.

Anonymous:2014:RSCa

[Ano14-28] Anonymous. Rock stars of cybersecurity [advertisement]. IEEE Software, 31 (4):c4, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:RSCb

[Ano14-29] Anonymous. Rock stars of cybersecurity [advertisement]. IEEE Software, 31(5):c2, September/October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso20140500c2.pdf.

Anonymous:2014:RSMa

[Ano14-30] Anonymous. Rock stars of mobile cloud [advertisement]. IEEE Software, 31 (2):c4, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:RSMb

[Ano14-31] Anonymous. Rock stars of mobile cloud [house advertisement]. IEEE Software, 31(3):c4, May/June 2014.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:SEC

[Ano14-32]  Anonymous.  Software

engineering for the 21st Century [house advertisement]. IEEE Software, 31(3):c2, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:SES

[Ano14-33]  Anonymous.  Software

experts summit. IEEE Software, 31(2):1, March/ April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:SPL

[Ano14-34] Anonymous. Software programming languages call for papers. IEEE Software, 31(1):1, January/ February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:SSU

[Ano14-35] Anonymous. Sweet spot for user involvement. IEEE Software, 31(3):9, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:TCa

[Ano14-36] Anonymous. Table of contents. IEEE Software, 31 (1):2–3, January/February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:TCb

[Ano14-37]  Anonymous.  Table of

contents. IEEE Software, 31(2):2–3, March/ April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:TCc

[Ano14-38] Anonymous. Table of contents. IEEE Software, 31 (3):2–3, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:TCd

[Ano14-39] Anonymous. Table of contents. IEEE Software, 31(4):2–3, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2014:TCe

[Ano14-40]  Anonymous.  Table of

contents. IEEE Software, 31(5):2–3, September/October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050002.pdf.

Anonymous:2014:VB

[Ano14-41] Anonymous. View from the bridge. IEEE Software, 31(2):10–13, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2015:R

[Ano15a]  Anonymous.  2014 re-

viewers. IEEE Software, 32(1):8–10, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010008.html.

Anonymous:2015:RMD

[Ano15b] Anonymous. 2016 Richard E. Merwin Distinguished Service Award house advertisement. IEEE Software, 32(5):c3, September/October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso20150500c3.pdf.

Anonymous:2015:CPa

[Ano15c]  Anonymous.  Call for

papers. IEEE Software, 32(1):1, January/ February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010001.pdf.

Anonymous:2015:CPb

[Ano15d] Anonymous. Call for papers. IEEE Software, 32 (2):c2, March/April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso20150200c2.pdf.

Anonymous:2015:CPH

[Ano15e] Anonymous. Call for papers house advertisement. IEEE Software, 32(2):c3, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso20150200c3.pdf.

Anonymous:2015:CPY

[Ano15f]  Anonymous.  Confer-

ences in the palm of your hand house advertisement. IEEE Software, 32(2):41, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020041.pdf.

Anonymous:2015:FYJ

[Ano15g] Anonymous. Focus on your job search [advertisement]. IEEE Software, 32(1):c3, January/ February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        01/mso20150100c3.pdf.

Anonymous:2015:FCa

[Ano15h] Anonymous. [front cover]. IEEE Software, 32(1):c1, January/February 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        01/mso20150100c1.pdf.

Anonymous:2015:FCb

[Ano15i] Anonymous. Front cover. IEEE Software, 32(2):c1, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        02/mso20150200c1.pdf.

Anonymous:2015:FCc

[Ano15j] Anonymous. [front cover]. IEEE Software, 32(3):c1, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso20150300c1.pdf.

Anonymous:2015:FCd

[Ano15k] Anonymous. Front cover.

IEEE Software, 32(4):c1, July/August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso20150400c1.pdf.

Anonymous:2015:FCe

[Ano15l] Anonymous. Front cover.

IEEE Software, 32(5):c1, September/October 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso20150500c1.pdf.

Anonymous:2015:FCf

[Ano15m] Anonymous. Front cover.

IEEE Software, 32(6): c1, November/December 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL http://csdl.computer.
org/csdl/mags/so/2015/
06/mso20150600c1.pdf.

Anonymous:2015:FSE

[Ano15n] Anonymous. The future of software engineering house advertisement. IEEE Software, 32(5):c2, September/October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso20150500c2.pdf.

Anonymous:2015:ISP

[Ano15o] Anonymous. IEEE Security & Privacy house advertisement. IEEE Software, 32(4):c2, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso20150400c2.pdf.

Anonymous:2015:IA

[Ano15p] Anonymous. Intelect [advertisement]. IEEE Software, 32(1):c2, January/ February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso20150100c2.pdf.

Anonymous:2015:KYCa

[Ano15q] Anonymous. Keep your career moving forward house advertisement. IEEE Software, 32(5):1, September/October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.

org/csdl/mags/so/2015/
05/mso2015050001.pdf.

Anonymous:2015:KYCb

[Ano15r] Anonymous. Keep your career moving forward house advertisement. IEEE Software, 32(6):1, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060001.pdf.

Anonymous:2015:Ma

[Ano15s] Anonymous. [masthead].

IEEE Software, 32(1):7, January/February 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010007.pdf.

Anonymous:2015:Mb

[Ano15t] Anonymous. Masthead.

IEEE Software, 32(2):6, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020006.pdf.

Anonymous:2015:Mc

[Ano15u] Anonymous. [masthead].

IEEE Software, 32(3):6, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        03/mso2015030006.pdf.

Anonymous:2015:Md

[Ano15v] Anonymous. Masthead. IEEE Software, 32(4):6, July/August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        04/mso2015040006.pdf.

Anonymous:2015:Me

[Ano15w] Anonymous. Masthead. IEEE Software, 32(5):8, September/October 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        05/mso2015050008.pdf.

Anonymous:2015:Mf

[Ano15x] Anonymous. Masthead. IEEE Software, 32(6): 7, November/December 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

                                        URL http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        06/mso2015060007.pdf.

Anonymous:2015:RSP

[Ano15y] Anonymous. Rock stars of 3D printing [advertisement]. IEEE Software, 32(1):c4, January/ February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso20150100c4.pdf.

Anonymous:2015:RSCa

[Ano15z] Anonymous. Rock stars of cybersecurity [advertisement]. IEEE Software, 32(3):c4, May/June 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso20150300c4.pdf.

Anonymous:2015:RSCb

[Ano15-27] Anonymous. Rock stars of cybersecurity house advertisement. IEEE Software, 32(6):c3, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso20150600c3.pdf.

Anonymous:2015:RSWa

[Ano15-28] Anonymous. Rock stars of wearables house advertisement. IEEE Software, 32 (4):c4, July/August 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.

                                        org/csdl/mags/so/2015/
                                        04/mso20150400c4.pdf.

Anonymous:2015:RSWb

[Ano15-29] Anonymous. Rock stars of wearables house advertisement. IEEE Software, 32(5):37, September/October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        05/mso2015050037.pdf.

Anonymous:2015:SES

[Ano15-30] Anonymous. Software experts summit house advertisement. IEEE Software, 32(6):c4, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        06/mso20150600c4.pdf.

Anonymous:2015:SRS

[Ano15-31] Anonymous. Startup rock stars house advertisement. IEEE Software, 32(2):c4, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        02/mso20150200c4.pdf.

Anonymous:2015:SCP

[Ano15-32] Anonymous. STC 2015:

Call for presentations. IEEE Software, 32(3):c2, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso20150300c2.pdf.

Anonymous:2015:TCa

[Ano15-33]  Anonymous.  Table of

contents. IEEE Software, 32(1):2–3, January/ February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010002.pdf.

Anonymous:2015:TCb

[Ano15-34] Anonymous. Table of contents. IEEE Software, 32 (3):1–2, May/June 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030001.pdf.

Anonymous:2015:ULH

[Ano15-35]  Anonymous.  USENIX

LISA 15 house advertisement. IEEE Software, 32(5):c4, September/October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso20150500c4.pdf.

Anonymous:2015:WWL

[Ano15-36] Anonymous. Watch the world’s leading experts take multi-core strategies to new heights house advertisement. IEEE Software, 32(6):c2, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso20150600c2.pdf.

Anonymous:2016:R

[Ano16a]  Anonymous.  2015 re-

viewers. IEEE Software, 33(1):6–8, January/ February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010006.html.

Anonymous:2016:RMA

[Ano16b] Anonymous. 2017 Richard E. Merwin Award for Distinguished Service. IEEE Software, 33(5):116, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.

org/csdl/mags/so/2016/
05/mso2016050116.pdf.

Anonymous:2016:AYC

[Ano16c] Anonymous. Achieve your career goals with the fit that’s right for you [Advertisement]. IEEE Software, 33(5):1, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050001.pdf.

Anonymous:2016:CNE

[Ano16d] Anonymous. Call for nominees education awards nominations. IEEE Software, 33(5):47, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050047.pdf.

Anonymous:2016:CEA

[Ano16e] Anonymous. Computer entrepreneur award. IEEE Software, 33(6):93, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060093.pdf.

Anonymous:2016:FYJb

[Ano16f] Anonymous. Focus on your job search. IEEE Software, 33(6):40, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        06/mso2016060040.pdf.

Anonymous:2016:FYJa

[Ano16g] Anonymous. Focus on your job search house advertisement. IEEE Software, 33(1):55, January/ February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2016/
                                        01/mso2016010055.pdf.

Anonymous:2016:FCa

[Ano16h] Anonymous. Front cover. IEEE Software, 33(1):c1, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2016/
                                        01/mso20160100c1.pdf.

Anonymous:2016:FCb

[Ano16i] Anonymous. Front cover. IEEE Software, 33(2):c1, March/April 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
02/mso20160200c1.pdf.

Anonymous:2016:FCc

[Ano16j] Anonymous. [front cover].

IEEE Software, 33(3):c1, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso20160300c1.pdf.

Anonymous:2016:FCd

[Ano16k] Anonymous. Front cover.

IEEE Software, 33(4):c1, July/August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso20160400c1.pdf.

Anonymous:2016:FCe

[Ano16l] Anonymous. Front cover.

IEEE Software, 33(5):c1, September/October 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso20160500c1.pdf.

Anonymous:2016:FCf

[Ano16m] Anonymous. [front cover].

IEEE Software, 33(6): c1, November/December 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2016/
06/mso20160600c1.pdf.

Anonymous:2016:ICC

[Ano16n] Anonymous. IEEE Cloud Computing call for papers house advertisement. IEEE Software, 33(1):23, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010023.pdf.

Anonymous:2016:ICSj

[Ano16o] Anonymous. IEEE Computer Society. IEEE Software, 33(6):c3, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso20160600c3.pdf.

Anonymous:2016:ICSd

[Ano16p] Anonymous. IEEE Computer Society 2016 call for major award nominations. IEEE Software, 33 (3):c3, May/June 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso20160300c3.pdf.

Anonymous:2016:ICSg

[Ano16q] Anonymous. IEEE Computer Society 2016 call for major award nominations. IEEE Software, 33(5):71, September/October 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050071.pdf.

Anonymous:2016:ICSf

[Ano16r] Anonymous. IEEE Computer Society 2016 call for major award nominations house advertisement. IEEE Software, 33 (4):c3, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso20160400c3.pdf.

Anonymous:2016:ICSh

[Ano16s] Anonymous. IEEE Computer Society: Be at the center of it all. IEEE Software, 33(5):c3, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso20160500c3.pdf.

Anonymous:2016:ICSa

[Ano16t] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. IEEE Software, 33(1):c4, January/ February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso20160100c4.pdf.

Anonymous:2016:ICSb

[Ano16u] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. IEEE Software, 33(2):c2, March/ April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http:

//csdl.computer.org/
csdl/mags/so/2016/02/
mso20160200c2.pdf.

Anonymous:2016:ICSe

[Ano16v] Anonymous. IEEE Computer Society get the recognition you deserve. IEEE Software, 33(3):c4, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso20160300c4.pdf.

Anonymous:2016:ICSi

[Ano16w] Anonymous. IEEE Computer Society is where you choose the resources that fit your career. IEEE Software, 33(6):1, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060001.pdf.

Anonymous:2016:ICSc

[Ano16x] Anonymous. IEEE Computer Society rock stars of big data. IEEE Software, 33(3):c2, May/June 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso20160300c2.pdf.

Anonymous:2016:ITS

[Ano16y] Anonymous. IEEE Transactions on Sustainable Computing house advertisement. IEEE Software, 33(1):81, January/ February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010081.pdf.

Anonymous:2016:Ma

[Ano16z] Anonymous. Masthead.

IEEE Software, 33(1):9, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.

                                        org/csdl/mags/so/2016/
                                        01/mso2016010009.pdf.

Anonymous:2016:Mb

[Ano16-27] Anonymous. Masthead. IEEE Software, 33(2):7, March/April 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        02/mso2016020007.pdf.

Anonymous:2016:Mc

[Ano16-28] Anonymous. [masthead]. IEEE Software, 33(3):1, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        03/mso2016030001.pdf.

Anonymous:2016:Md

[Ano16-29] Anonymous. Masthead. IEEE Software, 33(4):1, July/August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        04/mso2016040001.pdf.

Anonymous:2016:Me

[Ano16-30] Anonymous. Masthead. IEEE Software, 33(5):8, September/October 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050008.pdf.

Anonymous:2016:Mf

[Ano16-31] Anonymous. Masthead.

IEEE Software, 33(6): 7, November/December 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060007.pdf.

Anonymous:2016:MTC

[Ano16-32]  Anonymous.  Modeling

test cases in BPMN for behavior-driven development. IEEE Software, 33(5):15–21, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050015-abs.
html.

Anonymous:2016:NMOa

[Ano16-33] Anonymous. New membership options for a better fit. IEEE Software, 33(5):c2, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso20160500c2.pdf.

Anonymous:2016:NMOb

[Ano16-34] Anonymous. New membership options for a better fit. IEEE Software, 33(6):c2, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        06/mso20160600c2.pdf.

Anonymous:2016:RYF

[Ano16-35] Anonymous. Read your favorite publications your way house advertisement. IEEE Software, 33(2):1, March/April 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        02/mso2016020001.pdf.

Anonymous:2016:RSBb

[Ano16-36] Anonymous. Rock stars of big data. IEEE Software, 33(5):95, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        05/mso2016050095.pdf.

Anonymous:2016:RSBa

[Ano16-37] Anonymous. Rock stars of big data house advertisement. IEEE Software, 33 (4):c2, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso20160400c2.pdf.

Anonymous:2016:RSCb

[Ano16-38] Anonymous. Rock stars of cyber security. IEEE Software, 33(5):c4, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso20160500c4.pdf.

Anonymous:2016:RSCa

[Ano16-39] Anonymous. Rock stars of cybersecurity house advertisement. IEEE Software, 33(4):c4, July/ August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso20160400c4.pdf.

Anonymous:2016:RSP

[Ano16-40] Anonymous. Rock stars of pervasive, predictive analytics. IEEE Software, 33(6):c4, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.

                                        org/csdl/mags/so/2016/
                                        06/mso20160600c4.pdf.

Anonymous:2016:RSR

[Ano16-41] Anonymous. Rock stars of risk-based security house advertisement. IEEE Software, 33(2):c4, March/ April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http:

                                        //csdl.computer.org/
                                        csdl/mags/so/2016/02/
                                        mso20160200c4.pdf.

Anonymous:2016:SAR

[Ano16-42] Anonymous. The software architect’s role in the digital age. IEEE Software, 33 (2):73, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        02/mso2016020073.pdf.

Anonymous:2016:SEI

[Ano16-43] Anonymous. Software engineering for the Internet of Things house advertisement. IEEE Software, 33 (2):c3, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        02/mso20160200c3.pdf.

Anonymous:2016:SCH

[Ano16-44] Anonymous. Stay connected house advertisement. IEEE Software, 33 (1):c3, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL http://www.computer.
org/csdl/mags/so/2016/
01/mso20160100c3.pdf.

Anonymous:2016:TCa

[Ano16-45] Anonymous. Table of contents. IEEE Software, 33 (3):2–3, May/June 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030002.pdf.

Anonymous:2016:TCb

[Ano16-46]  Anonymous.  Table of

contents. IEEE Software, 33(4):2–3, July/ August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040002.pdf.

Anonymous:2016:TCc

[Ano16-47]  Anonymous.  Table of

contents. IEEE Software, 33(5):2–3, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        05/mso2016050002.pdf.

Anonymous:2016:TCd

[Ano16-48]  Anonymous.  Table of

contents. IEEE Software, 33(6):2–3, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        06/mso2016060002.pdf.

Anonymous:2016:TP

[Ano16-49] Anonymous. Time pressure. IEEE Software, 33(3):6, May/June 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        03/mso2016030006.html.

Anonymous:2016:WWL

[Ano16-50] Anonymous. Watch the world’s leading experts take multi-core strategies to new heights house advertisement. IEEE Software, 33(1):c2, January/ February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2016/
                                        01/mso20160100c2.pdf.

Anonymous:2017:R

[Ano17a] Anonymous. 2016 reviewers. IEEE Software, 34(2): 8–10, March/April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020008.html.

Anonymous:2017:ICSc

[Ano17b] Anonymous. 40th international conference on software engineering. IEEE Software, 34(4):c2, July/ August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso20170400c2.pdf.

Anonymous:2017:ASV

[Ano17c] Anonymous. Accessible software verification with Dafny. IEEE Software, 34(6):94–97, November/ December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060094-abs.
html.

Anonymous:2017:AYCb

[Ano17d] Anonymous. Achieve your career goals with the fit that’s right for you. IEEE Software, 34(6):1, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060001.pdf.

Anonymous:2017:AYCa

[Ano17e]  Anonymous.  Achieve

your career goals with the fit that’s right for you. house advertisement. IEEE Software, 34(3):1, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030001.pdf.

Anonymous:2017:AIC

[Ano17f]  Anonymous.  ACM–

IEEE CS Eckert–Mauchly Award. IEEE Software, 34(1):71, January/ February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010071.pdf.

Anonymous:2017:BTM

[Ano17g] Anonymous. Become a TCSE member TODAY! IEEE Software, 34(5):c2, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso20170500c2.pdf.

Anonymous:2017:CNE

[Ano17h] Anonymous. Call for nominees education awards nominations. IEEE Software, 34(4):48, July/ August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040048.pdf.

Anonymous:2017:CSA

[Ano17i]  Anonymous.  Call for

standards award nominations. IEEE Software, 34(6):41, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060041.pdf.

Anonymous:2017:CYI

[Ano17j] Anonymous. Can you invent a better world through technologies? IEEE Software, 34(1):c3, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso20170100c3.pdf.

Anonymous:2017:CPC

[Ano17k] Anonymous. CiSE is the perfect combination. IEEE Software, 34(1):81, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        01/mso2017010081.pdf.

Anonymous:2017:CPY

[Ano17l]  Anonymous.  Confer-

ences in the palm of your hand house advertisement. IEEE Software, 34(3):104, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        03/mso2017030104.pdf.

Anonymous:2017:CAS

[Ano17m]  Anonymous.  Context-

aware software variability through adaptable interpreters. IEEE Software, 34(6):83–88, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        06/mso2017060083-abs.
                                        html.

Anonymous:2017:DWS

[Ano17n] Anonymous. Did 32% waterfall surprise you? IEEE Software, 34(1):7, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010007.html.

Anonymous:2017:EML

[Ano17o]  Anonymous.  The end

of the manufacturing-line analogy. IEEE Software, 34(6):89–93, November/ December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060089-abs.
html.

Anonymous:2017:ESP

[Ano17p] Anonymous. Evaluating software project managers: A multidimensional perspective. IEEE Software, 34(6):104–108, November/ December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060104-abs.
html.

Anonymous:2017:FYJ

[Ano17q] Anonymous. Focus on your job search house advertisement. IEEE Software, 34(3):72, May/ June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL https://

                                        www.computer.org/csdl/
                                        mags/so/2017/03/mso2017030072.
                                        pdf.

Anonymous:2017:FCa

[Ano17r] Anonymous. [front cover]. IEEE Software, 34(1):c1, January/February 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        01/mso20170100c1.pdf.

Anonymous:2017:FCb

[Ano17s] Anonymous. Front cover. IEEE Software, 34(2):c1, March/April 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        02/mso20170200c1.pdf.

Anonymous:2017:FCc

[Ano17t] Anonymous. Front cover. IEEE Software, 34(3):c1, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        03/mso20170300c1.pdf.

Anonymous:2017:FCd

[Ano17u] Anonymous. Front cover. IEEE Software, 34(4):c1, July/August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso20170400c1.pdf.

Anonymous:2017:FCe

[Ano17v] Anonymous. Front cover.

IEEE Software, 34(5):c1, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso20170500c1.pdf.

Anonymous:2017:FCf

[Ano17w] Anonymous. Front cover.

IEEE Software, 34(6): c1, November/December 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2017/
06/mso20170600c1.pdf.

Anonymous:2017:ICSb

[Ano17x] Anonymous. IEEE Computer Society Harlan D. Mills Award house advertisement. IEEE Software, 34(3):c4, May/June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso20170300c4.pdf.

Anonymous:2017:ICSa

[Ano17y] Anonymous. IEEE Computer Society is where you choose the resources that fit your career. IEEE Software, 34(1):1, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        01/mso2017010001.pdf.

Anonymous:2017:ICSd

[Ano17z] Anonymous. International conference on software engineering. IEEE Software, 34(5):c3, September/October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        05/mso20170500c3.pdf.

Anonymous:2017:Ma

[Ano17-27] Anonymous. Masthead. IEEE Software, 34(1):8, January/February 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        01/mso2017010008.pdf.

Anonymous:2017:Mb

[Ano17-28] Anonymous. Masthead. IEEE Software, 34(2):7, March/April 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020007.pdf.

Anonymous:2017:Mc

[Ano17-29] Anonymous. Masthead.

IEEE Software, 34(3):8, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030008.pdf.

Anonymous:2017:Md

[Ano17-30] Anonymous. Masthead.

IEEE Software, 34(4):3, July/August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040003.pdf.

Anonymous:2017:Me

[Ano17-31] Anonymous. Masthead.

IEEE Software, 34(5):1, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050001.pdf.

Anonymous:2017:Mf

[Ano17-32] Anonymous. Masthead.

IEEE Software, 34(6): 13, November/December 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060013.pdf.

Anonymous:2017:Mg

[Ano17-33]  Anonymous.  myCS.

IEEE Software, 34(6): 71, November/December 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060071.pdf.

Anonymous:2017:NMOa

[Ano17-34] Anonymous. New membership options for a better fit. IEEE Software, 34(1):c2, January/ February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso20170100c2.pdf.

Anonymous:2017:NMOc

[Ano17-35] Anonymous. New membership options for a better fit. IEEE Software, 34 (4):c3, July/August 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso20170400c3.pdf.

Anonymous:2017:NMOd

[Ano17-36] Anonymous. New membership options for a better fit. IEEE Software, 34(5):c4, September/October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso20170500c4.pdf.

Anonymous:2017:NMOe

[Ano17-37] Anonymous. New membership options for a better fit. IEEE Software, 34(6):c2, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso20170600c2.pdf.

Anonymous:2017:NMOb

[Ano17-38] Anonymous. New membership options for a better fit. House advertisement. IEEE Software, 34(3):c2, May/June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso20170300c2.pdf.

Anonymous:2017:S

[Ano17-39]  Anonymous.  Not so

secure. IEEE Software, 34(1):29, January/ February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010029.pdf.

Anonymous:2017:OMUb

[Ano17-40] Anonymous. One membership. Unlimited knowledge. IEEE Software, 34 (6):c4, November/December 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2017/
06/mso20170600c4.pdf.

Anonymous:2017:OMUa

[Ano17-41] Anonymous. One membership. unlimited knowledge. house advertisement. IEEE Software, 34(3):48, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030048.pdf.

Anonymous:2017:PCa

[Ano17-42] Anonymous. Prepare to connect. IEEE Software, 34(4):c4, July/ August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.

org/csdl/mags/so/2017/
04/mso20170400c4.pdf.

Anonymous:2017:PCb

[Ano17-43] Anonymous. Prepare to connect. IEEE Software, 34(6):c3, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso20170600c3.pdf.

Anonymous:2017:PCH

[Ano17-44] Anonymous. Prepare to connect house advertisement. IEEE Software, 34(3):c3, May/June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso20170300c3.pdf.

Anonymous:2017:TCa

[Ano17-45]  Anonymous.  Table of

contents. IEEE Software, 34(1):2–3, January/ February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010002.pdf.

Anonymous:2017:TCb

[Ano17-46]  Anonymous.  Table of

contents. IEEE Software, 34(2):2–3, March/ April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL https://

www.computer.org/csdl/
mags/so/2017/02/mso2017020002.
pdf.

Anonymous:2017:TCc

[Ano17-47] Anonymous. Table of contents. IEEE Software, 34 (3):2–3, May/June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030002.pdf.

Anonymous:2017:TCd

[Ano17-48]  Anonymous.  Table of

contents. IEEE Software, 34(4):1–2, July/ August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040001.pdf.

Anonymous:2017:TCe

[Ano17-49]  Anonymous.  Table of

contents. IEEE Software, 34(5):2–3, September/October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050002.pdf.

Anonymous:2017:TCf

[Ano17-50]  Anonymous.  Table of

contents. IEEE Software, 34(6):2–3, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060002.pdf.

Anonymous:2017:T

[Ano17-51] Anonymous. TechIgnite.

IEEE Software, 34(1):c4, January/February 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso20170100c4.pdf.

Anonymous:2017:WWL

[Ano17-52] Anonymous. Watch the world’s leading experts take multi-core strategies to new heights. IEEE Software, 34(1):37, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010037.pdf.

Anonymous:2018:ICCa

[Ano18a] Anonymous. 2019 IEEE-CS Charles Babbage Award. IEEE Software, 35(4):43, July/August 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:ICCb

[Ano18b] Anonymous. 2019 IEEE-CS Charles Babbage Award. IEEE Software, 35(5):c3, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:ICSb

[Ano18c] Anonymous. 40th International Conference on Software Engineering. IEEE Software, 35(2):c2, March/ April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:ICSg

[Ano18d]  Anonymous.  41st In-

ternational Conference on Software Engineering. IEEE Software, 35(5):c4, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:ITS

[Ano18e] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 35(6):55, November/ December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:CSA

[Ano18f] Anonymous. Call for standards award nominations. IEEE Software, 35(1):c3, January/February 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso20180100c3.pdf.

Anonymous:2018:CDD

[Ano18g] Anonymous. COMPSAC 2019 data driven intelligence for a smarter world. IEEE Software, 35(6): c4, November/December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:CPY

[Ano18h] Anonymous. Conferences in the palm of your hand. IEEE Software, 35(6): 61, November/December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:CI

[Ano18i] Anonymous. Connect on interface. IEEE Software, 35(4):94, July/ August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:FCa

[Ano18j] Anonymous. Front cover.

IEEE Software, 35(1):c1, January/February 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2018/
                                        01/mso20180100c1.pdf.

Anonymous:2018:FCb

[Ano18k] Anonymous. Front cover. IEEE Software, 35(2):c1, March/April 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:FCc

[Ano18l] Anonymous. Front cover. IEEE Software, 35(3):c1, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:FCd

[Ano18m] Anonymous. Front cover. IEEE Software, 35(4):c1, July/August 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:FCe

[Ano18n] Anonymous. Front cover. IEEE Software, 35(5):c1, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:FCf

[Ano18o] Anonymous. Front cover. IEEE Software, 35(6): c1, November/December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:ICSa

[Ano18p] Anonymous. IEEE Computer Society. IEEE Software, 35(1):77, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010077.pdf.

Anonymous:2018:ICSc

[Ano18q] Anonymous. IEEE Computer Society. IEEE Software, 35(2):51, March/ April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:ICSe

[Ano18r] Anonymous. IEEE Computer Society. IEEE Software, 35(3):c3, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:ICSf

[Ano18s] Anonymous. IEEE Computer Society. IEEE Software, 35(4):c3, July/ August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:ICSd

[Ano18t] Anonymous. IEEE Computer Society: Be at the center of it all. IEEE Software, 35(2):95, March/ April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:ICSh

[Ano18u] Anonymous. IEEE Computer Society: Be at the center of it all. IEEE Software, 35(6):c3, November/ December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:LBT

[Ano18v]  Anonymous.  Looking

for the BEST tech job for you? IEEE Software, 35(6):c2, November/ December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:Ma

[Ano18w] Anonymous. Masthead. IEEE Software, 35(1):1, January/February 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2018/
                                        01/mso2018010001.pdf.

Anonymous:2018:Mc

[Ano18x] Anonymous. [masthead].

IEEE Software, 35(2):1, March/April 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:Me

[Ano18y] Anonymous. [masthead].

IEEE Software, 35(3):1, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:Mg

[Ano18z] Anonymous. [masthead].

IEEE Software, 35(4):1, July/August 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:Mi

[Ano18-27] Anonymous. [masthead].

IEEE Software, 35(5):c2, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:Mj

[Ano18-28] Anonymous. [masthead].

IEEE Software, 35(6): 1, November/December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:Mb

[Ano18-29] Anonymous. myCS. IEEE Software, 35(1):c4, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso20180100c4.pdf.

Anonymous:2018:Md

[Ano18-30] Anonymous. myCS. IEEE Software, 35(2):c3, March/ April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:Mf

[Ano18-31] Anonymous. myCS. IEEE Software, 35(3):c4, May/ June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:Mh

[Ano18-32] Anonymous. myCS. IEEE Software, 35(4):c4, July/ August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:Mk

[Ano18-33]  Anonymous.  myCS.

IEEE Software, 35(6): 39, November/December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:OMUa

[Ano18-34] Anonymous. One membership. Unlimited knowledge. IEEE Software, 35 (2):c4, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:OMUb

[Ano18-35] Anonymous. One membership. Unlimited knowledge. IEEE Software, 35(4):100, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:PC

[Ano18-36] Anonymous. Prepare to connect. IEEE Software, 35(1):c2, January/ February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso20180100c2.pdf.

Anonymous:2018:REHa

[Ano18-37] Anonymous. Recognizing excellence in high performance computing nominations are solicited for the Seymour Cray / Sidney Fernbach / Ken Kennedy Awards. IEEE Software, 35(3):c2, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:REHb

[Ano18-38] Anonymous. Recognizing excellence in high performance computing nominations are solicited for the Seymour Cray / Sidney Fernbach / Ken Kennedy Awards. IEEE Software, 35(4):c2, July/ August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:TCa

[Ano18-39]  Anonymous.  Table of

contents. IEEE Software, 35(1):2–3, January/ February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2018/
                                        01/mso2018010002.pdf.

Anonymous:2018:TCb

[Ano18-40]  Anonymous.  Table of

contents. IEEE Software, 35(2):2–3, March/ April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:TCc

[Ano18-41] Anonymous. Table of contents. IEEE Software, 35 (3):2–3, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:TCd

[Ano18-42] Anonymous. Table of contents. IEEE Software, 35(4):2–3, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:TCe

[Ano18-43]  Anonymous.  Table of

contents. IEEE Software, 35(5):1–2, September/October 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:TCf

[Ano18-44]  Anonymous.  Table of

contents. IEEE Software, 35(6):2–3, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ICSe

[Ano19a] Anonymous. 2019 IEEE Computer Society election. IEEE Software, 36 (5):28, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ICSf

[Ano19b] Anonymous. 2019 IEEE Computer Society election. IEEE Software, 36 (5):84, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:AIC

[Ano19c]  Anonymous.  ACM/

IEEE CS Eckert–Mauchly Award call for award nominations. IEEE Software, 36(2):C4, March/ April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:IAHa

[Ano19d] Anonymous. IEEE Annals of the History of Computing. IEEE Software, 36 (2):82, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:IAHb

[Ano19e] Anonymous. IEEE Annals of the History of Computing. IEEE Software, 36(6):22, November/ December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ICG

[Ano19f] Anonymous. IEEE Computer Graphics and Applications. IEEE Software, 36(2):109, March/ April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ICE

[Ano19g] Anonymous. IEEE Computing Edge magazine subscribe. IEEE Software, 36(6):16, November/ December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:IIS

[Ano19h] Anonymous. IEEE Intelligent Systems magazine. IEEE Software, 36(3):111, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ISPa

[Ano19i] Anonymous. IEEE Security & Privacy. IEEE Software, 36(2):57, March/ April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ISPb

[Ano19j] Anonymous. IEEE Software and Privacy magazine subscribe. IEEE Software, 36(6):42, November/ December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ITBa

[Ano19k] Anonymous. IEEE Transactions on Big Data. IEEE Software, 36(2):63, March/April 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ITBb

[Ano19l] Anonymous. IEEE Transactions on Big Data. IEEE Software, 36(5):45, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ITSa

[Ano19m] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 36(2):69, March/ April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ITSb

[Ano19n] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 36(3):91, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:SMC

[Ano19o]  Anonymous.  Software

magazine call for articles. IEEE Software, 36(3):119, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:TBDa

[Ano19p] Anonymous. Transactions on Big Data. IEEE Software, 36(3):27, May/June 2019. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:TBDb

[Ano19q] Anonymous. Transactions on Big Data. IEEE Software, 36(6):82, November/ December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:C

[Ano19r] Anonymous. COMPSAC 2019. IEEE Software, 36 (3):C4, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:FCa

[Ano19s] Anonymous. Front cover.

IEEE Software, 36(1):C1, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:FCb

[Ano19t] Anonymous. Front cover.

IEEE Software, 36(2):C1, March/April 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:FCc

[Ano19u] Anonymous. Front cover.

IEEE Software, 36(3):C1, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:FCd

[Ano19v] Anonymous. Front cover. IEEE Software, 36(4):C1, July/August 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:FCe

[Ano19w] Anonymous. Front cover. IEEE Software, 36(5):C1, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:FCf

[Ano19x] Anonymous. Front cover. IEEE Software, 36(6): C1, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ICSa

[Ano19y] Anonymous. IEEE Computer Society information. IEEE Software, 36(1):C2, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ICSb

[Ano19z] Anonymous. IEEE Computer Society information. IEEE Software, 36(2):C2, March/April 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ICSc

[Ano19-27] Anonymous. IEEE Computer Society information. IEEE Software, 36(3):C2, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ICSd

[Ano19-28] Anonymous. IEEE Computer Society information. IEEE Software, 36(4):C3, July/August 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ICSg

[Ano19-29] Anonymous. IEEE Computer Society information. IEEE Software, 36(5):C3, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ICSh

[Ano19-30] Anonymous. IEEE Computer Society information. IEEE Software, 36(6): C3, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:IJBa

[Ano19-31] Anonymous. IEEE Jobs Board. IEEE Software, 36 (4):C4, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:IJBb

[Ano19-32] Anonymous. IEEE Jobs Board. IEEE Software, 36(5):C4, September/October 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:IJBc

[Ano19-33] Anonymous. IEEE Jobs Board. IEEE Software, 36(6):C4, November/December 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:ISN

[Ano19-34] Anonymous. IEEE Social Networking. IEEE Software, 36(6):88, November/ December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:LBT

[Ano19-35] Anonymous. Looking for the BEST tech job for you? IEEE Software, 36 (2):C3, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:Ma

[Ano19-36] Anonymous. Masthead. IEEE Software, 36(1):1, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:Mb

[Ano19-37] Anonymous. Masthead.

IEEE Software, 36(2):1, March/April 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:Mc

[Ano19-38] Anonymous. Masthead.

IEEE Software, 36(3):1, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:Md

[Ano19-39] Anonymous. Masthead.

IEEE Software, 36(4):C2, July/August 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:Me

[Ano19-40] Anonymous. Masthead.

IEEE Software, 36(5):C2, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:Mf

[Ano19-41] Anonymous. Masthead.

IEEE Software, 36(6): C2, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:SLa

[Ano19-42] Anonymous. Staff listing. IEEE Software, 36(3):5–6, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:SLb

[Ano19-43] Anonymous. Staff listing. IEEE Software, 36(4):4–5, July/August 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:SLc

[Ano19-44] Anonymous. Staff listing. IEEE Software, 36(5):4–5, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:SLd

[Ano19-45] Anonymous. Staff listing. IEEE Software, 36(6): 4–5, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:SSPa

[Ano19-46] Anonymous. Subscribe to Security & Privacy magazine. IEEE Software, 36(3):8, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:SSPb

[Ano19-47] Anonymous. Subscribe to Security & Privacy magazine. IEEE Software, 36(5):75, September/October 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:TCa

[Ano19-48] Anonymous. Table of contents. IEEE Software, 36 (1):2–3, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:TCb

[Ano19-49]  Anonymous.  Table of

contents. IEEE Software, 36(2):2–3, March/ April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:TCc

[Ano19-50] Anonymous. Table of contents. IEEE Software, 36 (3):2–3, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:TCd

[Ano19-51] Anonymous. Table of contents. IEEE Software, 36(4):1–2, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:TCe

[Ano19-52]  Anonymous.  Table of

contents. IEEE Software, 36(5):1–2, September/October 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:TCf

[Ano19-53]  Anonymous.  Table of

contents. IEEE Software, 36(6):1–2, November/December 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2019:WCS

[Ano19-54]  Anonymous.  World

Congress On Services 2019. IEEE Software, 36 (3):C3, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Avelino:2019:WCM

[APPV19] G. Avelino, L. Passos, F. Petrillo, and M. T. Valente. Who can maintain this code?: Assessing the effectiveness of repositorymining techniques for identifying software maintainers. IEEE Software, 36(6):34–42, November/ December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Aerts:2017:HSC

[AS17] Hans Aerts and Han Schaminee. How software is changing the automotive landscape. IEEE Software, 34(6):7–12, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060007.html.

Almutairi:2012:DAC

[ASB^{+}12] Abdulrahman A. Almutairi, Muhammad I. Sarfraz, Saleh Basalamah, Walid G. Aref, and Arif Ghafoor. A distributed access control architecture for Cloud computing. IEEE Software, 29(2): 36–44, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Avgeriou:2013:ASG

[ASH13] Paris Avgeriou, Michael Stal, and Rich Hilliard. Architecture sustainability [Guest Editors’ introduction]. IEEE Software, 30(6):40–44, November/ December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Atkinson:2014:FGI

[ASK14] Colin Atkinson, Thomas Schulze, and Sonja Klingert.

Facilitating greener IT through green specifications. IEEE Software, 31 (3):56–63, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Abelein:2013:DIU

[ASP13] Ulrike Abelein, Helen Sharp, and Barbara Paech. Does involving users in software development really influence system success? IEEE Software, 30(6):17–23, November/ December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Abdalkareem:2017:WDD

[ASR17] Rabe Abdalkareem, Emad Shihab, and Juergen Rilling. What do developers use the crowd for? A study using Stack Overflow. IEEE Software, 34(2):53–60, March/April 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020053-abs.
html.

Adams:2011:OSS

[AtHR11] Michael Adams, Arthur H. M. ter Hofstede, and Marcello La Rosa. Open source software for workflow management: The case of YAWL. IEEE Software, 28(3):16–19, May/ June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Avery:2011:EFM

[Ave11] David Avery. The evolution of flight management systems. IEEE Software, 28(1):11–13, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Alves:2019:EPS

[AVF19] C. Alves, G. Valença, and X. Franch. Exercising power in software ecosystems. IEEE Software, 36 (3):50–54, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Benaroch:2010:FPS

[BA10] Michel Benaroch and Ajit Appari. Financial pricing of software development risk factors. IEEE Software, 27(5):65–73, September/October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2012:AQR

[BAA^{+}12] Frank Buschmann, David Ameller, Claudia P. Ayala, Jordi Cabot, and Xavier Franch. Architecture quality revisited. IEEE Software, 29(4):22–24, July/ August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Balazs:2018:CRL

[Bal18] A. Balázs. A comet revisited: Lessons learned from Philaes landing. IEEE Software, 35(4):89– 93, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Basten:2017:G

[Bas17]  Dirk Basten.  Gamifi-

cation. IEEE Software, 34(5):76–81, September/ October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050076-abs.
html.

Bass:2018:SAD

[Bas18] Len Bass. The software architect and DevOps. IEEE Software, 35(1):8–10, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010008-abs.
html.

Bavani:2012:DAA

[Bav12] Raja Bavani. Distributed agile, agile testing, and technical debt. IEEE Software, 29(6):28–33, November/December 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Boehm:2010:P

[BB10] Barry Boehm and Kent Beck. Perspectives. IEEE Software, 27(4):26–29, July/August 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2012:CM

[BB12] Frank Buschmann and Jorg Bartholdt. Code matters! IEEE Software, 29(2):81–83, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bjarnason:2017:ART

[BB17] Elizabeth Bjarnason and Markus Borg. Aligning requirements and testing: Working together toward the same goal. IEEE Software, 34(1):20–23, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010020-abs.
html.

Basu:2011:RCB

[BBB^{+}11] Ananda Basu, Saddek Bensalem, Marius Bozga, Jacques Combaz, Mohamad Jaber, Thanh-Hung Nguyen, and Joseph Sifakis. Rigorous componentbased system design using the BIP framework. IEEE Software, 28(3):41– 48, May/June 2011. CO-DEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Basili:2018:SER

[BBB^{+}18] V. Basili, L. Briand, D. Bianculli, S. Nejati, F. Pastore, and M. Sabetzadeh. Software engineering research and industry: A symbiotic relationship to foster impact. IEEE Software, 35(5):44–49, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Barati:2019:PLR

[BBB^{+}19] S. Barati, F. A. Bartha, S. Biswas, R. Cartwright, A. Duracz, D. Fussell, H. Hoffmann, C. Imes, J. Miller, N. Mishra, Arvind, D. Nguyen, K. V. Palem, Y. Pei, K. Pingali, R. Sai, A. Wright, Y. Yang, and S. Zhang. Proteus: Language and runtime support for selfadaptive software development. IEEE Software, 36(2):73–82, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Becker:2016:RKS

[BBC^{+}16] Christoph Becker, Stefanie Betz, Ruzanna Chitchyan, Leticia Duboc, Steve M. Easterbrook, Birgit Penzenstadler, Norbet Seyff, and Colin C. Venters. Requirements: The key to sustainability. IEEE Software, 33(1):56–65, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010056-abs.
html.

Backes:2019:OCF

[BBC^{+}19] J. Backes, P. Bolignano, B. Cook, A. Gacek, K. S. Luckow, N. Rungta, M. Schaef, C. Schlesinger, R. Tanash, C. Varming, and M. Whalen. One-click formal methods. IEEE Software, 36(6):61– 65, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ball:2011:PPT

[BBdH^{+}11] Thomas Ball, Sebastian Burckhardt, Peli de Halleux, Madan Musuvathi, and Shaz Qadeer. Predictable and progressive testing of multithreaded code. IEEE Software, 28(3):75–83, May/ June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Basiri:2016:CE

[BBdR^{+}16] Ali Basiri, Niosha Behnam, Ruud de Rooij, Lorin Hochstein, Luke Kosewski, Justin Reynolds, and Casey Rosenthal. Chaos engineering. IEEE Software, 33(3):35–41, May/ June 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http:

                                        //csdl.computer.org/
                                        csdl/mags/so/2016/03/
                                        mso2016030035-abs.html.

Biddle:2019:III

[BBG19] R. Biddle, J. M. Brown, and S. Greenspan. From incident to insight: Incident responders and software innovation. IEEE Software, 36(1):56–62, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bertolino:2015:SEI

[BBM^{+}15] Antonia Bertolino, M. Brian Blake, Pankaj Mehra, Hong Mei, and Tao Xie.

Software engineering for Internet computing: Internetware and beyond [Guest Editors’ introduction]. IEEE Software, 32(1):35–37, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010035.html.

Bang:2018:CDC

[BBM18] Jae young Bang, Yuriy Brun, and Nenad Medvidović. Collaborativedesign conflicts: Costs and solutions. IEEE Software, 35(6):25–31, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Briand:2017:CCD

[BBN^{+}17] Lionel Briand, Domenico Bianculli, Shiva Nejati, Fabrizio Pastore, and Mehrdad Sabetzadeh. The case for context-driven software engineering research: Generalizability is overrated. IEEE Software, 34(5):72–75, September/ October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050072.html.

Baksa:2016:SCP

[BBP^{+}16] Attila Baksa, Andras Balazs, Zoltan Palos, Pater Spanyi, Sandor Szalai, and Laszlo Varhalmi. Software on a comet: The Philae Lander’s central onboard computer. IEEE Software, 33(2):81– 84, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        02/mso2016020081-abs.
                                        html.

Begel:2013:BSC

[BBS13a] Andrew Begel, Jan Bosch, and Margaret-Anne Storey. Bridging software communities through social networking. IEEE Software, 30(1):26–28, January/February 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Begel:2013:SNM

[BBS13b] Andrew Begel, Jan Bosch, and Margaret-Anne Storey. Social networking meets software development: Perspectives from GitHub, MSDN, Stack Exchange, and TopCoder. IEEE Software, 30(1):52–66, January/February 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Beecham:2017:PTS

[BCB^{+}17] Sarah Beecham, Tony Clear, John Barr, Mats Daniels, Michael Oudshoorn, and John Noll. Preparing tomorrow’s software engineers for work in a global environment. IEEE Software, 34(1): 9–12, January/February 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010009-abs.
html.

Beecham:2017:HBT

[BCD^{+}17] Sarah Beecham, Tony Clear, Daniela Damian, John Barr, John Noll, and Walt Scacchi. How best to teach global software engineering? educators are divided. IEEE Software, 34(1):16–19, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010016.html.

Bosch:2015:TSS

[BCH15] Jan Bosch, Rafael Capilla, and Rich Hilliard. Trends in systems and software variability [Guest Editors’ introduction]. IEEE Software, 32(3):44–51, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        03/mso2015030044.html.

Barbier:2015:SAC

[BCLP15] Franck Barbier, Eric Cariou, Olivier Le Goaer, and Samson Pierre. Software adaptation: Classification and a case study with state chart XML. IEEE Software, 32(5):68–76, September/October 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        05/mso2015050068-abs.
                                        html.

Bachmann:2013:IEE

[BCMN13] Felix H. Bachmann, Luis Carballo, James McHale, and Robert L. Nord. Integrate end to end early and often. IEEE Software, 30(4):9–14, July/ August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Braun:2016:PMA

[BCN16] Susanne Braun, Ralf Carbon, and Matthias Naab. Piloting a mobile-app ecosystem for smart farming. IEEE Software, 33(4): 9–14, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040009-abs.
html.

Babar:2010:MVS

[BCS10] Muhammad Ali Babar, Lianping Chen, and Forrest Shull. Managing variability in software product lines. IEEE Software, 27 (3):89–91, 94, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boden:2014:TSB

[BDD^{+}14] Alexander Boden, Christian Dorner, Sebastian Draxler, Volkmar Pipek, Gunnar Stevens, and Volker Wulf. Tangible and screen-based interfaces for end-user workflow modeling. IEEE Software, 31(4): 65–71, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bucchiarone:2018:MME

[BDD^{+}18] A. Bucchiarone, N. Dragoni, S. Dustdar, S. T. Larsen, and M. Mazzara. From monolithic to microservices: An experience report from the banking domain. IEEE Software, 35(3):50–55, May/June 2018. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Bellamy:2019:TYA

[BDH^{+}19] R. K. E. Bellamy, K. Dey, M. Hind, S. C. Hoffman, S. Houde, K. Kannan, P. Lohia, S. Mehta, A. Mojsilovic, S. Nagar, K. N. Ramamurthy, J. Richards, D. Saha, P. Sattigeri, M. Singh, K. R. Varshney, and Y. Zhang. Think your artificial intelligence software is fair? think again. IEEE Software, 36(4):76–80, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Brolund:2011:CPS

[BE11a] Daniel Brolund and Ola Ellnestam. Code pickup sticks. IEEE Software, 28(4):11–14, July/ August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buglione:2011:ETT

[BE11b] Luigi Buglione and Christof Ebert. Estimation tools and techniques. IEEE Software, 28(3):91–94, May/June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bell:2019:JBL

[Bel19] A. G. Bell. Jonathan Boccara on legacy code. IEEE Software, 36(5):80– 84, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bhamidipati:2015:PPM

[BFKZ15] Sandilya Bhamidipati, Nadia Fawaz, Branislav Kveton, and Amy Zhang. PriView: Personalized media consumption meets privacy against inference attacks. IEEE Software, 32(4):53–59, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040053-abs.
html.

Bellomo:2015:TAA

[BGK15] Stephany Bellomo, Ian Gorton, and Rick Kazman. Toward agile architecture: Insights from 15 years of ATAM data. IEEE Software, 32(5):38–45, September/October 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050038-abs.
html.

Buschmann:2010:FCSa

[BH10a] Frank Buschmann and Kevlin Henney. Five considerations for software architecture, part 1. IEEE Software, 27(3):63– 65, May/June 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2010:FCSb

[BH10b] Frank Buschmann and Kevlin Henney. Five considerations for software architecture, part 2. IEEE Software, 27(4): 12–14, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2013:AAM

[BH13] Frank Buschmann and Kevlin Henney. Architecture and agility: Married, divorced, or just good friends? IEEE Software, 30(2):80–82, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Badihi:2017:CAG

[BH17] Sahar Badihi and Abbas Heydarnoori. Crowd-Summarizer: Automated generation of code summaries for Java programs through crowdsourcing. IEEE Software, 34(2):71– 80, March/April 2017.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020071-abs.
html.

Bass:2019:TPO

[BH19] J. M. Bass and A. Haxby.

Tailoring product ownership in large-scale agile projects: Managing scale, distance, and governance. IEEE Software, 36(2):58–63, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bhatia:2018:NTS

[Bha18] K. Bhatia. Nate Taggart on serverless. IEEE Software, 35(4):101–104, July/ August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Baysal:2013:DDN

[BHG13] Olga Baysal, Reid Holmes, and Michael W. Godfrey. Developer dashboards: The need for qualitative analytics. IEEE Software, 30(4):46–52, July/August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Burger:2013:ACS

[BHH13] Stefan Burger, Oliver Hummel, and Matthias Heinisch. Airbus cabin software. IEEE Software, 30(1):21–25, January/February 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Balalaie:2016:MAE

[BHJ16] Armin Balalaie, Abbas Heydarnoori, and Pooyan Jamshidi. Microservices architecture enables DevOps: Migration to a cloud-native architecture. IEEE Software, 33(3):42– 52, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030042-abs.
html.

Baaz:2010:ALL

[BHN^{+}10] Anders Baaz, Lena Holmberg, Agneta Nilsson, Helena Olsson, and Anna Sandberg. Appreciating lessons learned. IEEE Software, 27(4):72–79, July/August 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Babb:2014:ERL

[BHN14] Jeffry Babb, Rashina Hoda, and Jacob Norbjerg. Embedding reflection and learning into agile software development.

IEEE Software, 31(4):51– 57, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bjarnason:2014:REB

[BHS^{+}14] Elizabeth Bjarnason, Anne Hess, Richard Berntsson Svensson, Bjorn Regnell, and Joerg Doerr. Reflecting on evidence-based timelines. IEEE Software, 31(4):37–43, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Barnes:2011:CCC

[BJ11] Nicholas Barnes and David Jones. Clear climate code: Rewriting legacy science software for clarity. IEEE Software, 28(6):36– 42, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bibi:2012:BAA

[BKB12] Stamatia Bibi, Dimitrios Katsaros, and Panayiotis Bozanis. Business application acquisition: Onpremise or SaaS-based solutions? IEEE Software, 29(3):86–93, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Breu:2014:NPS

[BKCF14] Ruth Breu, Annie Kuntzmann-Combelles, and Michael Felderer. New perspectives on software quality. IEEE Software, 31(1): 32–38, January/February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Berczuk:2010:WAT

[BL10] Steve Berczuk and Yi Lv. We’re all in this together. IEEE Software, 27(6):12– 15, November/December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Blankenship:2017:JDS

[Bla17] Marcus Blankenship. Josh Doody on salary negotiation for software engineers. IEEE Software, 34(2):117– 120, March/April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        02/mso2017020117.html.

Black:2018:NPJ

[Bla18a] N. Black. Nicolai Parlog on Java 9 modules. IEEE Software, 35(3):101– 104, May/June 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Black:2018:MPR

[Bla18b] Nate Black. Managing programmers, with Ron Lichty. IEEE Software, 35(1):117–120, January/ February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010117.html.

Blumen:2016:JLA

[Blu16] Robert Blumen. Jürgen Laartz and Alexander Budzier on why large IT projects fail. IEEE Software, 33(4):117–120, July/August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040117.html.

Blumen:2019:BSD

[Blu19] R. Blumen. Ben Sigelman on distributed tracing [software engineering radio]. IEEE Software, 36(1):98–101, January/ February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bell:2015:VFB

[BMDK15] Jonathan Bell, Eric Melski, Mohan Dattatreya, and Gail E. Kaiser.

Vroom: Faster build processes for Java. IEEE Software, 32(2):97–104, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020097-abs.
html.

Bener:2014:GS

[BMM14] Ayse Basar Bener, Maurizio Morisio, and Andriy Miranskyy. Green software. IEEE Software, 31 (3):36–39, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bener:2014:DPG

[BMR14] Ayse Basar Bener, Andriy Miranskyy, and Steve Raspudic. Deploying and provisioning green software. IEEE Software, 31 (3):76–78, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Benincasa:2014:ACM

[BMS^{+}14] Giacomo Benincasa, Alessandro Morelli, Cesare Stefanelli, Niranjan Suri, and Mauro Tortonesi. Agile communication middleware for next-generation mobile heterogeneous networks. IEEE Software, 31(2):54–61, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Brunet:2015:FYS

[BMSF15] Joao Brunet, Gail C. Murphy, Dalton Serey, and Jorge Figueiredo. Five years of software architecture checking: A case study of Eclipse. IEEE Software, 32(5):30–36, September/October 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050030-abs.
html.

Biro:2018:SSS

[BMSS18] Miklos Biro, Atif Mashkoor, Johannes Sametinger, and Remzi Seker. Software safety and security risk mitigation in cyberphysical systems. IEEE Software, 35(1):24–29, January/February 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010024.html.

Boden:2010:OSL

[BNW10] Alexander Boden, Bernhard Nett, and Volker Wulf. Operational and strategic learning in global software development. IEEE Software, 27(6):58–65, November/December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bosch:2017:TEB

[BO17] Jan Bosch and Helena H. Olsson. Toward evidencebased organizations: Lessons from embedded systems, online games, and the Internet of Things. IEEE Software, 34(5):60– 66, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/so/2017/
                                        05/mso2017050060-abs.
                                        html.

Booch:2010:AO

[Boo10a] Grady Booch. An architectural oxymoron. IEEE Software, 27(5):96, 95, September/October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2010:ASH

[Boo10b]  Grady Booch.  Archi-

tecture as a shared hallucination. IEEE Software, 27(1):96, 95, January/February 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2010:AR

[Boo10c]  Grady Booch.  Archi-

tecture reviews. IEEE Software, 27(3):96, 95, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2010:EBP

[Boo10d] Grady Booch. The elephant and the blind programmers. IEEE Software, 27(6):88, 87, November/ December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2010:EAT

[Boo10e] Grady Booch. Enterprise architecture and technical architecture. IEEE Software, 27(2):96, 95, March/ April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2010:SA

[Boo10f] Grady Booch. Systems architecture. IEEE Software, 27(4):96, 95, July/ August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2011:AJ

[Boo11a] Grady Booch. The architect’s journey. IEEE Software, 28(3):10–11, May/ June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2011:AST

[Boo11b] Grady Booch. The architecture of small things. IEEE Software, 28(6):18– 19, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2011:DD

[Boo11c]  Grady Booch.  Domi-

nant design. IEEE Software, 28(2):8–9, March/ April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2011:DMP

[Boo11d] Grady Booch. Draw me a picture. IEEE Software, 28(1):6–7, January/ February 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2011:SNW

[Boo11e] Grady Booch. The soul of a new Watson. IEEE Software, 28(4):9–10, July/ August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2011:UUT

[Boo11f]  Grady Booch.  Unin-

tentional and unbalanced transparency. IEEE Software, 28(5):12–13, September/October 2011.

CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Booch:2012:ATU

[Boo12a] Grady Booch. All things unseen. IEEE Software, 29 (3):14–16, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2012:FF

[Boo12b]  Grady Booch.  Facing

future. IEEE Software, 29(2):20–21, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2012:HE

[Boo12c] Grady Booch. The human experience. IEEE Software, 29(4):13–15, July/ August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2012:PA

[Boo12d] Grady Booch. The professional architect. IEEE Software, 29(1):12–13, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2012:TS

[Boo12e]  Grady Booch.  Three

stories. IEEE Software, 29(5):16–18, September/ October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2012:WLS

[Boo12f]  Grady Booch.  Woven

on the loom of sorrow. IEEE Software, 29(6):13– 15, November/December 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2013:DEM

[Boo13a] Grady Booch. Deus ex machina. IEEE Software, 30(6):14–16, November/December 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2013:MM

[Boo13b]  Grady Booch.  From

minecraft to minds. IEEE Software, 30(2):11–13, March/April 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2013:GTO

[Boo13c]  Grady Booch.  The

great and terrible Oz. IEEE Software, 30(1): 14–15, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2013:DB

[Boo13d] Grady Booch. In defense of boring. IEEE Software, 30 (3):16–17, May/June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2013:SP

[Boo13e] Grady Booch. The stories of possibility. IEEE Software, 30(5):14–15, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2013:WY

[Boo13f] Grady Booch. The wonder years. IEEE Software, 30(4):16–17, July/ August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2014:BQ

[Boo14a] Grady Booch. The big questions. IEEE Software, 31(4):9–11, July/ August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2014:HEA

[Boo14b] Grady Booch. The human and ethical aspects of big data. IEEE Software, 31(1):20–22, January/February 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2014:ILS

[Boo14c] Grady Booch. The incredible lightness of software. IEEE Software, 31(3):88, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2014:PM

[Boo14d]  Grady Booch.  The

persistence of memory. IEEE Software, 31(6):38– 40, November/December 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060038.html.

Booch:2014:SWT

[Boo14e] Grady Booch. The stories we tell ourselves. IEEE Software, 31(2):18– 20, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Booch:2014:CCQ

[Boo14f] Grady Booch. To code or not to code, that is the question. IEEE Software, 31(5):9–11, September/October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050009.html.

Booch:2015:AWM

[Boo15a]  Grady Booch.  All

watched over by machines of loving grace. IEEE Software, 32(2):19– 21, March/April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020019.html.

Booch:2015:AOC

[Boo15b] Grady Booch. Anarchy and order [on computing]. IEEE Software, 32 (3):20–22, May/June 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030020-abs.
html.

Booch:2015:OWO

[Boo15c] Grady Booch. I, for one, welcome our new computer overlords. IEEE Software, 32(6):8–10, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060008.html.

Booch:2015:MA

[Boo15d] Grady Booch. My autobiography. IEEE Software, 32(5):13–15, September/ October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050013-abs.
html.

Booch:2015:BBB

[Boo15e] Grady Booch. Of boilers, bit, and bots. IEEE Software, 32(1):11–13, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010011.html.

Booch:2015:TMO

[Boo15f] Grady Booch. Then a miracle occurs. IEEE Software, 32(4):12–14, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040012.html.

Booch:2016:CH

[Boo16a]  Grady Booch.  The

computational human. IEEE Software, 33(2):8– 10, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
02/mso2016020008.html.

Booch:2016:DD

[Boo16b]  Grady Booch.  Dis-

rupting the disrupters. IEEE Software, 33(4):6–8, July/August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040006-abs.
html.

Booch:2016:ICL

[Boo16c] Grady Booch. It is cold.

and lonely. IEEE Software, 33(3):7–9, May/June 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030007-abs.
html.

Booch:2016:NLM

[Boo16d] Grady Booch. No laughing matter. IEEE Software, 33(5):9–11, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050009-abs.
html.

Booch:2016:OUT

[Boo16e] Grady Booch. Once upon a time. IEEE Software, 33(6):8–10, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        06/mso2016060008-abs.
                                        html.

Booch:2016:RTP

[Boo16f]  Grady Booch.  Re-

membrance of things past. IEEE Software, 33(1):10–12, January/ February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2016/
                                        01/mso2016010010.html.

Booch:2018:HSE

[Boo18]  G. Booch.  The his-

tory of software engineering. IEEE Software, 35 (5):108–114, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bosch:2010:TCS

[Bos10] Jan Bosch. Toward compositional software product lines. IEEE Software, 27(3):29–34, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bosch:2016:SDE

[Bos16] Jan Bosch. Speed, data, and ecosystems: The future of software engineering. IEEE Software, 33(1):82–88, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010082-abs.
html.

Bouchard:2010:MSM

[Bou10] Lionel Bouchard. Multimedia software for mobile phones. IEEE Software, 27 (3):8–10, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Breaux:2014:PRA

[Bre14] Travis Breaux. Privacy requirements in an age of increased sharing. IEEE Software, 31(5):24–27, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050024.html.

Briand:2012:EES

[Bri12] Lionel Briand. Embracing the engineering side of software engineering. IEEE Software, 29(4):96, July/August 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Broy:2018:YTT

[Bro18] M. Broy. Yesterday, today, and tomorrow: 50 years of software engineering. IEEE Software, 35(5):38–43, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Brosgol:2019:HSS

[Bro19] B. M. Brosgol. How to succeed in the software business while giving away the source code: The AdaCore experience. IEEE Software, 36(6):17–22, November/December 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Brant:2015:RTT

[BS15] John Brant and Friedrich Steimann. Refactoring tools are trustworthy enough and trust must be earned. IEEE Software, 32(6):80–83, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        06/mso2015060080.html.

Britto:2016:SAL

[BSD16] Ricardo Britto, Darja Smite, and Lars-Ola Damm. Software architects in large-scale distributed projects: An Ericsson case study. IEEE Software, 33(6):48–55, November/ December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060048-abs.
html.

Broring:2017:EIE

[BSS^{+}17] Arne Broring, Stefan Schmid, Corina-Kim Schindhelm, Abdelmajid Khelil, Sebastian Kabisch, Denis Kramer, Danh Le Phuoc, Jelena Mitic, Darko Anicic, and Ernest Teniente. Enabling IoT ecosystems through platform interoperability. IEEE Software, 34(1):54–61, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010054-abs.
html.

Blincoe:2019:PGD

[BSW19] K. Blincoe, O. Springer, and M. R. Wrobel. Perceptions of gender diversity’s impact on mood in software development teams. IEEE Software, 36(5):51–56, September/ October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Budgen:2013:DPM

[Bud13] David Budgen. Design patterns: Magic or myth? IEEE Software, 30(2):87– 90, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Buschmann:2010:LFPa

[Bus10a] Frank Buschmann. Learning from failure, part 2: Featuritis, performitis, and other diseases. IEEE Software, 27(1): 10–11, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2010:LFPb

[Bus10b] Frank Buschmann. Learning from failure, Part III: On hammers and nails, and falling in love with technology and design. IEEE Software, 27(2):49– 51, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2010:ASP

[Bus10c] Frank Buschmann. On architecture styles and paradigms. IEEE Software, 27(5):92–94, September/October 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2010:VFS

[Bus10d] Frank Buschmann. Valuefocused system quality. IEEE Software, 27(6):84– 86, November/December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2011:GYAa

[Bus11a] Frank Buschmann. Gardening your architecture, part 1: Refactoring. IEEE Software, 28(4): 92–94, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2011:GYAb

[Bus11b] Frank Buschmann. Gardening your architecture, part 2: Reengineering and rewriting. IEEE Software, 28(5):21–23, September/ October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2011:TAB

[Bus11c] Frank Buschmann. Tests:

The architect’s best friend. IEEE Software, 28(3):7–9, May/June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2011:PPT

[Bus11d] Frank Buschmann. To pay or not to pay technical debt. IEEE Software, 28(6):29–31, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2011:USUa

[Bus11e] Frank Buschmann. Unusable software is useless, part 1. IEEE Software, 28(1):92–94, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2011:USUb

[Bus11f] Frank Buschmann. Unusable software is useless, part 2. IEEE Software, 28 (2):100–102, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Buschmann:2012:BGW

[Bus12a] Frank Buschmann. To boldly go where no one has gone before. IEEE Software, 29(1):23–25, January/February 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Buschmann:2012:WLA

[Bus12b] Frank Buschmann. A week in the life of an architect. IEEE Software, 29 (3):94–96, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Buschmann:2013:IR

[Bus13] Frank Buschmann. Innovation reconsidered. IEEE Software, 30(1):18–20, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bouwers:2010:LSC

[BvD10] Eric Bouwers and Arie van Deursen. A lightweight sanity check for implemented architectures. IEEE Software, 27(4):44– 50, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Baker:2012:GEI

[BvdHOP12] Alex Baker, Andre van der Hoek, Harold Ossher, and Marian Petre. Guest Editors’ introduction: Studying professional software design. IEEE Software, 29(1):28–33, January/February 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bull:2014:SRP

[BW14] Christopher N. Bull and Jon Whittle. Supporting reflective practice in software engineering education through a studiobased approach. IEEE Software, 31(4):44–50, July/August 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bashroush:2017:APE

[BW17] Rabih Bashroush and Eoin Woods. Architectural principles for energyaware Internet-scale applications. IEEE Software, 34 (3):14–17, May/June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030014-abs.
html.

Blair:2010:RDA

[BWC10] Stuart Blair, Richard Watt, and Tim Cull. Responsibility-driven architecture. IEEE Software, 27(2):26–32, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bajwa:2017:SUM

[BWD^{+}17] Sohaib Shahid Bajwa, Xiaofeng Wang, Anh Nguyen Duc, Rafael Matone Chanin, Rafael Prikladnicki, Leandro Bento Pompermaier, and Pekka Abrahamsson. Start-ups must be ready to pivot. IEEE Software, 34 (3):18–22, May/June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030018-abs.
html.

Bashroush:2016:DCE

[BWN16] Rabih Bashroush, Eoin Woods, and Adel Noureddine. Data center energy demand: What got us here won’t get us there. IEEE Software, 33(2):18– 21, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
02/mso2016020018-abs.
html.

Bano:2018:UIS

[BZd18] M. Bano, D. Zowghi, and F. da Rimini. User involvement in software development: The good, the bad, and the ugly. IEEE Software, 35(6):8–11, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cabot:2018:WCM

[Cab18] J. Cabot. WordPress: A content management system to democratize publishing. IEEE Software, 35 (3):89–92, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Capretz:2014:BHF

[Cap14] Luiz Fernando Capretz. Bringing the human factor to software engineering. IEEE Software, 31(2): 104, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carter:2017:FRD

[Car17a] Kim Carter. François Raynaud on DevSecOps. IEEE Software, 34(5):93–96, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        05/mso2017050093.html.

Carver:2017:DRE

[Car17b] Jeffrey C. Carver. Developments in requirements engineering. IEEE Software, 34(4):8–10, July/ August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        04/mso2017040008.html.

Carver:2019:QNS

[Car19] J. Carver. Quality, nontechnical skills, blind programmers, and deep learning. IEEE Software, 36 (2):127–136, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Clements:2010:BGV

[CB10] Paul Clements and Len Bass. The business goals viewpoint. IEEE Software, 27(6):38–45, November/December 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chen:2013:CAS

[CBN13] Lianping Chen, Muhammad Ali Babar, and Bashar Nuseibeh. Characterizing architecturally significant requirements. IEEE Software, 30(2):38– 45, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Conboy:2019:ILS

[CC19] K. Conboy and N. Carroll.

Implementing large-scale agile frameworks: Challenges and recommendations. IEEE Software, 36(2):44–50, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cabot:2019:MSA

[CCC^{+}19] J. Cabot, R. Capilla, C. Carrillo, H. Muccini, and B. Penzenstadler.

Measuring systems and architectures: A sustainability perspective. IEEE Software, 36(3):98–100, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2017:GTD

[CCCM17] Jeffrey C. Carver, Jordi Cabot, Rafael Capilla, and Henry Muccini. GitHub, technical debt, code formatting, and more. IEEE Software, 34(2):105–107, March/April 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020105.html.

Ciccozzi:2017:MDE

[CCD^{+}17] Federico Ciccozzi, Ivica Crnkovic, Davide Di Ruscio, Ivano Malavolta, Patrizio Pelliccione, and Romina Spalazzese. Modeldriven engineering for mission-critical IoT systems. IEEE Software, 34(1):46–53, January/ February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010046-abs.
html.

Carver:2015:CCR

[CCH^{+}15] Jeffrey C. Carver, Bora Caglayan, Mayy Habayeb, Birgit Penzenstadler, and Aiko Yamashita. Collaborations and code reviews. IEEE Software, 32(5):27–29, September/ October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050027.html.

Carver:2016:RTS

[CCMT16] Jeffrey C. Carver, Jordi Cabot, Leandro L. Minku, and Marco Torchiano. Regression testing, spoken language, crash-inducing commits, UML, and legal policy. IEEE Software, 33(2):26–28, March/ April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http:

//csdl.computer.org/
csdl/mags/so/2016/02/
mso2016020026.html.

Carver:2018:GSE

[CCP^{+}18] J. Carver, R. Capilla, B. Penzenstadler, A. Serebrenik, and A. Valdezate. Gender, sentiment and emotions, and safetycritical systems. IEEE Software, 35(6):16–19, November/December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cabiddu:2011:TPA

[CCS^{+}11] Giovanni Cabiddu, Emanuele Cesena, Roberto Sassu, Davide Vernizzi, Gianluca Ramunno, and Antonio Lioy. The Trusted Platform Agent. IEEE Software, 28(2):35–41, March/ April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Choi:2010:MFI

[CCSW10] Namjoo Choi, Indushobha Chengalur-Smith, and Andrew Whitmore. Managing first impressions of new open source software projects. IEEE Software, 27(6):73–77, November/December 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Conboy:2011:PPK

[CCWP11] Kieran Conboy, Sharon Coyle, Xiaofeng Wang, and Minna Pikkarainen. People over process: Key challenges in agile development. IEEE Software, 28(4):48–57, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Carver:2016:PLE

[CdAC^{+}16] Jeffrey C. Carver, Eduardo Santana de Almeida, Rafael Capilla, Leandro L. Minku, Henry Muccini, and Birgit Penzenstadler. Product lines, energy conservation, use cases, agile development, and infotainment. IEEE Software, 33 (3):29–31, May/June 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030029.html.

Calefato:2019:ACD

[CE19] F. Calefato and C. Ebert.

Agile collaboration for distributed teams [software technology]. IEEE Software, 36(1):72–78, January/February 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carrozza:2013:EAT

[CFF^{+}13] Gabriella Carrozza, Mauro Faella, Francesco Fucci, Roberto Pietrantuono, and Stefano Russo. Engineering air traffic control systems with a modeldriven approach. IEEE Software, 30(3):42–48, May/June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Camara:2019:BLL

[CGE19] J. Camara, D. Garlan, and G. Eakman. Building long-lived adaptive systems. IEEE Software, 36(2):70–72, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cano:2013:SIS

[CH13a] Jesús Cano and Roberto Hernandez. SCEPYLT: An information system for fighting terrorism. IEEE Software, 30(3):73– 79, May/June 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cleland-Huang:2013:RAK

[CH13b] Jane Cleland-Huang. Are requirements alive and kicking? IEEE Software, 30(3):13–15, May/June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cleland-Huang:2013:MEP

[CH13c] Jane Cleland-Huang. Meet Elaine: A persona-driven approach to exploring architecturally significant requirements. IEEE Software, 30(4):18–21, July/ August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cleland-Huang:2013:TAQ

[CH13d] Jane Cleland-Huang. Thinking about quoins: Strategic traceability of architecturally significant requirements. IEEE Software, 30(5):16–18, September/ October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cleland-Huang:2014:DFA

[CH14a] Jane Cleland-Huang. Don’t fire the architect! Where were the requirements? IEEE Software, 31(2):27– 29, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cleland-Huang:2014:HWD

[CH14b] Jane Cleland-Huang. How well do you know your personae non gratae? IEEE Software, 31(4):28– 31, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cleland-Huang:2015:IVT

[CH15a] Jane Cleland-Huang. Injecting value-thinking into prioritization decisions. IEEE Software, 32(2):14– 18, March/April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.

org/csdl/mags/so/2015/
02/mso2015020014-abs.
html.

Cleland-Huang:2015:MDK

[CH15b] Jane Cleland-Huang. Mining domain knowledge [requirements]. IEEE Software, 32(3):16–19, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030016-abs.
html.

Cleland-Huang:2015:TMI

[CH15c] Jane Cleland-Huang. Toward meaningful industrial– academic partnerships. IEEE Software, 32(1): 18–21, January/February 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010018-abs.
html.

Cleland-Huang:2016:RRS

[CH16a] Jane Cleland-Huang. Requirements that reflect social responsibility. IEEE Software, 33(1):109–111, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010109.html.

Cleland-Huang:2016:SP

[CH16b] Jane Cleland-Huang. Stakeholders on the prowl. IEEE Software, 33(2):29– 31, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
02/mso2016020029.html.

Cleland-Huang:2017:SSA

[CH17] Jane Cleland-Huang. Safety stories in agile development. IEEE Software, 34(4):16–19, July/ August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040016-abs.
html.

Cleland-Huang:2016:KAO

[CHDK^{+}16] Jane Cleland-Huang, Tamara Denning, Tadayoshi Kohno, Forrest Shull, and Samuel Weber. Keeping ahead of our adversaries. IEEE Software, 33(3):24–28, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030024.html.

Chernak:2014:DMC

[Che14] Yuri Chernak. Dataflow modeling with crosscutting concerns and a concept lattice. IEEE Software, 31(6):70–78, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060070-abs.
html.

Chen:2015:CDH

[Che15]  Lianping Chen.  Con-

tinuous delivery: Huge benefits, but challenges too. IEEE Software, 32(2): 50–54, March/April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020050-abs.
html.

Cleland-Huang:2013:TPR

[CHHSM13] Jane Cleland-Huang, Robert S. Hanmer, Sam Supakkul, and Mehdi Mirakhorli. The twin peaks of requirements and architecture. IEEE Software, 30(2):24–29, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cleland-Huang:2014:RGW

[CHL14] Jane Cleland-Huang and Paula Laurent. Requirements in a global world. IEEE Software, 31(6):34–37, November/ December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060034-abs.
html.

Chen:2016:SPD

[CKH16] Hong-Mei Chen, Rick Kazman, and Serge Haziyev. Strategic prototyping for developing big data systems. IEEE Software, 33(2):36–43, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL http://csdl.computer.
org/csdl/mags/so/2016/
02/mso2016020036-abs.
html.

Clements:2010:CSA

[Cle10] Paul Clements. Certified software architects. IEEE Software, 27(6):6– 8, November/December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Canovas:2010:ADM

[CM10] Javier Canovas and Jesus Molina. An architecturedriven modernization tool for calculating metrics. IEEE Software, 27(4):37– 43, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cormode:2012:ADC

[CM12] Graham Cormode and Muthu Muthukrishnan. Approximating data with the count-min sketch. IEEE Software, 29(1): 64–69, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2019:HIS

[CM19] J. C. Carver and L. L. Minku. Highlights from ICSE 2019: Software security and mobile app energy consumption. IEEE Software, 36(5):29–31, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Crichton:2012:SSO

[CMC^{+}12] Daniel J. Crichton, Chris A. Mattmann, Luca Cinquini, Amy Braverman, Duane Waliser, Michael Gunson, Andrew F. Hart, Cameron E. Goodale, Peter Lean, and Jinwon Kim. Sharing satellite observations with the climate-modeling community: Software and architecture. IEEE Software, 29(5):73–81, September/ October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cantor:2016:SSD

[CMM16] Murray Cantor, Bruce MacIsaac, and Rosanna Mannan. Steering software development workflow: Lessons from the Internet. IEEE Software, 33(5):96–102, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050096-abs.
html.

Carver:2017:RHV

[CMP17] Jeffrey C. Carver, Leandro L. Minku, and Birgit Penzenstadler. Requirements, human values, and the development technology landscape. IEEE Software, 34(1):13–15, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010013.html.

Chapresto:2011:CAR

[CMS11] Eduardo Jimenez Chapresto, Kenny Mitchell, and Francisco Jose Seron. Capture and analysis of racing gameplay metrics. IEEE Software, 28(5):46– 52, September/October 2011. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic).

Carver:2017:DTD

[CMY17] Jeffrey C. Carver, Henry Muccini, and Aiko Yamashita. Distributed teams, developer participation, and more. IEEE Software, 34(3):114–116, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        03/mso2017030114.html.

Czerwonka:2013:CBS

[CNSM13] Jacek Czerwonka, Nachiappan Nagappan, Wolfram Schulte, and Brendan Murphy. CODEM-INE: Building a software development data analytics platform at Microsoft. IEEE Software, 30(4):64– 71, July/August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Capilla:2017:TAK

[CNZC17] Rafael Capilla, Elisa Yumi Nakagawa, Uwe Zdun, and Carlos Carrillo. Toward architecture knowledge sustainability: Extending system longevity. IEEE Software, 34(2):108– 111, March/April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020108-abs.
html.

Conroy:2012:TDW

[Con12] Patrick Conroy. Technical debt: Where are the shareholders’ interests? IEEE Software, 29(6):88, November/ December 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cook:2010:CRC

[Coo10] Mary Rose Cook. Creative requirements conversations. IEEE Software, 27(2):90–91, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Correa:2013:HAM

[Cor13] Bett Correa. How are architects made? IEEE Software, 30(5):11–13, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2018:IAC

[CP18] J. C. Carver and R. Prikladnicki. Industry academia collaboration in software engineering. IEEE Software, 35(5):120–124, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Casino:2017:HRS

[CPB^{+}17] Fran Casino, Constantinos Patsakis, Edgar Batista, Frederic Borras, and Antoni Martinez-Balleste. Healthy routes in the smart city: A contextaware mobile recommender. IEEE Software, 34(6):42– 47, November/December 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060042.html.

Christou:2010:UAU

[CPP10] Ioannis Christou, Stavros Ponis, and Eleni Palaiologou. Using the agile unified process in banking. IEEE Software, 27 (3):72–79, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Carver:2016:PQP

[CPP16] Jeffrey C. Carver, Maria Paasivaara, and Birgit Penzenstadler. Probing questions, participatory democracy, quality assurance, and customer data. IEEE Software, 33(5):12–14, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050012.html.

Carver:2018:SAE

[CPS18] J. Carver, B. Penzenstadler, and A. Serebrenik. Software analysis, evolution, and reengineering, and ICT sustainability. IEEE Software, 35(4): 78–80, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2017:HF

[CPSY17] Jeffrey C. Carver, Birgit Penzenstadler, Alexander Serebrenik, and Aiko Yamashita. The human factor. IEEE Software, 34(5):90–92, September/ October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050090.html.

Clune:2011:STV

[CR11] Thomas Clune and Richard Rood. Software testing and verification in climate model development. IEEE Software, 28(6):49– 55, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Cantor:2014:EGS

[CR14] Murray Cantor and Walker Royce. Economic governance of software delivery. IEEE Software, 31(1): 54–61, January/February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Crabb:2014:BCC

[Cra14] Joel Crabb. The BestBuy.com cloud architecture. IEEE Software, 31(2):91–96, March/April 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Callanan:2016:DMI

[CS16] Matt Callanan and Alexandra Spillane. DevOps: Making it easy to do the right thing. IEEE Software, 33(3):53–59, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        03/mso2016030053-abs.
                                        html.

Carver:2017:IH

[CS17] Jeffrey C. Carver and Alexander Serebrenik. ICSE highlights. IEEE Software, 34(6):18–20, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060018.html.

Carver:2018:SME

[CS18] J. C. Carver and A. Serebrenik. Software maintenance and evolution and automated software engineering. IEEE Software, 35(2):102–104, March/ April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2019:GSE

[CS19] J. C. Carver and A. Serebrenik. Gender in software engineering. IEEE Software, 36(6):76–78, November/December 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2018:ESE

[CSC^{+}18] J. C. Carver, E. Santana de Almeida, R. Capilla, L. Minku, M. Torchiano, and A. Valdezate. Empirical software engineering, predictive models, and product lines. IEEE Software, 35(3):8–11, May/ June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Crnkovic:2011:SCB

[CSS11] Ivica Crnkovic, Judith Stafford, and Clemens Szyperski. Software components beyond programming: From routines to services. IEEE Software, 28(3):22–26, May/June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Curtis:2012:EPA

[CSS12] Bill Curtis, Jay Sappidi, and Alexandra Szynkarski. Estimating the principal of an application’s technical debt. IEEE Software, 29(6):34–42, November/December 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Capiluppi:2013:ATC

[CSS13] Andrea Capiluppi, Alexander Serebrenik, and Leif Singer. Assessing technical candidates on the Social Web. IEEE Software, 30(1):45–51, January/February 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ceccato:2011:CRS

[CT11] Mariano Ceccato and Paolo Tonella. CodeBender: Remote software protection using orthogonal replacement. IEEE Software, 28(2):28–34, March/ April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Callau:2013:PG

[CT13] Oscar Callaú and Éric Tanter. Programming with ghosts. IEEE Software, 30(1):74–80, January/February 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cervantes:2013:PWU

[CVEK13] Humberto Cervantes, Perla Velasco-Elizondo, and Rick Kazman. A principled way to use frameworks in architecture design. IEEE Software, 30(2):46–53, March/ April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2015:SQE

[CYM^{+}15] Jeffrey C. Carver, Aiko Yamashita, Leandro Minku, Mayy Habayeb, and Sedef Akinli Kocak. Software quality, energy awareness, and more. IEEE Software, 32(6):98–100, November/ December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060098.html.

delBianco:2011:SOS

[dBLMT11] Vieri del Bianco, Luigi Lavazza, Sandro Morasca, and Davide Taibi. A survey on open source software trustworthiness. IEEE Software, 28(5):67– 75, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Debreceni:2018:SVC

[DBRV18] C. Debreceni, G. Bergmann, I. Ráth, and D. Varró. Secure views for collaborative modeling. IEEE Software, 35(6):32–38, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Djuric:2010:MPI

[DD10] Dragan Djuric and Vladan Devedzic. Magic potion: Incorporating new development paradigms through metaprogramming. IEEE Software, 27(5):38–44, September/ October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dilmaghani:2012:SES

[DD12] Ania Dilmaghani and Jim Dibble. Strategies for early-stage collaborative design. IEEE Software, 29(1):39–45, January/February 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

DiRocco:2015:CRM

[DDIP15] Juri Di Rocco, Davide Di Ruscio, Ludovico Iovino, and Alfonso Pierantonio. Collaborative repositories in model-driven engineering [software technology]. IEEE Software, 32(3):28– 34, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030028-abs.
html.

DeMarco:2011:ALP

[DeM11] Tom DeMarco. All late projects are the same. IEEE Software, 28(6):104, 103, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

DeMarco:2013:BWP

[DeM13]  Tom DeMarco.  Bells,

whistles, power, and the requirements process. IEEE Software, 30(4):104, July/August 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

DeMarco:2014:SBE

[DeM14] Tom DeMarco. Sigil, Blue-Griffon, and the evolving software market. IEEE Software, 31(4):100, July/ August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dingsoyr:2016:TPS

[DFD^{+}16] Torgeir Dingsoyr, Tor Erlend Faegri, Tore Dyba, Borge Haugset, and Yngve Lindsjorn. Team performance in software development: Research results versus agile principles. IEEE Software, 33(4):106– 110, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040106-abs.
html.

Dalpiaz:2018:NLP

[DFFP18] F. Dalpiaz, A. Ferrari, X. Franch, and C. Palomares. Natural language processing for requirements engineering: The best is yet to come. IEEE Software, 35(5):115–119, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dingsoeyr:2019:ADS

[DFP19] T. Dingsoeyr, D. Falessi, and K. Power. Agile development at scale: The next frontier. IEEE Software, 36(2):30–38, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

deGea:2011:RET

[dGNA^{+}11] Juan M. Carrillo de Gea, Joaquin Nicolas, Jose L. Fernandez Aleman, Ambrosio Toval, Christof Ebert, and Aurora Vizcaino. Requirements engineering tools. IEEE Software, 28(4):86–91, July/ August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Daneva:2015:CQR

[DHB15] Maya Daneva, Andrea Herrmann, and Luigi Buglione. Coping with quality requirements in large, contract-based projects. IEEE Software, 32(6):84–91, November/ December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060084-abs.
html.

Dekhtyar:2019:RER

[DHH^{+}19] A. Dekhtyar, J. Huffman Hayes, I. Hadar, E. Combs, A. Ferrari, S. Gregory, J. Horkoff, M. Levy, M. Nayebi, B. Paech, J. Payne, M. Primrose, P. Spoletini, S. Clarke, C. Brophy, D. Amyot, W. Maalej, G. Ruhe, J. Cleland-Huang, and D. Zowghi. Requirements engineering (RE) for social good: RE cares [requirements]. IEEE Software, 36(1): 86–94, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dig:2011:RAP

[Dig11] Danny Dig. A refactoring approach to parallelism. IEEE Software, 28(1): 17–22, January/February 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dig:2015:RAE

[Dig15] Danny Dig. Refactoring for asynchronous execution on mobile devices. IEEE Software, 32(6):52–61, November/ December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        06/mso2015060052-abs.
                                        html.

DiRuscio:2012:CEM

[DIP12] Davide Di Ruscio, Ludovico Iovino, and Alfonso Pierantonio. Coupled evolution in model-driven engineering. IEEE Software, 29(6):78–84, November/December 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Denault:2011:JMM

[DK11] Alexandre Denault and Jorg Kienzle. Journey: a massively multiplayer online game middleware. IEEE Software, 28(5):38– 44, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Duboc:2010:DTS

[DLR10] Leticia Duboc, Emmanuel Letier, and David S. Rosenblum. Death, taxes, & scalability. IEEE Software, 27(4):20–21, July/ August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dyba:2014:RSE

[DMG14] Tore Dyba, Neil Maiden, and Robert Glass. The reflective software engineer: Reflective practice. IEEE Software, 31(4):32– 36, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

deMoura:2016:WSI

[dMHSC16] Edleno Silva de Moura, Mauro Rojas Herrera, Leonardo Santos, and Tayana Conte. When software impacts the economy and environment. IEEE Software, 33(6):23– 26, November/December 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060023-abs.
html.

Neto:2011:TSP

[dMSNRdCM^{+}11] Paulo Anselmo da Mota Silveira Neto, Per Runeson, Ivan do Carmo Machado, Eduardo Santana de Almeida, Silvio Romero de Lemos Meira, and Emelie Engstrom. Testing software product lines. IEEE Software, 28(5):16–20, September/ October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

deOliveira:2016:ELL

[dOdA16] Raphael Pereira de Oliveira and Eduardo Santana de Almeida. Evaluating Lehman’s laws of software evolution for software product lines. IEEE Software, 33(3):90– 93, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/

03/mso2016030090-abs.
html.

Melo:2016:BEF

[dOMFP16] Claudia de O. Melo, Ronaldo Ferraz, and Rebecca J. Parsons. Brazil and the emerging future of software engineering. IEEE Software, 33(1):45–47, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010045.html.

Dornenburg:2018:PD

[Dör18]  E. Dörnenburg.  The

path to DevOps. IEEE Software, 35(5):71–75, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Duarte:2014:SRV

[DPCL14] Tiago Duarte, Rafael Prikladnicki, Fabio Calefato, and Filippo Lanubile. Speech recognition for voice-based machine translation. IEEE Software, 31(1):26–31, January/ February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dybaa:2012:WEL

[DS12] Tore Dybå and Helen Sharp. What’s the evidence for lean? IEEE Software, 29(5):19–21, September/October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dhote:2013:PTC

[DS13] Manish Rajendra Dhote and G. G. Sarate. Performance testing complexity analysis on Ajaxbased Web applications. IEEE Software, 30(6):70– 74, November/December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Azevedo:2014:AAA

[dSAPW^{+}14] Luis da Silva Azevedo, David Parker, Martin Walker, Yiannis Papadopoulos, and Rui Esteves Araujo. Assisted assignment of automotive safety requirements. IEEE Software, 31(1):62–68, January/February 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Debski:2018:SRA

[DSM^{+}18] A. Debski, B. Szczepanik, M. Malawski, S. Spahr, and D. Muthig. A scalable, reactive architecture for cloud applications. IEEE Software, 35(2):62–71, March/April 2018. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

daSilva:2016:HSD

[dSMdSC16] Fabio Q. B. da Silva, Cleviton V. F. Monteiro, Igor Ebrahim dos Santos, and Luiz Fernando Capretz. How software development group leaders influence team members’ innovative behavior. IEEE Software, 33(5):106–109, September/October 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050106-abs.
html.

Dabbish:2013:LT

[DSTH13] Laura Dabbish, Colleen Stuart, Jason Tsay, and James Herbsleb. Leveraging transparency. IEEE Software, 30(1):37–43, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dullemond:2014:CSV

[DvGvS14] Kevin Dullemond, Ben van Gameren, and Rini van Solingen. Collaboration spaces for virtual software teams. IEEE Software, 31(6):47–53, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        06/mso2014060047-abs.
                                        html.

Dybaa:2013:CEE

[Dyb13] Tore Dybå. Contextualizing empirical evidence. IEEE Software, 30(1): 81–83, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Devanbu:2018:BEH

[DZB18] P. Devanbu, T. Zimmermann, and C. Bird. Belief and evidence: How software engineers form their opinions. IEEE Software, 35(6):72–76, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Edmondson:2014:NGM

[EAG^{+}14] James Edmondson, William Anderson, Jeff Gray, Joseph P. Loyall, Klaus Schmid, and Jules White. Next-generation mobile computing. IEEE Software, 31(2):44–47, March/ April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2012:LSD

[EAO12] Christof Ebert, Pekka Abrahamsson, and Nilay Oza. Lean software development. IEEE Software, 29(5):22–25, September/ October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2014:SPM

[Ebe14]  Christof Ebert.  Soft-

ware product management. IEEE Software, 31 (3):21–24, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2015:IFS

[Ebe15a] Christof Ebert. Implementing functional safety. IEEE Software, 32(5):84– 89, September/October 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050084.html.

Ebert:2015:LF

[Ebe15b] Christof Ebert. Looking into the future. IEEE Software, 32(6):92–97, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060092.html.

Ebert:2018:YSE

[Ebe18] C. Ebert. 50 years of software engineering: Progress and perils. IEEE Software, 35(5):94–101, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Eito-Brun:2016:IDS

[EBS16] Ricardo Eito-Brun and Miguel-Angel Sicilia. Innovationdriven software development: Leveraging small companies’ product-development capabilities. IEEE Software, 33(5):38–46, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050038-abs.
html.

Ebert:2016:CC

[EC16] Christof Ebert and James Cain. Cyclomatic complexity. IEEE Software, 33(6):27–29, November/ December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060027-abs.
html.

Ebert:2017:TST

[EC17] Christof Ebert and Steve Counsell. Toward software technology 2050. IEEE Software, 34(4):82– 88, July/August 2017.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040082-abs.
html.

Ebert:2018:DT

[ED18] C. Ebert and C. H. C.

Duarte. Digital transformation. IEEE Software, 35(4):16–21, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2019:CEI

[ED19] C. Ebert and A. Dubey.

Convergence of enterprise IT and embedded systems. IEEE Software, 36(3):92– 97, May/June 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Easterbrook:2011:GEI

[EEBB11] Steve M. Easterbrook, Paul N. Edwards, Venkatramani Balaji, and Reinhard Budich. Guest Editors’ introduction: Climate change — science and software. IEEE Software, 28(6):32–35, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2017:AS

[EF17] Christof Ebert and John Favaro. Automotive software. IEEE Software, 34 (3):33–39, May/June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030033.html.

Erdweg:2014:ESS

[EFO14] Sebastian Erdweg, Stefan Fehrenbach, and Klaus Ostermann. Evolution of software systems with extensible languages and DSLs. IEEE Software, 31(5):68–75, September/ October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050068-abs.
html.

Ebert:2016:D

[EGHS16] Christof Ebert, Gorka Gallardo, Josune Hernantes, and Nicolas Serrano. DevOps. IEEE Software, 33(3):94–100, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030094.html.

Ebert:2019:DST

[EHMFT19] C. Ebert, J. Heidrich, S. Martinez-Fernandez, and A. Trendowicz. Data science: Technologies for better software. IEEE Software, 36(6):66–72, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2015:WNA

[EHS15] Christof Ebert, Gerd Hoefner, and Mani V. S. What next? Advances in software-driven industries. IEEE Software, 32(1):22–28, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010022.html.

Ebert:2012:RSR

[EJ12] Christof Ebert and Michael Jastram. ReqIF: Seamless requirements interchange format between business partners. IEEE Software, 29(5):82–87, September/ October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2016:GSE

[EKP16] Christof Ebert, Marco Kuhrmann, and Rafael Prikladnicki. Global software engineering: An industry perspective. IEEE Software, 33(1):105–108, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2016/
                                        01/mso2016010105.html.

Eldh:2015:COP

[EM15] Sigrid Eldh and Brendan Murphy. Code ownership perspectives. IEEE Software, 32(6):18–19, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        06/mso2015060018.html.

Erdogmus:2018:CBB

[EM18] H. Erdogmus and N. Medvidović. A conversation with Barry Boehm: Recollections from 50 years of software engineering. IEEE Software, 35(5):14– 19, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Erdogmus:2018:YSE

[EMP18] H. Erdogmus, N. Medvidović, and F. Paulisch. 50 years of software engineering. IEEE Software, 35(5):20–24, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Engelfriet:2010:COS

[Eng10] Arnoud Engelfriet. Choosing an open source license. IEEE Software, 27(1): 48–49, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erder:2016:WAR

[EP16] Murat Erder and Pierre Pureur. What’s the architect’s role in an agile, cloud-centric world? IEEE Software, 33(5):30–33, September/October 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050030-abs.
html.

Ebert:2017:SA

[EP17a] Christof Ebert and Maria Paasivaara. Scaling agile. IEEE Software, 34 (6):98–103, November/ December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060098-abs.
html.

Erder:2017:WTP

[EP17b] Murat Erder and Pierre Pureur. What type of people are software architects?

IEEE Software, 34(4):20– 22, July/August 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        04/mso2017040020-abs.
                                        html.

Erdogmus:2010:CAS

[Erd10a] Hakan Erdogmus. Can all sequential processes grow up to be iterative and incremental? IEEE Software, 27(4):2–5, July/ August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2010:DVL

[Erd10b] Hakan Erdogmus. Deja vu: The life of software engineering ideas. IEEE Software, 27(1):2–5, January/ February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2010:HIE

[Erd10c] Hakan Erdogmus. How important is evidence, really? IEEE Software, 27(3):2–5, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2010:PR

[Erd10d] Hakan Erdogmus. Passing the reins. IEEE Software, 27(6):3–5, November/December 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2010:RPS

[Erd10e] Hakan Erdogmus. Regress or progress? Seeing good software engineering ideas through. IEEE Software, 27(2):4–7, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Erdogmus:2010:TPT

[Erd10f] Hakan Erdogmus. Tracking progress through earned value. IEEE Software, 27(5):2–7, September/ October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2014:FSE

[ES14] Christof Ebert and Hassan Soubra. Functional size estimation technologies for software maintenance. IEEE Software, 31(6):24–29, November/ December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060024.html.

Elbanna:2016:RAS

[ES16] Amany Elbanna and Suprateek Sarker. The risks of agile software development: Learning from adopters. IEEE Software, 33(5):72–79, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050072-abs.
html.

Ebert:2017:IT

[ES17] Christof Ebert and Kris Shankar. Industry trends 2017. IEEE Software, 34(2):112–116, March/ April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL https://

www.computer.org/csdl/
mags/so/2017/02/mso2017020112-
abs.html.

Eugster:2015:DIT

[ESZ15] Patrick Eugster, Vinaitheerthan Sundaram, and Xiangyu Zhang. Debugging the Internet of Things: The case of wireless sensor networks. IEEE Software, 32(1):38–49, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/

01/mso2015010038-abs.
html.

Eveleens:2010:RFC

[EV10] J. Laurenz Eveleens and Chris Verhoef. The rise and fall of the chaos report figures. IEEE Software, 27(1):30–36, January/February 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ebert:2019:VAS

[EW19] C. Ebert and M. Weyrich.

Validation of autonomous systems. IEEE Software, 36(5):15–23, September/ October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Faber:2010:ASP

[Fab10]  Roland Faber.  Archi-

tects as service providers. IEEE Software, 27(2):33– 40, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fairbanks:2014:AH

[Fai14] George Fairbanks. Architectural hoisting. IEEE Software, 31(4):12–15, July/August 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2019:BCR

[Fai19a]  G. Fairbanks.  Better

code reviews with design by contract. IEEE Software, 36(6):53–56, November/December 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2019:HCR

[Fai19b] G. Fairbanks. Healthy code reveals the problem and solution. IEEE Software, 36(5):76–79, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2019:IRR

[Fai19c]  G. Fairbanks.  Ignore,

refactor, or rewrite. IEEE Software, 36(2):133–136, March/April 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2019:ICP

[Fai19d] G. Fairbanks. Intellectual control [pragmatic designer]. IEEE Software, 36(1):91–94, January/February 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2019:PLE

[Fai19e]  G. Fairbanks.  Princi-

ple of least expressiveness.

IEEE Software, 36(3):116– 119, May/June 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2019:SYT

[Fai19f]  G. Fairbanks.  Scale

your team horizontally. IEEE Software, 36(4):88– 90, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Farwell:2017:YFA

[Far17] Matthew Farwell. Yakov Fain on Angular. IEEE Software, 34(6):109–112, November/December 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060109.html.

Favaro:2010:GEI

[Fav10] John Favaro. Guest Editor’s introduction: Renewing the software project management life cycle. IEEE Software, 27(1): 17–19, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Favaro:2012:ESI

[Fav12]  John Favaro.  Excel-

lence in search: An interview with David Chaiken. IEEE Software, 29(1):

84–86, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Felderer:2015:UDT

[FB15] Michael Felderer and Armin Beer. Using defect taxonomies for testing requirements. IEEE Software, 32(3):94–101, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030094-abs.
html.

Falessi:2019:ESU

[FC19] D. Falessi and G. Cantone. The effort savings from using NLP to classify equivalent requirements. IEEE Software, 36(1): 48–55, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Falcarin:2011:GEI

[FCAJ11] Paolo Falcarin, Christian Collberg, Mikhail Atallah, and Mariusz Jakubowski. Guest Editors’ introduction: Software protection. IEEE Software, 28(2):24– 27, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Frey:2016:HSA

[FCN16] Soren Frey, Lambros Charissis, and Jens Nahm. How software architects drive connected vehicles. IEEE Software, 33(6):41–47, November/ December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060041-abs.
html.

Falessi:2010:PCA

[FCS^{+}10] Davide Falessi, Giovanni Cantone, Salvatore Alessandro Sarcia’, Giuseppe Calavaro, Paolo Subiaco, and Cristiana D’Amore. Peaceful coexistence: Agile developer perspectives on software architecture. IEEE Software, 27(2):23– 25, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ferrari:2017:NLR

[FDE^{+}17] Alessio Ferrari, Felice DellOrletta, Andrea Esuli, Vincenzo Gervasi, and Stefania Gnesi. Natural language requirements processing: A 4D vision. IEEE Software, 34(6):28–35, November/ December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        06/mso2017060028-abs.
                                        html.

Fernandez:2018:SRE

[Fer18] Daniel Mendez Fernandez. Supporting requirementsengineering research that industry needs: The NaPiRE initiative. IEEE Software, 35(1):112–116, January/February 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2018/
                                        01/mso2018010112-abs.
                                        html.

Ferrari:2013:MBD

[FFGM13] Alessio Ferrari, Alessandro Fantechi, Stefania Gnesi, and Gianluca Magnani. Model-based development and formal methods in the railway industry. IEEE Software, 30 (3):28–34, May/June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Furda:2018:MEL

[FFZ^{+}18] A. Furda, C. Fidge, O. Zimmermann, W. Kelly, and A. Barros. Migrating enterprise legacy source code to microservices: On multitenancy, statefulness, and data consistency. IEEE Software, 35(3):63–72, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fagerholm:2014:OOS

[FGBM14] Fabian Fagerholm, Alejandro Sanchez Guinea, Jay Borenstein, and Jurgen Munch. Onboarding in open source projects. IEEE Software, 31(6):54–61, November/ December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060054-abs.
html.

Fricker:2010:HIP

[FGBS10] Samuel Fricker, Tony Gorschek, Carl Byman, and Armin Schmidle. Handshaking with implementation proposals: Negotiating requirements understanding. IEEE Software, 27(2):72–80, March/ April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fonseca:2019:MEA

[FKL19] A. Fonseca, R. Kazman, and P. Lago. A manifesto for energy-aware software. IEEE Software, 36(6):79– 82, November/December 2019. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Falcini:2017:DLA

[FLC17] Fabio Falcini, Giuseppe Lami, and Alessandra Mitidieri Costanza. Deep learning in automotive software. IEEE Software, 34(3):56–63, May/ June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL https://

www.computer.org/csdl/
mags/so/2017/03/mso2017030056.
html.

Fysarakis:2016:NDE

[FMMP16] Konstantinos Fysarakis, Damianos Mylonakis, Charalampos Manifavas, and Ioannis Papaefstathiou. Node.DPWS: Efficient Web services for the Internet of Things. IEEE Software, 33(3):60–67, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030060-abs.
html.

Fenton:2014:DSS

[FN14] Norman Fenton and Martin Neil. Decision support software for probabilistic risk assessment using Bayesian networks. IEEE Software, 31(2):21– 26, March/April 2014.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Favaro:2011:GEI

[FP11] John Favaro and Shari Lawrence Pfleeger. Guest Editors’ introduction: Software as a business. IEEE Software, 28(4):22–25, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fox:2013:NSE

[FP13] Armando Fox and David Patterson. Is the new software engineering curriculum agile? IEEE Software, 30(5):88, September/October 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Freudenthal:2010:DSL

[Fre10] Margus Freudenthal. Domainspecific languages in a customs information system. IEEE Software, 27(2):65– 71, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fontana:2015:ACT

[FRM15] Rafaela Mantovani Fontana, Sheila Reinehr, and Andreia Malucelli. Agile compass: A tool for identifying maturity in agile software-development teams. IEEE Software, 32(6):20–23, November/ December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060020-abs.
html.

Freudenberg:2010:TBR

[FS10] Sallyann Freudenberg and Helen Sharp. The top 10 burning research questions from practitioners. IEEE Software, 27(5):8–9, September/October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ferragina:2012:FAA

[FS12] Paolo Ferragina and Ugo Scaiella. Fast and accurate annotation of short texts with Wikipedia pages. IEEE Software, 29(1): 70–75, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Falessi:2012:PSS

[FSB^{+}12] Davide Falessi, Mehrdad Sabetzadeh, Lionel Briand, Emanuele Turella, Thierry Coq, and Rajwinder Kaur Panesar-Walawege. Planning for safety standards compliance: a modelbased tool-supported approach. IEEE Software, 29 (3):64–70, May/June 2012.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Falessi:2014:AMC

[FSM14] Davide Falessi, Michele Shaw, and Kathleen Mullen. Achieving and maintaining CMMI Maturity Level 5 in a small organization. IEEE Software, 31(5):80–86, September/ October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050080-abs.
html.

Furtado:2011:IDG

[FSRdA11] Andre W. B. Furtado, Andre L. M. Santos, Geber L. Ramalho, and Eduardo Santana de Almeida. Improving digital game development with software product lines. IEEE Software, 28(5):30–37, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Feinbube:2011:JFM

[FTP11] Frank Feinbube, Peter Troger, and Andreas Polze. Joint forces: From multithreaded programming to GPU computing. IEEE Software, 28(1): 51–57, January/February 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Femmer:2019:RQQ

[FV19] H. Femmer and A. Vogelsang. Requirements quality is quality in use. IEEE Software, 36(3):83– 91, May/June 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fernandez:2015:NPR

[FWK^{+}15] Daniel Mendez Fernandez, Stefan Wagner, Marcos Kalinowski, Andre Schekelmann, Ahmet Tuzcu, Tayana Conte, Rodrigo Spinola, and Rafael Prikladnicki. Naming the pain in requirements engineering: Comparing practices in Brazil and Germany. IEEE Software, 32(5):16–23, September/ October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        05/mso2015050016-abs.
                                        html.

Farias:2019:WCI

[FWS19] V. Farias, I. Wiese, and R. Santos. What characterizes an influencer in software ecosystems? IEEE Software, 36(1): 42–47, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gomez:2017:ASC

[GAM^{+}17] Maria Gomez, Bram Adams, Walid Maalej, Martin Monperrus, and Romain Rouvoy. App Store 2.0: From crowdsourced information to actionable feedback in mobile ecosystems. IEEE Software, 34(2):81–89, March/April 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020081-abs.
html.

Gatrell:2016:VSS

[Gat16] Matt Gatrell. The value of a single solution for end-to-end ALM tool support. IEEE Software, 33(5):103–105, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050103-abs.
html.

Gaumer:2013:PRA

[Gau13] Dale Gaumer. Peer review anxiety. IEEE Software, 30(6):11–13, November/December 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gonzalez-Barahona:2013:UHC

[GBICMR13] Jesus M. Gonzalez-Barahona, Daniel Izquierdo-Cortazar, Stefano Maffulli, and Gregorio Robles. Understanding how companies interact with free software communities. IEEE Software, 30(5):38–45, September/ October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gorton:2016:SEB

[GBM16] Ian Gorton, Ayse Basar Bener, and Audris Mockus. Software engineering for big data systems. IEEE Software, 33(2):32–35, March/April 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        02/mso2016020032.html.

Gotel:2013:REN

[GCH13] Olly Gotel and Jane Cleland-Huang. Requirements engineering’s next top model. IEEE Software, 30(6):24–29, November/ December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gangadharan:2011:MCM

[GD11] G. R. Gangadharan and Vincenzo D’Andrea. Managing copyrights and moral rights of service-based software. IEEE Software, 28(2):48–55, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

German:2012:MOS

[GD12] Daniel German and Massimiliano Di Penta. A method for open source license compliance of Java applications. IEEE Software, 29(3):58–63, May/ June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gavalas:2011:DPM

[GE11] Damianos Gavalas and Daphne Economou. Development platforms for mobile applications: Status and trends. IEEE Software, 28(1):77–86, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gat:2012:PC

[GE12] Israel Gat and Christof Ebert. Point counterpoint. IEEE Software, 29(6):52– 55, November/December 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Garousi:2017:TAJ

[GE17] Vahid Garousi and Frank Elberzhager. Test automation: Not just for test execution. IEEE Software, 34(2):90–96, March/April 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020090-abs.
html.

Garousi:2016:DVM

[GF16] Vahid Garousi and Michael Felderer. Developing, verifying, and maintaining high-quality automated test scripts. IEEE Software, 33(3):68–75, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030068-abs.
html.

Garousi:2017:WAI

[GF17] Vahid Garousi and Michael Felderer. Worlds apart: Industrial and academic focus areas in software testing. IEEE Software, 34(5):38–45, September/ October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.

org/csdl/mags/so/2017/
05/mso2017050038-abs.
html.

Garousi:2018:WWKa

[GFH18] Vahid Garousi, Michael Felderer, and Tuna Hacaloglu. What we know about software test maturity and test process improvement. IEEE Software, 35(1):84–92, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010084-abs.
html.

Garousi:2018:WWKb

[GFKY18] V. Garousi, M. Felderer, Çağri M. Karapıçak, and Uğur Yılmaz. What we know about testing embedded software. IEEE Software, 35(4):62– 69, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gorschek:2010:LIP

[GFPK10] Tony Gorschek, Samuel Fricker, Kenneth Palm, and Steven Kunsman. A lightweight innovation process for softwareintensive product development. IEEE Software, 27(1):37–45, January/February 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gamez:2015:CSA

[GFT15] Nadia Gamez, Lidia Fuentes, and Jose M. Troya. Creating self-adapting mobile systems with dynamic software product lines. IEEE Software, 32(2):105– 112, March/April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        02/mso2015020105-abs.
                                        html.

Gutierrez:2019:SME

[GGGM19] G. Gutierrez, J. Garzas, M. T. Gonzalez de Lena, and J. M. Moguerza. Selfmanaging: An empirical study of the practice in agile teams. IEEE Software, 36(1):23–27, January/February 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ghaisas:2014:PWW

[Gha14]  Smita Ghaisas.  Prac-

ticing what we preach. IEEE Software, 31(1): 88–92, January/February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Graydon:2012:AC

[GHH^{+}12] Patrick Graydon, Ibrahim Habli, Richard Hawkins, Tim Kelly, and John Knight. Arguing conformance. IEEE Software, 29 (3):50–57, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ghosh:2010:MDS

[Gho10] Debasish Ghosh. Multiparadigm data storage for enterprise applications. IEEE Software, 27(5):57– 60, September/October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Groenewegen:2010:SCL

[GHV10] Danny M. Groenewegen, Zef Hemel, and Eelco Visser. Separation of concerns and linguistic integration in WebDSL. IEEE Software, 27(5):31– 37, September/October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gimpel:2014:SCS

[Gim14]  James Gimpel.  Soft-

ware that checks software: The impact of PClint. IEEE Software, 31(1): 15–19, January/February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Goues:2018:BGR

[GJO^{+}18] C. L. Goues, C. Jaspan, I. Ozkaya, M. Shaw, and K. T. Stolee. Bridging the gap: From research to practical advice. IEEE Software, 35(5):50–57, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gorton:2015:DDD

[GK15a] Ian Gorton and John Klein. Distribution, data, deployment: Software architecture convergence in big data systems. IEEE Software, 32 (3):78–85, May/June 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030078-abs.
html.

Grandison:2015:SPW

[GK15b] Tyrone Grandison and Larry Koved. Security and privacy on the Web [Guest Editors’ introduction]. IEEE Software, 32(4):36–39, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040036.html.

Goldsteen:2015:ASM

[GKD^{+}15] Abigail Goldsteen, Ksenya Kveler, Tamar Domany, Igor Gokhman, Boris Rozenberg, and Ariel Farkash. Applicationscreen masking: A hybrid approach. IEEE Software, 32(4):40–45, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040040-abs.
html.

Garousi:2019:WWK

[GKF19] V. Garousi, B. Kucuk, and M. Felderer. What we know about smells in software test code. IEEE Software, 36(3):61–73, May/ June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Grundy:2012:GEI

[GKKL12] John Grundy, Gerald Kaefer, Jacky Keong, and Anna Liu. Guest Editors’ introduction: Software engineering for the Cloud. IEEE Software, 29(2):26–29, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Glass:2012:GGI

[Gla12] Robert L. Glass. The gender gap: Is it a computing problem or simply a computer science problem? IEEE Software, 29(2):88, 87, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). See response [Ris12b].

Galster:2014:SVI

[GLA14] Matthias Galster, Laurens Lapre, and Paris Avgeriou. SOA in variabilityintensive environments: Pitfalls and best practices. IEEE Software, 31(1): 77–84, January/February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gotterbarn:2010:UYS

[GM10] Donald Gotterbarn and Keith W. Miller. Unmasking your software’s ethical risks. IEEE Software, 27(1):12–13, January/February 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gotel:2011:RT

[GM11] Olly Gotel and Stephen Morris. Requirements tracery. IEEE Software, 28(5):92–94, September/ October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Guillen:2014:PSM

[GMB^{+}14] Joaquin Guillen, Javier Miranda, Javier Berrocal, Jose Garcia-Alonso, Juan Manuel Murillo, and Carlos Canal. People as a service: A mobilecentric model for providing collective sociological profiles. IEEE Software, 31(2):48–53, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ghazi:2014:LHG

[GMP14] Parisa Ghazi, Ana M.

Moreno, and Lawrence Peters. Looking for the Holy Grail of software development. IEEE Software, 31(1):96, January/ February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Griswold:2015:BRR

[GO15] William G. Griswold and William F. Opdyke. The birth of refactoring: A retrospective on the nature of high-impact software engineering research. IEEE Software, 32(6):30– 38, November/December 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060030-abs.
html.

Gorton:2013:CDD

[Gor13] Ian Gorton. Cyber dumpster diving: Creating new software systems for less. IEEE Software, 30(1): 9–13, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Giordano:2013:VCM

[GP13] Massimiliano Giordano and Giuseppe Polese. Visual computer-managed security: A framework for developing access control in enterprise applications. IEEE Software, 30(5):62–69, September/ October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Garrido:2011:RUW

[GRD11] Alejandra Garrido, Gustavo Rossi, and Damiano Distante. Refactoring for usability in Web applications. IEEE Software, 28 (3):60–67, May/June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Garcia-Rosello:2012:TLW

[GRDL^{+}12]  Emilio  Garcia-Rosello,

Jacinto G. Dacosta, Maria J. Lado, Arturo J. Mendez, and Jose Garcia Perez-Schofield. Two-layer wrapping for COTS software integration: An experience with Matlab. IEEE Software, 29(4):76–82, July/ August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gregory:2017:RMC

[Gre17a] Sarah Gregory. RE@40:

Midlife crisis or graceful maturity? IEEE Software, 34(6):14–17, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060014.html.

Gregory:2017:UJR

[Gre17b] Sarah Gregory. The unplanned journey of a requirements engineer in industry: An introduction. IEEE Software, 34(5):16–19, September/ October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050016-abs.
html.

Gregory:2018:HCC

[Gre18a] S. Gregory. How common is common enough in requirements-engineering practice? IEEE Software, 35(3):20–23, May/June 2018. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Gregory:2018:IDH

[Gre18b] S. Gregory. It depends: Heuristics for commonenough requirements practice. IEEE Software, 35(4): 12–15, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gregory:2019:REQ

[Gre19] S. Gregory. Requirements engineering: The quest for meaningful metrics: Time for a change? IEEE Software, 36(6):7–11, November/December 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Giacaman:2011:OOP

[GS11] Nasser Giacaman and Oliver Sinnen. Objectoriented parallelisation of Java desktop programs. IEEE Software, 28(1): 32–38, January/February 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Groen:2017:CRE

[GSA^{+}17] Eduard C. Groen, Norbert Seyff, Raian Ali, Fabiano Dalpiaz, Joerg Doerr, Emitza Guzman, Mahmood Hosseini, Jordi Marco, Marc Oriol, Anna Perini, and Melanie Stade.

The crowd in requirements engineering: The landscape and challenges. IEEE Software, 34(2):44– 52, March/April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020044-abs.
html.

Gelbard:2010:OOA

[GTS10] Roy Gelbard, Dov Teeeni, and Matti Sade. Objectoriented analysis: Is it just theory? IEEE Software, 27(1):64–71, January/February 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Garousi:2019:ATS

[GTS^{+}19] V. Garousi, S. Tasli, O. Sertel, M. Tokgoz, K. Herkiloglu, H. F. E. Arkin, and O. Bilir. Automated testing of simulation software in the aviation industry: An experience report. IEEE Software, 36(4):63–75, July/ August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Guerra:2014:DFT

[Gue14]  Eduardo Guerra.  De-

signing a framework with test-driven development:

A journey. IEEE Software, 31(1):9–14, January/February 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gunther:2010:MDA

[Gun10] Sebastian Gunther. Multi-DSL applications with Ruby. IEEE Software, 27(5):25–30, September/ October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Giardino:2014:WDW

[GUP^{+}14] Carmine Giardino, Michael Unterkalmsteiner, Nicolo Paternoster, Tony Gorschek, and Pekka Abrahamsson. What do we know about software development in startups? IEEE Software, 31(5):28–32, September/ October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        05/mso2014050028-abs.
                                        html.

Glass:2011:NSS

[GV11] Robert L. Glass and Iris Vessey. Naivete squared: In search of two taxonomies and a mapping between them. IEEE Software, 28(5):14– 15, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Garcia:2011:PMT

[GVE11] Felix Garcia, Aurora Vizcaino, and Christof Ebert. Process management tools. IEEE Software, 28(2):15–18, March/ April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Garcia:2017:CBA

[GVSP^{+}17] Ana Cristina Bicharra Garcia, Adriana Santarosa Vivacqua, Nayat Sanchez-Pi, Luis Marti, and Jose M. Molina. Crowd-based ambient assisted living to monitor the elderly’s health outdoors. IEEE Software, 34(6):53–57, November/ December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060053-abs.
html.

Graziotin:2014:SDM

[GWA14] Daniel Graziotin, Xiaofeng Wang, and Pekka Abrahamsson. Software developers, moods, emotions, and performance. IEEE Software, 31(4):24– 27, July/August 2014.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gorton:2011:CP

[GWLY11] Ian Gorton, Adam Wynne, Yan Liu, and Jian Yin. Components in the pipeline. IEEE Software, 28(3):34– 40, May/June 2011. CO-DEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gu:2011:PC

[GWP^{+}11] Yuan Xiang Gu, Brecht Wyseur, Bart Preneel, Jean-Daniel Aussel, and Reiner Sailer. Point/ counterpoint. IEEE Software, 28(2):56–59, March/ April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Harrison:2011:PBA

[HA11] Neil B. Harrison and Paris Avgeriou. Patternbased architecture reviews. IEEE Software, 28(6):66–71, November/ December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hedin:2011:ELL

[HAE11] Gorel Hedin, Johan Akesson, and Torbjorn Ekman. Extending languages by leveraging compilers: From Modelica to Optimica.

IEEE Software, 28(3):68– 74, May/June 2011. CO-DEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hamilton:2018:WET

[Ham18] M. H. Hamilton. What the errors tell us. IEEE Software, 35(5):32–37, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hazzan:2010:PHA

[Haz10] Orit Hazzan. Putting human aspects of software engineering in university curricula. IEEE Software, 27(4):90–91, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hall:2011:DFP

[HBB^{+}11] Tracy Hall, Sarah Beecham, David Bowes, David Gray, and Steve Counsell. Developing fault-prediction models: What the research can show industry. IEEE Software, 28(6):96– 99, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hirzalla:2014:BAT

[HBCH14] Mamoun Hirzalla, Peter Bahrs, and Jane Cleland-Huang. Beyond anecdotal thinking: Deepening our understanding for achieving quality goals. IEEE Software, 31(3):26– 29, May/June 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hoda:2013:TLT

[HBN13] Rashina Hoda, Jeffry Babb, and Jacob Norbjerg. Toward learning teams. IEEE Software, 30(4):95–98, July/August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hannay:2017:BPB

[HBS17a] Jo Erskine Hannay, Hans Christian Benestad, and Kjetil Strand. Benefit points: The best part of the story. IEEE Software, 34(3):73– 85, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030073-abs.
html.

Hannay:2017:EBV

[HBS17b] Jo Erskine Hannay, Hans Christian Benestad, and Kjetil Strand. Earned business value: See that you deliver value to your customer. IEEE Software, 34(4):58– 70, July/August 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040058-abs.
html.

Hannay:2019:AUA

[HBS19] J. E. Hannay, H. C. Benestad, and K. Strand. Agile uncertainty assessment for benefit points and story points. IEEE Software, 36(4):50–62, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Haberle:2015:CCC

[HCF^{+}15] Tobias Haberle, Lambros Charissis, Christoph Fehling, Jens Nahm, and Frank Leymann. The connected car in the cloud: A platform for prototyping telematics services. IEEE Software, 32(6):11–17, November/ December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060011-abs.
html.

Henry:2019:HCL

[Hen19] G. Henry. Howard Chu on lightning memory-mapped database. IEEE Software, 36(6):83–87, November/December 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). See minor corrections [?].

Hermans:2017:PHN

[Her17] Felienne Hermans. Peter Hilton on naming. IEEE Software, 34(3):117–120, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        03/mso2017030117.html.

Hernantes:2015:IIM

[HGS15] Josune Hernantes, Gorka Gallardo, and Nicolas Serrano. IT infrastructuremonitoring tools. IEEE Software, 32(4):88–93, July/August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        04/mso2015040088.html.

Huygens:2010:SDN

[HHLJ10] Christophe Huygens, Danny Hughes, Bert Lagaisse, and Wouter Joosen. Streamlining development for networked embedded systems using multiple paradigms. IEEE Software, 27(5):45– 52, September/October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hassan:2013:RWN

[HHR^{+}13] Ahmed E. Hassan, Abram Hindle, Per Runeson, Martin Shepperd, Prem Devanbu, and Sunghun Kim. Roundtable: What’s next in software analytics. IEEE Software, 30(4): 53–56, July/August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hindle:2019:CLM

[Hin19] A. Hindle. Complexity:

Let’s not make this complicated. IEEE Software, 36(2):130–132, March/ April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hatton:2017:DGE

[HJ17] Leo Hatton and Alan John. Delivering genuine emails in an ocean of spam. IEEE Software, 34(4):11– 15, July/August 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040011-abs.
html.

Hill:2016:AAF

[HJP16] Emily Hill, Philip M.

Johnson, and Daniel Port. Is an athletic approach the future of software engineering education? IEEE Software, 33(1):97–100, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010097-abs.
html.

Herbsleb:2016:ITS

[HKB16] James Herbsleb, Christian Kastner, and Christopher Bogart. Intelligently transparent software ecosystems. IEEE Software, 33(1):89–96, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010089-abs.
html.

Happe:2011:FPP

[HKR11] Jens Happe, Heiko Koziolek, and Ralf Reussner. Facilitating performance predictions using software components. IEEE Software, 28(3):27–33, May/ June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hneif:2011:UGI

[HL11] Malik Hneif and Sai Peck Lee. Using guidelines to improve quality in software nonfunctional attributes. IEEE Software, 28(6):72–77, November/ December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hsu:2017:AVG

[HLS17] Chia-Wei Hsu, Shan-Hsin Lee, and Shiuhpyng Winston Shieh. Adaptive virtual gestures for GUI testing on Smartphones. IEEE Software, 34(5):22–29, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050022-abs.
html.

Hollis:2013:EAP

[HM13] Bianca Hollis and Neil Maiden. Extending agile processes with creativity techniques. IEEE Software, 30(5):78–84, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hackbarth:2016:ISQ

[HMPS16] Randy Hackbarth, Audris Mockus, John Palframan, and Ravi Sethi. Improving software quality as customers perceive it. IEEE Software, 33(4): 40–45, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        04/mso2016040040-abs.
                                        html.

Hafiz:2015:RM

[HO15] Munawar Hafiz and Jeffrey Overbey. Refactoring myths. IEEE Software, 32(6):39–43, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        06/mso2015060039-abs.
                                        html.

Haghighatkhah:2017:ISA

[HOBK17] Alireza Haghighatkhah, Markku Oivo, Ahmad Banijamali, and Pasi Kuvaja. Improving the state of automotive software engineering. IEEE Software, 34(5):82–86, September/ October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        05/mso2017050082.html.

Hoda:2013:PP

[Hod13] Rashina Hoda. Power to the people. IEEE Software, 30(2):92, March/ April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hofstedt:2010:CBO

[Hof10] Petra Hofstedt. Constraintbased object-oriented programming. IEEE Software, 27(5):53–56, September/October 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Holzmann:2013:LSM

[Hol13] Gerard J. Holzmann.

Landing a spacecraft on Mars. IEEE Software, 30(2):83–86, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2014:FIR

[Hol14] Gerard J. Holzmann.

Fault intolerance [Reliable code]. IEEE Software, 31(6):16–20, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060016.html.

Holzmann:2015:ATR

[Hol15a] Gerard J. Holzmann. Assertive testing [reliable code]. IEEE Software, 32 (3):12–15, May/June 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/

                                        03/mso2015030012-abs.
                                        html.

Holzmann:2015:CE

[Hol15b] Gerard J. Holzmann. Code evasion. IEEE Software, 32(5):77–80, September/ October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        05/mso2015050077.html.

Holzmann:2015:CI

[Hol15c] Gerard J. Holzmann. Code inflation. IEEE Software, 32(2):10–13, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        02/mso2015020010.html.

Holzmann:2015:B

[Hol15d] Gerard J. Holzmann. Out of bounds. IEEE Software, 32(6):24–26, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        06/mso2015060024.html.

Holzmann:2015:PT

[Hol15e] Gerard J. Holzmann. Points of truth. IEEE Software, 32(4):18–21, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040018.html.

Holzmann:2015:CH

[Hol15f] Gerard J. Holzmann. To code is human. IEEE Software, 32(1):14–17, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010014.html.

Holzmann:2016:BY

[Hol16a] Gerard J. Holzmann.

Brace yourself. IEEE Software, 33(5):34–37, September/October 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050034-abs.
html.

Holzmann:2016:CC

[Hol16b] Gerard J. Holzmann.

Code clarity. IEEE Software, 33(2):22–25, March/April 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/

                                        02/mso2016020022-abs.
                                        html.

Holzmann:2016:FUQ

[Hol16c] Gerard J. Holzmann. Frequently unanswered questions. IEEE Software, 33 (3):10–12, May/June 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        03/mso2016030010.html.

Holzmann:2016:HM

[Hol16d] Gerard J. Holzmann. Hi maintenance. IEEE Software, 33(6):98–101, November/December 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        06/mso2016060098-abs.
                                        html.

Holzmann:2016:TT

[Hol16e] Gerard J. Holzmann. Tiny tools. IEEE Software, 33(1):24–28, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2016/
                                        01/mso2016010024.html.

Holzmann:2016:WL

[Hol16f] Gerard J. Holzmann. The weakest link. IEEE Software, 33(4):18–21, July/ August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040018-abs.
html.

Holzmann:2017:CC

[Hol17a] Gerard J. Holzmann.

Code craft. IEEE Software, 34(2):18–21, March/ April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL https://

www.computer.org/csdl/
mags/so/2017/02/mso2017020018-
abs.html.

Holzmann:2017:DP

[Hol17b] Gerard J. Holzmann. Dead programs. IEEE Software, 34(4):89–91, July/ August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040089-abs.
html.

Holzmann:2017:RR

[Hol17c] Gerard J. Holzmann. Randomly right. IEEE Software, 34(5):87–89, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/

                                        05/mso2017050087-abs.
                                        html.

Holzmann:2017:TTP

[Hol17d] Gerard J. Holzmann. A tale of three programs. IEEE Software, 34(3):23– 26, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        03/mso2017030023-abs.
                                        html.

Holzmann:2017:VD

[Hol17e] Gerard J. Holzmann. The value of doubt. IEEE Software, 34(1):106–109, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        01/mso2017010106-abs.
                                        html.

Holzmann:2018:CV

[Hol18a] G. J. Holzmann. Code vault. IEEE Software, 35(5):85–87, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2018:CB

[Hol18b] G. J. Holzmann. Curve balls. IEEE Software, 35(2):18–21, March/April 2018. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2018:SC

[Hol18c] G. J. Holzmann. Software components. IEEE Software, 35(3):80–82, May/ June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2018:EM

[Hol18d] Gerard J. Holzmann. Ends and means. IEEE Software, 35(1):14–17, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010014-abs.
html.

Hole:2019:DSS

[Hol19a]  K. J. Hole.  Domi-

nating software systems: How to overcome online information asymmetry. IEEE Software, 36(4):81– 87, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2019:CM

[Hol19b] G. J. Holzmann. Code mining. IEEE Software, 36(2):25–29, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2019:CO

[Hol19c] G. J. Holzmann. Code overload. IEEE Software, 36(6):73–75, November/December 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2019:DC

[Hol19d] G. J. Holzmann. Does not compute. IEEE Software, 36(3):14–16, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hohpe:2016:SAR

[HOZZ16] Gregor Hohpe, Ipek Ozkaya, Uwe Zdun, and Olaf Zimmermann. The software architect’s role in the digital age. IEEE Software, 33(6):30–39, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        06/mso2016060030.html.

Hatzivasilis:2016:SSP

[HPM16] George Hatzivasilis, Ioannis Papaefstathiou, and Charalampos Manifavas. Software security, privacy, and dependability: Metrics and measurement. IEEE Software, 33(4):46– 54, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040046-abs.
html.

Held:2011:PCO

[HR11] Isaac Held and David Randall. Point/counterpoint: Open source climate model development is worth it / should climate models be open source? IEEE Software, 28(6):62–65, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hathhorn:2019:DCO

[HR19a] C. Hathhorn and G. Rosu.

Dealing with C’s original sin. IEEE Software, 36(5):24–28, September/ October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hatton:2019:LML

[HR19b] L. Hatton and A. Rutkowski.

“Lessons must be learned”

— but are they? IEEE Software, 36(4):91–95, July/August 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hill:2010:TCE

[HSEG10] James Hill, D. Schmidt, James Edmondson, and Aniruddha Gokhale. Tools for continuously evaluating distributed system qualities. IEEE Software, 27(4):65–71, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hermans:2016:LTF

[HSF^{+}16] Felienne Hermans, Janet Siegmund, Thomas Fritz, Gabriele Bavota, Meiyappan Nagappan, Abram Hindle, Yasutaka Kamei, Ali Mesbah, and Bram Adams. Leaders of tomorrow on the future of software engineering: A roundtable. IEEE Software, 33(2):99–104, March/April 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
02/mso2016020099.html.

Hoda:2018:REA

[HSG18] R. Hoda, N. Salleh, and J. Grundy. The rise and evolution of agile software development. IEEE Software, 35(5):58– 63, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Heidrich:2016:EBD

[HTE16] Jens Heidrich, Adam Trendowicz, and Christof Ebert. Exploiting Big Data’s benefits. IEEE Software, 33(4):111–116, July/August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040111-abs.
html.

Hurlburt:2016:SDS

[HV16] George Hurlburt and Jeffrey Voas. Software is driving software engineering? IEEE Software, 33(1):101–104, January/ February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010101.html.

Hatton:2019:CRE

[Hv19] L. Hatton and M. van Genuchten. Computational reproducibility: The elephant in the room. IEEE Software, 36(2):137– 144, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hofland:2010:SMS

[HvdL10] Lennart Hofland and Joop van der Linden. Software in MRI scanners. IEEE Software, 27(4):87– 89, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Hatton:2012:EDD

[HvG12] Les Hatton and Michiel van Genuchten. Early design decisions. IEEE Software, 29(1):87–89, January/February 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hatton:2016:WSC

[HvG16] Les Hatton and Michiel van Genuchten. When software crosses a line. IEEE Software, 33(1): 29–31, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/so/2016/
                                        01/mso2016010029.html.

Hohpe:2013:TYP

[HWBYZ13] Gregor Hohpe, Rebecca Wirfs-Brock, Joseph W. Yoder, and Olaf Zimmermann. Twenty years of patterns’ impact. IEEE Software, 30(6):88, November/December 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Izquierdo:2019:SDA

[IGBKR19] D. Izquierdo, J. M. Gonzalez-Barahona, L. Kurth, and G. Robles. Software development analytics for Xen: Why and how. IEEE Software, 36 (3):28–32, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Izquierdo:2019:OGD

[IHSR19] D. Izquierdo, N. Huesman, A. Serebrenik, and G. Robles. OpenStack gender diversity report. IEEE Software, 36(1): 28–33, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Introne:2011:EOD

[ILM11] Joshua Introne, Robert Laubacher, and Thomas Malone. Enabling open development methodologies in climate change assessment modeling. IEEE Software, 28(6):56–61, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Itkonen:2016:TBE

[IML16] Juha Itkonen, Mika V.

Mantyla, and Casper Lassenius. Test better by exploring: Harnessing human skills and knowledge. IEEE Software, 33(4):90– 96, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/

                                        04/mso2016040090-abs.
                                        html.

Jansen:2019:MSP

[JCP19] S. Jansen, M. Cusumano, and K. M. Popp. Managing software platforms and ecosystems. IEEE Software, 36(3):17–21, May/ June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Johnson:2012:WTS

[JEJ12] Pontus Johnson, Mathias Ekstedt, and Ivar Jacobson. Where’s the theory for software engineering? IEEE Software, 29(5):96, September/October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jeppu:2013:FCS

[Jep13] Yogananda Jeppu. Flight control software: Mistakes made and lessons learned. IEEE Software, 30(3):67– 72, May/June 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jones:2010:CWS

[JN10] Lawrence G. Jones and Linda M. Northrop. Clearing the way for software product line success. IEEE Software, 27 (3):22–28, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

John:2010:UDP

[Joh10] Isabel John. Using documentation for product line scoping. IEEE Software, 27(3):42–47, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Johnson:2013:SUS

[Joh13] Philip M. Johnson. Searching under the streetlight for useful software analytics. IEEE Software, 30(4): 57–63, July/August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Johann:2015:SAD

[Joh15] Sven Johann. Software architecture for developers. IEEE Software, 32(5):93–96, September/ October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050093.html.

Johann:2016:DTI

[Joh16]  Sven Johann.  Dave

Thomas on innovating legacy systems. IEEE Software, 33(2):105–108, March/April 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
02/mso2016020105.html.

Johann:2017:KMI

[Joh17] Sven Johann. Kief Morris on infrastructure as code. IEEE Software, 34(1):117–120, January/ February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010117-abs.
html.

Jorgensen:2013:RES

[Jor13] Magne Jorgensen. Relative estimation of software development effort: It matters with what and how you compare. IEEE Software, 30(2):74–79, March/ April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jorgensen:2014:WWD

[Jor14] Magne Jorgensen. What we do and don’t know about software development effort estimation. IEEE Software, 31(2):37– 40, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jorgensen:2016:BSS

[Jor16] Magne Jorgensen. Better selection of software providers through trialsourcing. IEEE Software, 33(5):48–53, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050048-abs.
html.

Jorgensen:2019:RBP

[Jor19] M. Jorgensen. Relationships between project size, agile practices, and successful software development: Results and analysis. IEEE Software, 36(2): 39–43, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jamshidi:2018:MJF

[JPM^{+}18] P. Jamshidi, C. Pahl, N. C. Mendonça, J. Lewis, and S. Tilkov. Microservices: The journey so far and challenges ahead. IEEE Software, 35(3):24– 35, May/June 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jung:2018:JDF

[Jun18] J. Jung. John Doran on fixing a broken development process. IEEE Software, 35(6):77–80, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jolak:2018:DDS

[JWCR18] R. Jolak, A. Wortmann, M. Chaudron, and B. Rumpe. Does distance still matter? Revisiting collaborative distributed software design. IEEE Software, 35(6):40–47, November/ December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kaatz:2014:HSI

[Kaa14]  Tobias Kaatz.  Hir-

ing in the software industry. IEEE Software, 31(6):96, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060096.html.

Khomh:2018:SEM

[KAC^{+}18] F. Khomh, B. Adams, J. Cheng, M. Fokaefs, and G. Antoniol. Software engineering for machinelearning applications: The road ahead. IEEE Software, 35(5):81–84, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

King:2012:CKF

[KB12] Clive King and Chris Beal.

CSI kernel: Finding a needle in a multiterabyte haystack. IEEE Software, 29(6):9–12, November/December 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kwon:2016:PSC

[KC16] Shinil Kwon and Sungdeok Cha. A paradigm shift for the CAPTCHA race: Adding uncertainty to the process. IEEE Software, 33(6):80–85, November/ December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060080-abs.
html.

Kwan:2012:CLR

[KCD12] Irwin Kwan, Marcelo Cataldo, and Daniela Damian. Conway’s Law revisited: The evidence for a task-based perspective. IEEE Software, 29(1): 90–93, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kim:2013:VSR

[KCK^{+}13] Youngjoo Kim, Okjoo Choi, Moonzoo Kim, Jongmoon Baik, and Tai-Hyo Kim. Validating software reliability early through statistical model checking. IEEE Software, 30(3):35–41, May/June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kalinowski:2012:EBG

[KCT12] Marcos Kalinowski, David N. Card, and Guilherme H. Travassos. Evidence-based guidelines to defect causal analysis. IEEE Software, 29(4):16–18, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Koziolek:2013:MAS

[KDGV13] Heiko Koziolek, Dominik Domis, Thomas Goldschmidt, and Philipp Vorst. Measuring architecture sustainability. IEEE Software, 30(6):54– 62, November/December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kassab:2017:STS

[KDL17] Mohamad Kassab, Joanna F. DeFranco, and Phillip A. Laplante. Software testing: The state of the practice. IEEE Software, 34(5):46–52, September/ October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050046-abs.
html.

Kuhrmann:2019:HSD

[KDM^{+}19] M. Kuhrmann, P. Diebold, J. Munch, P. Tell, K. Trektere, F. McCaffery, V. Garousi, M. Felderer, O. Linssen, E. Hanser, and C. R. Prause. Hybrid software development approaches in practice: A European perspective. IEEE Software, 36(4):20–31, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Koru:2010:TRD

[KE10] A. Gunes Koru and Khaled El Emam. The theory of relative dependency: Higher coupling concentration in smaller modules. IEEE Software, 27(2):81–89, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Keeling:2015:AHC

[Kee15a] Michael Keeling. Architecture haiku: A case study in lean documentation [the pragmatic architect]. IEEE Software, 32 (3):35–39, May/June 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        03/mso2015030035-abs.
                                        html.

Keeling:2015:LFE

[Kee15b] Michael Keeling. Lightweight and flexible: Emerging trends in software architecture from the SATURN conferences. IEEE Software, 32(3):7–11, May/ June 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/so/2015/03/mso2015030007.
                                        html.

Keeling:2018:DYT

[Kee18] M. Keeling. Design with your team, not for your team. IEEE Software, 35(4):86–88, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kersten:2017:VSA

[Ker17]  Mik Kersten.  Value

stream architecture. IEEE Software, 34(5):10–12, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/

05/mso2017050010-abs.
html.

Kersten:2018:CED

[Ker18a] M. Kersten. A Cambrian explosion of DevOps tools. IEEE Software, 35(2):14–17, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kersten:2018:FPC

[Ker18b] M. Kersten. Five predictions for the coming decades of software. IEEE Software, 35(5):7–9, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kersten:2018:MGT

[Ker18c] M. Kersten. Mining the ground truth of enterprise toolchains. IEEE Software, 35(3):12–17, May/ June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kersten:2018:WFT

[Ker18d] M. Kersten. What flows through a software value stream? IEEE Software, 35(4):8–11, July/ August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Knoche:2018:UML

[KH18] H. Knoche and W. Hasselbring. Using microservices for legacy software modernization. IEEE Software, 35(3):44–49, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kude:2019:SGS

[KHD19] T. Kude, T. Huber, and J. Dibbern. Successfully governing software ecosystems: Competence profiles of partnership managers. IEEE Software, 36 (3):39–44, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kazman:2018:MEC

[KHYT18] R. Kazman, S. Haziyev, A. Yakuba, and D. A. Tamburri. Managing energy consumption as an architectural quality attribute. IEEE Software, 35 (5):102–107, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kienle:2010:ATT

[Kie10] Holger M. Kienle. It’s about time to take JavaScript (more) seriously. IEEE Software, 27(3):60–62, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kimmel:2019:JNS

[Kim19] T. Kimmel. Johnathan Nightingale on scaling engineering management. IEEE Software, 36(3):120– 124, May/June 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Klein:2016:WMA

[Kle16]  John Klein.  What

makes an architect successful? IEEE Software, 33(1):20–22, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010020-abs.
html.

Kotonya:2011:SSD

[KLM11] Gerald Kotonya, Simon Lock, and John Mariani. Scrapheap software development: Lessons from an experiment on opportunistic reuse. IEEE Software, 28(2):68–74, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Knauss:2019:MRP

[Kna19] E. Knauss. The missing requirements perspective in large-scale agile system development. IEEE Software, 36(3):9–13, May/ June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Khalid:2016:ERB

[KNH16] Hammad Khalid, Meiyappan Nagappan, and Ahmed E. Hassan. Examining the relationship between Find-Bugs warnings and app ratings. IEEE Software, 33(4):34–39, July/ August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040034-abs.
html.

Kruchten:2012:TDM

[KNO12] Philippe Kruchten, Robert L. Nord, and Ipek Ozkaya. Technical debt: From metaphor to theory and practice. IEEE Software, 29(6):18–21, November/ December 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Komssi:2015:WH

[KPR^{+}15] Marko Komssi, Danielle Pichlis, Mikko Raatikainen, Klas Kindstrom, and Janne Jarvinen. What are hackathons for? IEEE Software, 32(5):60–67, September/October 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050060-abs.
html.

Kornstadt:2010:CSE

[KR10] Andreas Kornstadt and Eugen Reiswich. Composing systems with Eclipse rich client platform plugins. IEEE Software, 27(6):78–81, November/ December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Krishnamurthy:2017:BMW

[Kri17] Raghuraman Krishnamurthy.

Breezing my way as a solution architect: A retrospective on skill development and use. IEEE Software, 34(3):9–13, May/ June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL https://

www.computer.org/csdl/
mags/so/2017/03/mso2017030009-
abs.html.

Kruchten:2010:C

[Kru10] Philippe Kruchten. Certification 1, 2, 3. IEEE Software, 27(3):92–94, May/ June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kruchten:2015:LLL

[Kru15] Philippe Kruchten. Lifelong learning for lifelong employment. IEEE Software, 32(4):85–87, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        04/mso2015040085.html.

Khalid:2015:WDM

[KSNH15] Hammad Khalid, Emad Shihab, Meiyappan Nagappan, and Ahmed E. Hassan. What do mobile app users complain about? IEEE Software, 32 (3):70–77, May/June 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        03/mso2015030070-abs.
                                        html.

Kumar:2017:SAS

[KSTS17] Vinay Kumar, Lalit Kumar Singh, Anil Kumar Tripathi, and Pooja Singh. Safety analysis of safety-critical systems using state-space models. IEEE Software, 34(4): 38–47, July/August 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/

04/mso2017040038-abs.
html.

Kern:2010:DAN

[KSW10] Walter Kern, Christian Silberbauer, and Christian Wolff. The dimension architecture: a new approach to resource access. IEEE Software, 27(5):74–81, September/ October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Karac:2018:WDW

[KT18] I. Karac and B. Turhan.

What do we (really) know about test-driven development? IEEE Software, 35(4):81–85, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kelly:2011:SST

[KTH11] Diane Kelly, Stefan Thorsteinson, and Daniel Hook. Scientific software testing: Analysis with four dimensions. IEEE Software, 28 (3):84–90, May/June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Klotins:2018:SIP

[KUG18] E. Klotins, M. Unterkalmsteiner, and T. Gorschek. Software-intensive product engineering in start-ups: A taxonomy. IEEE Software, 35(4):44–52, July/ August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Klotins:2019:SEA

[KUG19] E. Klotins, M. Unterkalmsteiner, and T. Gorschek. Software engineering antipatterns in start-ups. IEEE Software, 36(2):118– 126, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kim:2011:MDP

[KV11] Wooyoung Kim and Michael Voss. Multicore desktop programming with Intel Threading Building Blocks. IEEE Software, 28(1):23–31, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kshetri:2018:BEV

[KV18] N. Kshetri and J. Voas. Blockchain-enabled e-voting. IEEE Software, 35(4):95– 99, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kolb:2010:PC

[KvdL10] Ronny Kolb and Frank van der Linden. Point/ counterpoint. IEEE Software, 27(3):56–59, May/ June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Khademi:2015:EEW

[KZW15] Amin Faiz Khademi, Mohammad Zulkernine, and Komminist Weldemariam. An empirical evaluation of Web-based fingerprinting. IEEE Software, 32(4): 46–52, July/August 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040046-abs.
html.

Lanovaz:2019:CCT

[LA19] M. J. Lanovaz and B. Adams.

Comparing the communication tone and responses of users and developers in two R mailing lists: Measuring positive and negative emails. IEEE Software, 36(5):46– 50, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Litecky:2010:MCJ

[LAAN10] Chuck Litecky, Andrew Aken, Altaf Ahmad, and H. James Nelson. Mining for computing jobs. IEEE Software, 27(1): 78–85, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lago:2010:GEI

[LAH10] Patricia Lago, Paris Avgeriou, and Rich Hilliard. Guest Editors’ introduction: Software architecture: Framing stakeholders’ concerns. IEEE Software, 27(6):20–24, November/December 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lattanze:2012:IAT

[Lat12] Anthony J. Lattanze. Infusing architectural thinking into organizations. IEEE Software, 29(1): 19–22, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Leicht:2017:LPC

[LBL17] Niklas Leicht, Ivo Blohm, and Jan Marco Leimeister. Leveraging the power of the crowd for software testing. IEEE Software, 34(2):62–69, March/April 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        02/mso2017020062-abs.
                                        html.

Larrucea:2017:RE

[LBWX17] Xabier Larrucea, Fabien Belmonte, Adam Welc, and Tao Xie. Reliability engineering. IEEE Software, 34(4):26–29, July/ August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040026.html.

Land:2011:ODW

[LC11] Rikard Land and Ivica Crnković. Oh dear, we bought our competitor: Integrating similar software systems. IEEE Software, 28(2):75–82, March/ April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Laplante:2017:DBM

[LCA17] Phillip Laplante, Steve Counsell, and Giuliano Antoniol. Disseminating the best material to practitioners. IEEE Software, 34(3):111–113, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030111.html.

Lanubile:2013:GAG

[LCE13] Filippo Lanubile, Fabio Calefato, and Christof Ebert. Group awareness in global software engineering. IEEE Software, 30(2):

18–23, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Larrucea:2013:SCS

[LCF13] Xabier Larrucea, Annie Combelles, and John Favaro. Safety-critical software [Guest Editors’ introduction]. IEEE Software, 30(3):25–27, May/ June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Larrucea:2017:SEI

[LCFT17] Xabier Larrucea, Annie Combelles, John Favaro, and Kunal Taneja. Software engineering for the Internet of Things. IEEE Software, 34(1):24–28, January/February 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        01/mso2017010024.html.

Lutz:2017:ROS

[LCH17] Robyn Lutz and Jane Cleland-Huang. The risk of overly strict requirements. IEEE Software, 34(2):26–29, March/ April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL https://

                                        www.computer.org/csdl/

mags/so/2017/02/mso2017020026-
abs.html.

Lin:2014:IAA

[LCYL14] Ying-Dar Lin, Edward T.- H. Chu, Shang-Che Yu, and Yuan-Cheng Lai. Improving the accuracy of automated GUI testing for embedded systems. IEEE Software, 31(1): 39–45, January/February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lichtenstein:2019:SCA

[LDBF19] Y. Lichtenstein, S. Dujmovic, and C. Baden-Fuller. Strategies for competing in the automotive industry’s software ecosystem: Standards and bottlenecks. IEEE Software, 36(3):45–49, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Levendovszky:2014:DRT

[LDO^{+}14] Tihamer Levendovszky, Abhishek Dubey, William R. Otte, Daniel Balasubramanian, Alessandro Coglio, Sandor Nyako, William Emfinger, Pranav Kumar, Aniruddha Gokhale, and Gabor Karsai. Distributed real-time managed systems: A model-driven distributed secure information architecture platform for managed embedded systems. IEEE Software, 31(2):62–69, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Louridas:2013:EAS

[LE13] Panos Louridas and Christof Ebert. Embedded analytics and statistics for big data. IEEE Software, 30(6):33–39, November/ December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Louridas:2016:ML

[LE16] Panos Louridas and Christof Ebert. Machine learning. IEEE Software, 33(5):110–115, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        05/mso2016050110-abs.
                                        html.

Lanubile:2010:CTG

[LEPV10] Filippo Lanubile, Christof Ebert, Rafael Prikladnicki, and Aurora Vizcaino. Collaboration tools for global software engineering. IEEE Software, 27(2):52–55, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Letouzey:2012:MTD

[LI12] Jean-Louis Letouzey and Michel Ilkiewicz. Managing technical debt with the SQALE method. IEEE Software, 29(6):44–51, November/December 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Li:2014:DPH

[Li14] Jingquan Li. Data protection in healthcare social networks. IEEE Software, 31(1):46–53, January/February 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lee:2013:SES

[LIL^{+}13] Dong-Hyun Lee, Hoh Peter In, Keun Lee, Sooyong Park, and Mike Hinchey. Sustainable embedded software life-cycle planning. IEEE Software, 30(4):72–80, July/August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lee:2010:CSO

[LK10] Jaejoon Lee and Gerald Kotonya. Combining service-orientation with product line engineering. IEEE Software, 27(3):35– 41, May/June 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lindman:2016:BRE

[LKR16] Juho Lindman, Tomi Kinnari, and Matti Rossi. Business roles in the emerging open-data ecosystem. IEEE Software, 33(5):54–59, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050054-abs.
html.

Lah:2016:HEA

[LL16] Urska Lah and James R. Lewis. How expertise affects a digital-rights-management-sharing application’s usability. IEEE Software, 33(3):76–82, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030076-abs.
html.

Loo:2012:UED

[LLC12] Keen Ngee Loo, Sai Peck Lee, and Thiam Kian Chiew. UML extension for defining the interaction variants of design patterns. IEEE Software, 29(5):64–72, September/ October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Laplante:2016:CUS

[LLV16] Nancy Laplante, Phillip A.

Laplante, and Jeffrey Voas. Caring: An undiscovered “Super-ility” of smart healthcare. IEEE Software, 33(6):16–19, November/December 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060016-abs.
html.

Leopold:2016:LQI

[LMG16] Henrik Leopold, Jan Mendling, and Oliver Gunther. Learning from quality issues of BPMN models from industry. IEEE Software, 33(4):26– 33, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040026-abs.
html.

Leppanen:2015:RSD

[LML^{+}15] Marko Leppanen, Simo Makinen, Samuel Lahtinen, Outi Sievi-Korte, Antti-Pekka Tuovinen, and Tomi Mannisto. Refactoring — a shot in the dark? IEEE Software, 32(6):62–70, November/ December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060062-abs.
html.

Lago:2015:RAA

[LMM^{+}15] Patricia Lago, Ivano Malavolta, Henry Muccini, Patrizio Pelliccione, and Antony Tang. The road ahead for architectural languages. IEEE Software, 32(1):98–105, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010098-abs.
html.

Laporte:2018:ASE

[LMMO18] Claude Y. Laporte, Mirna Munoz, Jezreel Mejia Miranda, and Rory V. OConnor. Applying software engineering standards in very small entities: From startups to grownups. IEEE Software, 35(1):99–103, January/February 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010099-abs.
html.

Leppanen:2015:HCR

[LMP^{+}15] Marko Leppanen, Simo Makinen, Max Pagels, Veli-Pekka Eloranta, Juha Itkonen, Mika V. Mantyla, and Tomi Mannisto. The highways and country roads to continuous deployment. IEEE Software, 32(2):64–72, March/ April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http://

www.computer.org/csdl/
mags/so/2015/02/mso2015020064-
abs.html.

Lanza:2016:TDP

[LMP16] Michele Lanza, Andrea Mocci, and Luca Ponzanelli. The tragedy of defect prediction, prince of empirical software engineering research. IEEE Software, 33(6):102–105, November/December 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060102.html.

Lewis:2011:SOS

[LMSS11] Grace Lewis, Edwin Morris, Soumya Simanta, and Dennis Smith. Service orientation and systems of systems. IEEE Software, 28(1):58–63, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Luk:2011:SAT

[LNH^{+}11] Chi-Keung Luk, Ryan Newton, William Hasenplaugh, Mark Hampton, and Geoff Lowney. A synergetic approach to throughput computing on x86-based multicore desktops. IEEE Software, 28(1):39–50, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lobur:2011:SCC

[Lob11] Julia M. Lobur. The success of a COTS caseload management system in state government. IEEE Software, 28(2):10–14, March/April 2011. CO-DEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Larrucea:2016:SPI

[LOCPL16] Xabier Larrucea, Rory V. O’Connor, Ricardo Colomo-Palacios, and Claude Y. Laporte. Software process improvement in very small organizations. IEEE Software, 33(2):85–89, March/April 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/

02/mso2016020085-abs.
html.

Louridas:2010:AMY

[Lou10] Panos Louridas. Up in the air: Moving your applications to the cloud. IEEE Software, 27(4):6– 11, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Louridas:2011:TM

[Lou11] Panos Louridas. Test management. IEEE Software, 28(5):86–91, September/ October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Louridas:2016:CSE

[Lou16] Panos Louridas. Component stacks for enterprise applications. IEEE Software, 33(2):93–98, March/ April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http:

//csdl.computer.org/
csdl/mags/so/2016/02/
mso2016020093.html.

Luthria:2012:SOA

[LR12] Haresh Luthria and Fethi A.

Rabhi. Service-oriented architectures: Myth or reality? IEEE Software, 29(4): 46–52, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lu:2015:RFS

[LRB^{+}15] Jian Lu, David S. Rosenblum, Tevfik Bultan, Valerie Issarny, Schahram Dustdar, Margaret-Anne Storey, and Dongmei Zhang. Roundtable: The future of software engineering for Internet computing. IEEE Software, 32(1):91–97, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010091-abs.
html.

Lindman:2011:MOS

[LRP11] Juho Lindman, Matti Rossi, and Anna Puustell. Matching open source software licenses with corresponding business models. IEEE Software, 28(4): 31–35, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Li:2012:EDT

[LSCK12] Jingyue Li, Tor Stalhane, Reidar Conradi, and Jan M. W. Kristiansen. Enhancing defect tracking systems to facilitate software quality improvement. IEEE Software, 29(2):59– 66, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Larrucea:2018:M

[LSCPE18] X. Larrucea, I. Santamaria, R. Colomo-Palacios, and C. Ebert. Microservices. IEEE Software, 35(3):96–100, May/ June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lebeuf:2018:SB

[LSZ18] Carlene Lebeuf, Margaret-Anne Storey, and Alexey Zagalsky. Software bots. IEEE Software, 35(1): 18–23, January/February 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010018-abs.
html.

Lim:2012:BAW

[LTS12] Erin Lim, Nitin Taksande, and Carolyn Seaman. A balancing act: What software practitioners have to say about technical debt. IEEE Software, 29(6):22– 27, November/December 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lewicki:2019:RTO

[LTv19] P. Lewicki, J. Tochowicz, and J. van Genuchten. Are robots taking our jobs? A RoboPlatform at a bank. IEEE Software, 36(3):101– 104, May/June 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

LaToza:2016:CSE

[LvdH16] Thomas D. LaToza and Andre van der Hoek. Crowdsourcing in software engineering: Models, motivations, and challenges. IEEE Software, 33(1):74–80, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010074.html.

Lewis:2011:RGR

[LW11] Chris Lewis and Jim Whitehead. Repairing games at runtime or, how we learned to stop worrying and love emergence. IEEE Software, 28(5):53– 59, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Larrucea:2017:SMR

[LWCP17] Xabier Larrucea, Alastair Walker, and Ricardo Colomo-Palacios. Supporting the management of reusable automotive software. IEEE Software, 34(3):40–47, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030040-abs.
html.

Lu:2017:ABB

[LX17] Qinghua Lu and Xiwei Xu. Adaptable blockchainbased systems: A case study for product traceability. IEEE Software, 34(6):21–27, November/ December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060021-abs.
html.

Lu:2015:TTC

[LXB^{+}15] Qinghua Lu, Xiwei Xu, Len Bass, Liming Zhu, and Weishan Zhang. A tail-tolerant cloud API wrapper. IEEE Software, 32(1):76–82, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010076-abs.
html.

Liu:2015:DEE

[LXC15] Yepang Liu, Chang Xu, and Shing-Chi Cheung. Diagnosing energy efficiency and performance for mobile Internetware applications. IEEE Software, 32(1):67–75, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        01/mso2015010067-abs.
                                        html.

Madison:2010:AAI

[Mad10] James Madison. Agile architecture interactions. IEEE Software, 27(2):41– 48, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mellblom:2019:CBB

[MAGT19] E. Mellblom, I. Arason, L. Gren, and R. Torkar. The connection between burnout and personality types in software developers. IEEE Software, 36(5):57–64, September/ October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maiden:2010:SDA

[Mai10a]  Neil Maiden.  Service

design: It’s all in the brand. IEEE Software, 27(5):18–19, September/ October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2010:TMI

[Mai10b] Neil Maiden. Trust me, I’m an analyst. IEEE Software, 27(1):46–47, January/February 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2011:IA

[Mai11a] Neil Maiden. The inhibited analyst. IEEE Software, 28(6):100–102, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2011:RA

[Mai11b] Neil Maiden. Requirements and aesthetics. IEEE Software, 28(3):20– 21, May/June 2011. CO-DEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2011:WTI

[Mai11c] Neil Maiden. What time is it, Eccles? IEEE Software, 28(4):84–85, July/ August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maiden:2012:CA

[Mai12a] Neil Maiden. Cherishing ambiguity. IEEE Software, 29(6):16–17, November/December 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maiden:2012:EHR

[Mai12b] Neil Maiden. Exactly how are requirements written? IEEE Software, 29(1): 26–27, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maiden:2012:FRW

[Mai12c] Neil Maiden. Framing requirements work as learning. IEEE Software, 29 (3):8–9, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maiden:2012:PI

[Mai12d]  Neil Maiden.  Politics

are inescapable. IEEE Software, 29(5):88–90, September/October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maiden:2012:SKR

[Mai12e] Neil Maiden. Spocks and Kirks in the requirements universe. IEEE Software, 29(4):84–85, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maiden:2013:MOR

[Mai13a]  Neil Maiden.  Moni-

toring our requirements. IEEE Software, 30(1): 16–17, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maiden:2013:WRW

[Mai13b] Neil Maiden. So, what is requirements work? IEEE Software, 30(2):14– 15, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mojica:2014:LSE

[MAN^{+}14] Israel J. Mojica, Bram Adams, Meiyappan Nagappan, Steffen Dienst, Thorsten Berger, and Ahmed E. Hassan. A large-scale empirical study on software reuse in mobile apps. IEEE Software, 31(2):78–86, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Marasco:2012:SBN

[Mar12] Joe Marasco. Silver bullets: No secret ingredients. IEEE Software, 29(3):102– 104, May/June 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mavin:2012:LTU

[Mav12]  Alistair Mavin.  Lis-

ten, then use EARS. IEEE Software, 29(2):17– 18, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mahfouz:2010:RDD

[MBLN10] Ayman Mahfouz, Leonor Barroca, Robin Laney, and Bashar Nuseibeh. Requirements-driven design of service-oriented interactions. IEEE Software, 27(6):25–32, November/ December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Muccini:2018:CMS

[MBv18] H. Muccini, J. Bosch, and A. van der Hoek. Collaborative modeling in software engineering. IEEE Software, 35(6):20–24, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Misirli:2013:RSS

[MCBT13] Ayse Tosun Misirli, Bora Caglayan, Ayse Bener, and Burak Turhan. A retrospective study of software analytics projects: In-depth interviews with practitioners. IEEE Software, 30(5):54– 61, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mens:2019:SDF

[MCD19] T. Mens, M. Cataldo, and D. Damian. The social developer: The future of software development [Guest Editors’ introduction]. IEEE Software, 36(1):11–14, January/February 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

McGraw:2011:TTS

[McG11]  Gary McGraw.  Tech-

nology transfer: a software security marketplace case study. IEEE Software, 28(5):9–11, September/October 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mirakhorli:2013:TTP

[MCH13] Mehdi Mirakhorli and Jane Cleland-Huang. Traversing the twin peaks. IEEE Software, 30(2):30–36, March/April 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mader:2015:RPD

[MCH15] Patrick Mader and Jane Cleland-Huang. From raw project data to business intelligence. IEEE Software, 32(4):22–25, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040022-abs.
html.

Mens:2017:MMC

[MCHK17] Kim Mens, Rafael Capilla, Herman Hartmann, and Thomas Kropf. Modeling and managing contextaware systems’ variability. IEEE Software, 34(6):58–63, November/ December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060058.html.

McHugh:2012:API

[MCL12] Orla McHugh, Kieran Conboy, and Michael Lang. Agile practices: The impact on trust in software project teams. IEEE Software, 29(3):71– 76, May/June 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Michael:2011:VVT

[MDOS11] James Bret Michael, Doron Drusinsky, Thomas W. Otani, and Man-Tak Shing. Verification and validation for trustworthy software systems. IEEE Software, 28(6):86– 92, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Menzies:2013:BDM

[Men13] Tim Menzies. Beyond data mining. IEEE Software, 30(3):92, May/June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Menzies:2018:UES

[Men18] T. Menzies. The unreasonable effectiveness of software analytics. IEEE Software, 35(2):96–98, March/ April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Meyer:2014:CIT

[Mey14a] Mathias Meyer. Continuous integration and its tools. IEEE Software, 31 (3):14–16, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Meyerson:2014:GPL

[Mey14b] Jeff Meyerson. The Go programming language. IEEE Software, 31(5):104, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        05/mso2014050104.html.

Meyerson:2016:BHA

[Mey16] Jeff Meyerson. Ben Hindman on Apache Mesos. IEEE Software, 33(1):117– 120, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/so/2016/
                                        01/mso2016010117.html.

Meyer:2018:MSA

[Mey18] B. Meyer. Making sense of agile methods. IEEE Software, 35(2):91–94, March/ April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mancl:2013:STR

[MF13] Dennis Mancl and Steven Fraser. Software tools research: SPLASH panel discussion. IEEE Software, 30(2):16–17, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Michlmayr:2015:WHS

[MFS15] Martin Michlmayr, Brian Fitzgerald, and Klaas-Jan Stol. Why and how should open source projects adopt time-based releases? IEEE Software, 32(2):55–63, March/ April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http://

www.computer.org/csdl/
mags/so/2015/02/mso2015020055-
abs.html.

Martins:2017:RES

[MG17] Luiz Eduardo G. Martins and Tony Gorschek. Requirements engineering for safety-critical systems: Overview and challenges. IEEE Software, 34(4):49– 57, July/August 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040049-abs.
html.

Mens:2010:GEI

[MGFRD10] Tom Mens, Yann-Gael Gueheneuc, Juan Fernandez-Ramil, and Maja D’Hondt. Guest Editors’ introduction: Software evolution. IEEE Software, 27(4):22– 25, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Menasce:2011:SFS

[MGMS11] Daniel Menasce, Hassan Gomaa, Sam Malek, and Joao Sousa. SASSY: a framework for selfarchitecting service-oriented systems. IEEE Software, 28(6):78–85, November/ December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

MacLeod:2018:CRT

[MGS^{+}18a] L. MacLeod, M. Greiler, M. Storey, C. Bird, and J. Czerwonka. Code reviewing in the trenches: Challenges and best practices. IEEE Software, 35(4):34–42, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mead:2018:HCS

[MGS18b] N. R. Mead, D. Garlan, and M. Shaw. Half a century of software engineering education: The CMU exemplar. IEEE Software, 35(5):25–31, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Murer:2014:FYS

[MH14] Stephan Murer and Claus Hagen. Fifteen years of service-oriented architecture at Credit Suisse. IEEE Software, 31(6):9–15, November/ December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        06/mso2014060009-abs.
                                        html.

Morin:2017:MBS

[MHF17] Brice Morin, Nicolas Harrand, and Franck Fleurey. Model-based software engineering to tame the IoT jungle. IEEE Software, 34(1):30–36, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010030-abs.
html.

Mao:2017:RTM

[MHJ17] Ke Mao, Mark Harman, and Yue Jia. Robotic testing of mobile apps for truly black-box automation. IEEE Software, 34(2):11– 16, March/April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020011-abs.
html.

Miranskyy:2016:OLA

[MHLCL16] Andriy Miranskyy, Abdelwahab Hamou-Lhadj, Enzo Cialini, and Alf Larsson. Operational-log analysis for big data systems: Challenges and solutions. IEEE Software, 33(2):52– 59, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        02/mso2016020052-abs.
                                        html.

Maipradit:2019:SCU

[MHM19] R. Maipradit, H. Hata, and K. Matsumoto. Sentiment classification using N-gram inverse document frequency and automated machine learning. IEEE Software, 36(5):65– 70, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Murphy-Hill:2015:RGE

[MHRSO15]  Emerson  Murphy-Hill,

Don Roberts, Peter Sommerlad, and William F. Opdyke. Refactoring [Guest Editors’ introduction]. IEEE Software, 32(6):27–29, November/ December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        06/mso2015060027.html.

Mueller:2019:ASE

[MIC19] D. Mueller and D. Izquierdo-Cortazar. From art to science: The evolution of community development. IEEE Software, 36(6):23– 28, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maiden:2010:AR

[MJ10] Neil Maiden and Sara Jones. Agile requirements. IEEE Software, 27(3):87– 88, May/June 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mader:2013:STS

[MJZCH13] Patrick Mader, Paul L.

Jones, Yi Zhang, and Jane Cleland-Huang. Strategic traceability for safetycritical projects. IEEE Software, 30(3):58–66, May/June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Madria:2014:SCC

[MKD14] Sanjay Madria, Vimal Kumar, and Rashmi Dalvi. Sensor cloud: A cloud of virtual sensors. IEEE Software, 31(2):70–77, March/ April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Moy:2013:TFV

[MLD^{+}13] Yannick Moy, Emmanuel Ledinot, Herve Delseny, Virginie Wiels, and Benjamin Monate. Testing or formal verification: DO-178C alternatives and industrial experience. IEEE Software, 30 (3):50–57, May/June 2013.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mijailovic:2013:RUI

[MM13] Zarko Mijailovic and Dragan Milicev. A retrospective on user interface development technology. IEEE Software, 30(6):76– 83, November/December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mathew:2018:SET

[MM18a] G. Mathew and T. Menzies. Software engineering’s top topics, trends, and researchers. IEEE Software, 35(5):88–93, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maudoux:2018:CET

[MM18b] G. Maudoux and K. Mens. Correct, efficient, and tailored: The future of build systems. IEEE Software, 35(2):32–37, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Martens:2019:RER

[MM19] D. Martens and W. Maalej. Release early, release often, and watch your users’ emotions: Lessons from emotional patterns.

IEEE Software, 36(5):32– 37, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

McGregor:2010:GEI

[MMYJ10] John D. McGregor, Dirk Muthig, Kentaro Yoshimura, and Paul Jensen. Guest Editors’ introduction: Successful software product line practices. IEEE Software, 27(3):16–21, May/ June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Maalej:2016:TDD

[MNJR16] Walid Maalej, Maleknaz Nayebi, Timo Johann, and Guenther Ruhe. Toward data-driven requirements engineering. IEEE Software, 33(1):48–54, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010048-abs.
html.

Maglyas:2012:LSS

[MNS12] Andrey Maglyas, Uolevi Nikula, and Kari Smolander. Lean solutions to software product management problems. IEEE Software, 29(5):40–46, September/ October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Makitalo:2018:SSE

[MOK^{+}18] Niko Makitalo, Aleksandr Ometov, Joona Kannisto, Sergey Andreev, Yevgeni Koucheryavy, and Tommi Mikkonen. Safe, secure executions at the network edge: Coordinating cloud, edge, and fog computing. IEEE Software, 35(1):30–37, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2018/
                                        01/mso2018010030-abs.
                                        html.

Monden:2011:GGU

[MOMM11] Akito Monden, Satoshi Okahara, Yuki Manabe, and Kenichi Matsumoto. Guilty or not guilty: Using clone metrics to determine open source licensing violations. IEEE Software, 28(2):42–47, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Moore:2016:FTL

[MONF16] Andrew Moore, Tim O’Reilly, Paul D. Nielsen, and Kevin Fall. Four thought leaders on where the industry is headed.

IEEE Software, 33(1): 36–39, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010036.html.

Morozoff:2010:ULC

[Mor10] Edmund Morozoff. Using a line-of-code metric to understand software rework. IEEE Software, 27(1):72–77, January/February 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mossinger:2010:SAS

[Mos10] Jurgen Mossinger. Software in automotive systems. IEEE Software, 27(2):92–94, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Marr:2014:WTY

[MPD14] Stefan Marr, Tobias Pape, and Wolfgang De Meuter. Are we there yet?: Simple language implementation techniques for the 21st Century. IEEE Software, 31(5):60–67, September/ October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/

                                        05/mso2014050060-abs.
                                        html.

Musson:2013:LCH

[MRF^{+}13] Robert Musson, Jacqueline Richards, Danyel Fisher, Christian Bird, Brian Bussone, and Sandipan Ganguly. Leveraging the crowd: How 48,000 users helped improve Lync performance. IEEE Software, 30(4):38–45, July/ August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mohan:2010:ISP

[MRS10] Kannan Mohan, Balasubramaniam Ramesh, and Vijayan Sugumaran. Integrating software product line engineering and agile development. IEEE Software, 27(3):48–55, May/ June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Morris:2012:LLS

[MS12] Chris Morris and Judith Segal. Lessons learned from a scientific software development project. IEEE Software, 29(4):9– 12, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mathiassen:2014:PMC

[MS14] Lars Mathiassen and Anna Borjesson Sandberg. Process mass customization in a global software firm. IEEE Software, 31(6):62–69, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060062-abs.
html.

McIlroy:2017:IWR

[MSAH17] Stuart McIlroy, Weiyi Shang, Nasir Ali, and Ahmed E. Hassan. Is it worth responding to reviews? Studying the top free apps in Google Play. IEEE Software, 34(3):64– 71, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030064-abs.
html.

Mitchell:2010:FTL

[MSS10] Nick Mitchell, Edith Schonberg, and Gary Sevitsky. Four trends leading to Java runtime bloat. IEEE Software, 27(1): 56–63, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Moreno:2014:PIA

[MSSMDC14] Ana M. Moreno, Maria-Isabel Sanchez-Segura, Fuensanta Medina-Dominguez, and Gonzalo Cuevas. Process improvement from an academic perspective: How could software engineering education contribute to CMMI practices? IEEE Software, 31(4):91–97, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mikkonen:2019:SRE

[MT19] T. Mikkonen and A. Taivalsaari. Software reuse in the era of opportunistic design. IEEE Software, 36(3):105–111, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Marsan:2019:TSS

[MTM^{+}19] J. Marsan, M. Templier, P. Marois, B. Adams, K. Carillo, and G. L. Mopenza. Toward solving social and technical problems in open source software ecosystems: Using cause-and-effect analysis to disentangle the causes of complex problems. IEEE Software, 36(1):34–41, January/February 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Murguzur:2015:RTV

[MTT^{+}15] Aitor Murguzur, Salvador Trujillo, Hong-Linh Truong, Schahram Dustdar, Oscar Ortiz, and Goiuria Sagardui. Run-time variability for context-aware smart workflows. IEEE Software, 32 (3):52–60, May/June 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030052-abs.
html.

Maiden:2011:RAN

[MW11] Neil Maiden and Anja Wever. Requirements analysis: The next generation. IEEE Software, 28(2):22–23, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

MavinMav:2019:TYE

[MW19] A. Mavin Mav and P. Wilkinson. Ten years of EARS. IEEE Software, 36(5):10– 14, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Menzies:2013:MFS

[MZ13a] Tim Menzies and Thomas Zimmermann. The many faces of software analytics. IEEE Software, 30(5):28– 29, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Menzies:2013:SAW

[MZ13b] Tim Menzies and Thomas Zimmermann. Software analytics: So what? IEEE Software, 30(4):31– 37, July/August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Menzies:2018:SAW

[MZ18] T. Menzies and T. Zimmermann. Software analytics: What’s next? IEEE Software, 35(5):64– 70, September/October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Meservy:2012:BRA

[MZLD12] Thomas O. Meservy, Chen Zhang, Euntae T. Lee, and Jasbir Dhaliwal. The business rules approach and its effect on software testing. IEEE Software, 29(4): 60–66, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nagappan:2018:RWG

[Nag18] M. Nagappan. Reconsidering whether GOTO is harmful. IEEE Software, 35(3):93–95, May/June 2018. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Niu:2018:REC

[NBF^{+}18] N. Niu, S. Brinkkemper, X. Franch, J. Partanen, and J. Savolainen. Requirements engineering and continuous deployment. IEEE Software, 35(2):86–90, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nagy:2016:OEN

[NBS^{+}16] Janos Nagy, Kalman Balajthy, Sandor Szalai, Balint Sodor, Istvan Horvath, and Csaba Lipusz. Obstanovka: Exploring nearby space. IEEE Software, 33(4):101–105, July/August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040101-abs.
html.

Nunes:2011:IEI

[NCK11] Nuno Jardim Nunes, Larry Constantine, and Rick Kazman. iUCP: Estimating interactive-software project size with enhanced use-case points. IEEE Software, 28(4):64– 73, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Nabulsi:2015:TFC

[NH15] Mohammad A. Nabulsi and Robert M. Hierons. A test framework for communications-critical largescale systems. IEEE Software, 32(3):86–93, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        03/mso2015030086-abs.
                                        html.

Nakasai:2019:DBA

[NHM19] K. Nakasai, H. Hata, and K. Matsumoto. Are donation badges appealing?: A case study of developer responses to Eclipse bug reports. IEEE Software, 36 (3):22–27, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nichols:2019:EMI

[Nic19] W. R. Nichols. The end to the myth of individual programmer productivity. IEEE Software, 36(5):71–75, September/ October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nayebi:2018:HLN

[NKR^{+}18] Maleknaz Nayebi, Shaikh Jeeshan Kabeer, Guenther Ruhe, Chris Carlson, and Francis Chew. Hybrid labels are the new measure! IEEE Software, 35(1):54–57, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010054-abs.
html.

Nord:2012:MAV

[NOS12] Robert L. Nord, Ipek Ozkaya, and Raghvinder S. Sangwan. Making architecture visible to improve flow management in lean software development. IEEE Software, 29(5):33–39, September/ October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Neema:2019:BRA

[NPJ19] S. Neema, R. Parikh, and S. Jagannathan. Building resource adaptive software systems. IEEE Software, 36(2):103–109, March/April 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nowak:2017:TEB

[NS17] Artur Nowak and Holger J. Schünemann. Toward evidence-based software engineering: Lessons learned in healthcare application development. IEEE Software, 34(5):67–71, September/ October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050067-abs.
html.

Novielli:2019:SES

[NS19] N. Novielli and A. Serebrenik. Sentiment and emotion in software engineering. IEEE Software, 36(5):6–23, September/October 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nakakoji:2012:TUS

[NYMS12] Kumiyo Nakakoji, Yasuhiro Yamamoto, Nobuto Matsubara, and Yoshinari Shirai. Toward unweaving streams of thought for reflection in professional software design. IEEE Software, 29(1): 34–38, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Obrenovic:2013:SSB

[Obr13] Zeljko Obrenovic. Software sketchifying: Bringing innovation into software development. IEEE Software, 30(3):80–86, May/June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Obrenovic:2017:IPI

[Obr17]  Zeljko Obrenovic.  In-

sights from the past: The IEEE Software history experiment. IEEE Software, 34(4):71–78, July/ August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040071-abs.
html.

Obrenovic:2018:QIS

[Obr18] Željko Obrenović. Quotes from IEEE Software history. IEEE Software, 35(5):10–13, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ortin:2014:RPI

[OCFLI14] Francisco Ortin, Patricia Conde, Daniel Fernandez-Lanvin, and Raul Izquierdo. The runtime performance

of invokedynamic: An

evaluation with a Java library. IEEE Software, 31(4):82–90, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Offutt:2013:PES

[Off13] Jeff Offutt. Putting the engineering into software engineering education. IEEE Software, 30(1):96, January/February 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ochoa:2017:WDC

[ORM^{+}17] Sergio F. Ochoa, Romain Robbes, Maira Marques, Luis Silvestre, and Alcides Quispe. What differentiates Chilean niche software companies: Business knowledge and reputation. IEEE Software, 34(3):96– 103, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030096-abs.
html.

Osbourn:2011:GMW

[Osb11] Toby Osbourn. Getting the most out of the Web. IEEE Software, 28(1): 96, 95, January/February 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Olsson:2014:SRL

[OSBA14] Helena Olsson, Anna Sandberg, Jan Bosch, and Hiva Alahyari. Scale and responsiveness in largescale software development. IEEE Software, 31(5):87–93, September/ October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050087-abs.
html.

Ojala:2011:DCB

[OT11] Arto Ojala and Pasi Tyrvainen. Developing cloud business models: a case study on cloud gaming. IEEE Software, 28(4):42– 47, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ozkaya:2019:DAO

[Ozk19a] I. Ozkaya. Are DevOps and automation our next silver bullet? IEEE Software, 36(4):3–95, July/ August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2019:ESD

[Ozk19b] I. Ozkaya. Ethics is a software design concern. IEEE Software, 36(3):4–8, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2019:GAS

[Ozk19c] I. Ozkaya. The golden age of software engineering [from the Editor]. IEEE Software, 36(1):4–10, January/February 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2019:IID

[Ozk19d] I. Ozkaya. If it does not scale, it does not work! IEEE Software, 36(2):4–7, March/April 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2019:ICD

[Ozk19e] I. Ozkaya. Interact, collaborate, debate. IEEE Software, 36(6):3–6, November/December 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2019:VD

[Ozk19f] I. Ozkaya. The voice of the developer. IEEE Software, 36(5):3–5, September/October 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Papazoglou:2011:MES

[PAB11] Michael P. Papazoglou, Vasilios Andrikopoulos, and Salima Benbernou. Managing evolving services. IEEE Software, 28 (3):49–55, May/June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Parnas:2018:SSC

[Par18] David Lorge Parnas. Software structures: A careful look. IEEE Software, 35(6):68–71, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Prikladnicki:2010:PED

[PAS10] Rafael Prikladnicki, Jorge Luis Nicolas Audy, and Forrest Shull. Patterns in effective distributed software development. IEEE Software, 27(2):12–15, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pinciroli:2016:BPL

[PB16] Carlo Pinciroli and Giovanni Beltrame. Buzz: A programming language for robot swarms. IEEE Software, 33(4):97–100, July/August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        04/mso2016040097-abs.
                                        html.

Peng:2014:CSD

[PBE14] Xin Peng, Muhammad Ali Babar, and Christof Ebert. Collaborative software development platforms for crowdsourcing. IEEE Software, 31(2):30–36, March/ April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Poppendieck:2012:LSD

[PC12] Mary Poppendieck and Michael A. Cusumano. Lean software development: a tutorial. IEEE Software, 29(5):26–32, September/October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Perez-Castillo:2011:RT

[PCdGPE11] Ricardo Perez-Castillo, Ignacio Garcia-Rodriguez de Guzman, Mario Piattini, and Christof Ebert. Reengineering technologies. IEEE Software, 28(6):13–17, November/ December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Perez-Castillo:2014:AHE

[PCP14] Ricardo Perez-Castillo and Mario Piattini. Analyzing the harmful effect of God class refactoring on power consumption. IEEE Software, 31(3):48–54, May/ June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Perez-Castillo:2019:EA

[PCRPE19] R. Perez-Castillo, F. Ruiz, M. Piattini, and C. Ebert. Enterprise architecture. IEEE Software, 36(4):12– 19, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Passos:2012:OCS

[PDNdSB12] Odette Mestrinho Passos, Arilo Claudio Dias-Neto, and Raimundo da Silva Barreto. Organizational culture and success in SPI initiatives. IEEE Software, 29(3):97– 99, May/June 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Penzenstadler:2018:SES

[PDV^{+}18] B. Penzenstadler, L. Duboc, C. C. Venters, S. Betz, N. Seyff, K. Wnuk, R. Chitchyan, S. M. Easterbrook, and C. Becker. Software engineering for sustainability: Find the leverage points! IEEE Software, 35(4):22–33, July/August 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pech:2013:SAI

[Pec13]  Stephan Pech.  Soft-

ware agents in industrial automation systems. IEEE Software, 30(3):20– 24, May/June 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pelkola:2012:FMP

[Pel12] David Pelkola. A framework for managing privacyenhancing technology. IEEE Software, 29(3):45–49, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Penzenstadler:2015:SRM

[Pen15] Birgit Penzenstadler. Sustainability and requirements: A manifesto. IEEE Software, 32(5):90–92, September/October 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        05/mso2015050090.html.

Piedrahita:2018:LSD

[PGG^{+}18] Andres F. Murillo Piedrahita, Vikram Gaur, Jairo Giraldo, Alvaro A. Cardenas, and Sandra Julieta Rueda. Leveraging software-defined networking for incident response in industrial control systems. IEEE Software, 35(1):44–50, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010044-abs.
html.

Parnin:2017:TAC

[PHA^{+}17] Chris Parnin, Eric Helms, Chris Atlee, Harley Boughton, Mark Ghattas, Andy Glover, James Holman, John Micco, Brendan Murphy, Tony Savor, Michael Stumm, Shari Whitaker, and Laurie Williams. The top 10 adages in continuous deployment. IEEE Software, 34(3):86–95, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030086-abs.
html.

Pang:2016:WDP

[PHAH16] Candy Pang, Abram Hindle, Bram Adams, and Ahmed E. Hassan. What do programmers know about software energy consumption? IEEE Software, 33(3):83–89, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2016/
                                        03/mso2016030083-abs.
                                        html.

Paasivaara:2019:EYA

[PL19] M. Paasivaara and C. Lassenius. Empower your agile organization: Communitybased decision making in large-scale agile development at Ericsson. IEEE Software, 36(2):64– 69, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Prikladnicki:2018:TAU

[PLC18] Rafael Prikladnicki, Casper Lassenius, and Jeffrey C. Carver. Trends in agile updated: Perspectives from the practitioners. IEEE Software, 35(1):109–111, January/February 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2018/
                                        01/mso2018010109.html.

Prikladnicki:2019:TAO

[PLC19] R. Prikladnicki, C. Lassenius, and J. C. Carver. Trends in agile: From operational to strategic agility [practitioners’ digest]. IEEE Software, 36(1):95–97, January/ February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Prikladnicki:2016:TAP

[PLTC16] Rafael Prikladnicki, Casper Lassenius, Evelyn Tian, and Jeffrey C. Carver. Trends in agile: Perspectives from the practitioners. IEEE Software, 33(6):20–22, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060020.html.

Persson:2010:PMR

[PM10] John Stouby Persson and Lars Mathiassen. A process for managing risks in distributed teams. IEEE Software, 27(1):20–29, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Prikladnicki:2018:VER

[PM18] Rafael Prikladnicki and Tim Menzies. From voice of evidence to redirections. IEEE Software, 35(1):11–13, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2018/
                                        01/mso2018010011-abs.
                                        html.

Prikladnicki:2012:TSC

[PMCE12] Rafael Prikladnicki, Sabrina Marczak, Erran Carmel, and Christof Ebert. Technologies to support collaboration across time zones. IEEE Software, 29(3):10– 13, May/June 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Parsons:2014:CRP

[PMSL14] David Parsons, Anuradha Mathrani, Teo Susnjak, and Arno Leist. Coderetreats: Reflective practice and the Game of Life. IEEE Software, 31(4):58–64, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Poort:2014:DAA

[Poo14] Eltjo R. Poort. Driving agile architecting with cost and risk. IEEE Software, 31(5):20–23, September/ October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        05/mso2014050020-abs.
                                        html.

Poort:2016:JEA

[Poo16] Eltjo Poort. Just enough anticipation: Architect your time dimension. IEEE Software, 33(6):11– 15, November/December 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060011-abs.
html.

Popp:2011:SIB

[Pop11] Karl Michael Popp. Software industry business models. IEEE Software, 28(4):26–30, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Prikladnicki:2017:BSD

[PPMD17] Rafael Prikladnicki, Marcelo Gattermann Perin, Sabrina Marczak, and Alessandra Costa Smolenaars Dutra. The best software development teams might be temporary. IEEE Software, 34(2):22–25, March/April 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020022-abs.
html.

Prasad:2017:DRS

[Pra17] Balaji Prasad. Darkitecture: The reality skirted by architecture. IEEE Software, 34(1):103–105, January/February 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010103-abs.
html.

Polo:2013:TA

[PRPE13] Macario Polo, Pedro Reales, Mario Piattini, and Christof Ebert. Test automation. IEEE Software, 30(1):84–89, January/February 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Penzenstadler:2014:SSN

[PRRT14] Birgit Penzenstadler, Ankita Raturi, Debra Richardson, and Bill Tomlinson. Safety, security, now sustainability: The nonfunctional requirement for the 21st Century. IEEE Software, 31(3):40–47, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Poth:2014:EQM

[PS14] Alexander Poth and Ali Sunyaev. Effective quality management: Value- and risk-based software quality management. IEEE Software, 31(6):79–85, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060079-abs.
html.

Perez-Soler:2018:CMG

[PSGd18] S. Pérez-Soler, E. Guerra, and J. de Lara. Collaborative modeling and group decision making using chatbots in social networks. IEEE Software, 35(6):48–54, November/ December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pankratius:2011:GEI

[PSK11] Victor Pankratius, Wolfram Schulte, and Kurt Keutzer. Guest Editors’ introduction: Parallelism on the desktop. IEEE Software, 28(1):14–16, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Prodan:2012:EHP

[PSO12] Radu Prodan, Michael Sperk, and Simon Ostermann. Evaluating highperformance computing on Google App Engine. IEEE Software, 29(2):52– 58, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pike:2017:SAS

[PST^{+}17] Lee Pike, Jamey Sharp, Mark Tullsen, Patrick C. Hickey, and James Bielman. Secure automotive software: The next steps. IEEE Software, 34 (3):49–55, May/June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030049-abs.
html.

Port:2018:AAS

[PT18] Dan Port and Bill Taber. Actionable analytics for strategic maintenance of critical software: An industry experience report. IEEE Software, 35(1):58–63, January/ February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010058-abs.
html.

Pe-Than:2019:DCH

[PTNF^{+}19] E. P. P. Pe-Than, A. Nolte, A. Filippova, C. Bird, S. Scallen, and J. D. Herbsleb. Designing corporate hackathons with a purpose: The future of software development.

IEEE Software, 36(1): 15–22, January/February 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Passos:2010:SAC

[PTV^{+}10] Leonardo Passos, Ricardo Terra, Marco Tulio Valente, Renato Diniz, and Nabor Mendonca. Static architecture-conformance checking: An illustrative overview. IEEE Software, 27(5):82–89, September/ October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Papazoglou:2015:RAK

[PvdHM15] Michael P. Papazoglou, Willem-Jan van den Heuvel, and Julien Etienne Mascolo. A reference architecture and knowledgebased structures for smart manufacturing networks. IEEE Software, 32(3):61– 69, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030061-abs.
html.

Pfeffer:2019:SAU

[PWF^{+}19] A. Pfeffer, C. Wu, G. Fry, K. Lu, S. Marotta, M. Reposa, Y. Shi, T. K. S. Kumar, C. A. Knoblock, D. Parker, I. Muhammad, and C. Novakovic. Software adaptation for an unmanned undersea vehicle. IEEE Software, 36(2): 91–96, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pautasso:2015:SYI

[PZ15] Cesare Pautasso and Olaf Zimmermann. Seeking your insights. IEEE Software, 32(2):7–9, March/ April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http://

                                        www.computer.org/csdl/
                                        mags/so/2015/02/mso2015020007.
                                        html.

Pautasso:2018:WSC

[PZ18] Cesare Pautasso and Olaf Zimmermann. The Web as a software connector: Integration resting on linked resources. IEEE Software, 35(1):93–98, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2018/
                                        01/mso2018010093-abs.
                                        html.

Pautasso:2017:MPPa

[PZA^{+}17a] Cesare Pautasso, Olaf Zimmermann, Mike Amundsen, James Lewis, and Nicolai Josuttis. Microservices in practice, part 1: Reality check and service design. IEEE Software, 34(1):91–98, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010091-abs.
html.

Pautasso:2017:MPPb

[PZA^{+}17b] Cesare Pautasso, Olaf Zimmermann, Mike Amundsen, James Lewis, and Nicolai Josuttis. Microservices in practice, part 2: Service integration and sustainability. IEEE Software, 34(2):97– 104, March/April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020097-abs.
html.

Prencipe:2012:GEI

[PZTF12] Giuseppe Prencipe, Cesare Zavattari, Alessandro Tommasi, and John Favaro. Guest Editors’ introduction: Algorithms and today’s practitioner. IEEE Software, 29(1): 61–63, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ronzon:2019:MCN

[RBE19] T. Ronzon, J. Buck, and J. Eckstein. Making companies nimble — from software agility to business agility: A conversation with the authors [insights]. IEEE Software, 36(1):79–85, January/February 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rao:2018:PTD

[RCLR18] Aakarsh Rao, Nadir Carreon, Roman Lysecky, and Jerzy Rozenblit. Probabilistic threat detection for risk management in cyber-physical medical systems. IEEE Software, 35(1):38–43, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2018/
                                        01/mso2018010038.html.

Rigby:2012:CPR

[RCP^{+}12] Peter Rigby, Brendan Cleary, Frederic Painchaud, Margaret-Anne Storey, and Daniel German. Contemporary peer review in action: Lessons from open source development. IEEE Software, 29(6):56– 61, November/December 2012. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Rugaber:2011:MSC

[RDMA11] Spencer Rugaber, Rocky Dunlap, Leo Mark, and Sameer Ansari. Managing software complexity and variability in coupled climate models. IEEE Software, 28(6):43–48, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reinero:2018:HSP

[Rei18] B. Reinero. Harsh Sinha on product management. IEEE Software, 35(2):105– 108, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rodriguez:2010:TTD

[REV10] Javier Portillo Rodriguez, Christof Ebert, and Aurora Vizcaino. Technologies and tools for distributed teams. IEEE Software, 27(5):10–14, September/October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rothman:2013:LLL

[RH13] Johanna Rothman and Shane Hastie. Lessons learned from leading workshops about geographically distributed agile teams. IEEE Software, 30(2):7–10, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rooksby:2012:CFD

[RI12] John Rooksby and Nozomi Ikeya. Collaboration in formative design: Working together at a whiteboard. IEEE Software, 29(1):56–60, January/February 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Richardson:2011:WAA

[Ric11] Eric Richardson. What an agile architect can learn from a hurricane meteorologist. IEEE Software, 28(6):9–12, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Riepula:2011:SSC

[Rie11] Mikko Riepula. Sharing source code with clients: a hybrid business and development model. IEEE Software, 28(4):36–41, July/ August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rising:2010:BP

[Ris10a] Linda Rising. The benefit of patterns. IEEE Software, 27(5):15–17, September/October 2010.

CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rising:2010:TOS

[Ris10b] Linda Rising. Telling our stories. IEEE Software, 27(3):6–7, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rising:2010:WPS

[Ris10c] Linda Rising. What programmers should know. IEEE Software, 27(4):15– 17, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Risi:2012:NGA

[Ris12a] Walter Ariel Risi. Nextgeneration architects for a harsh business world. IEEE Software, 29(2):9– 12, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rising:2012:GG

[Ris12b] Linda Rising. The gender gap. IEEE Software, 29 (4):8, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). See [Gla12].

Rising:2012:WCW

[Ris12c]  Linda Rising.  Why

can’t we all play nice?

IEEE Software, 29(5):7– 10, September/October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rising:2014:FWA

[Ris14] Linda Rising. A final word about stories. IEEE Software, 31(2):14–17, March/ April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rabkin:2013:HHC

[RK13] Ariel Rabkin and Randy Howard Katz. How Hadoop clusters break. IEEE Software, 30(4):88–94, July/August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Riungu-Kalliosaari:2012:TCE

[RKTS12] Leah Riungu-Kalliosaari, Ossi Taipale, and Kari Smolander. Testing in the Cloud: Exploring the practice. IEEE Software, 29(2):46–51, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ramasubbu:2015:MTD

[RKW15] Narayan Ramasubbu, Chris F. Kemerer, and C. Jason Woodard. Managing technical debt: Insights from recent empirical evidence. IEEE Software, 32(2):22– 25, March/April 2015.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020022-abs.
html.

Razavian:2016:FEA

[RL16] Maryam Razavian and Patricia Lago. Feminine expertise in architecting teams. IEEE Software, 33(4):64–71, July/ August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040064-abs.
html.

Reid:2012:DIH

[RLBR12] Louise Reid, Marti Lotter, John Burton, and Ita Richardson. Designing and implementing a hospital quality assurance program. IEEE Software, 29 (3):37–44, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Raatikainen:2012:MCS

[RMM^{+}12] Mikko Raatikainen, Tommi Mikkonen, Varvara Myllarniemi, Niko Makitalo, Tomi Mannisto, and Juha Savolainen. Mobile content as a service: a blueprint for a vendorneutral cloud of mobile devices. IEEE Software, 29(4):28–32, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ruiz:2014:IAL

[RNA^{+}14] Israel J. Mojica Ruiz, Meiyappan Nagappan, Bram Adams, Thorsten Berger, Steffen Dienst, and Ahmed E. Hassan. Impact of ad libraries on ratings of Android mobile apps. IEEE Software, 31(6):86–92, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060086-abs.
html.

Ruiz:2016:AAL

[RNA^{+}16a] Israel J. Mojica Ruiz, Meiyappan Nagappan, Bram Adams, Thorsten Berger, Steffen Dienst, and Ahmed E. Hassan. Analyzing ad library updates in Android apps. IEEE Software, 33(2):74– 80, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
02/mso2016020074-abs.
html.

Ruiz:2016:ERS

[RNA^{+}16b] Israel J. Mojica Ruiz, Meiyappan Nagappan, Bram Adams, Thorsten Berger, Steffen Dienst, and Ahmed E. Hassan. Examining the rating system used in mobile-app stores. IEEE Software, 33(6):86–92, November/ December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060086-abs.
html.

Redondo:2015:CEC

[RO15] Jose Manuel Redondo and Francisco Ortin. A comprehensive evaluation of common Python implementations. IEEE Software, 32(4):76–84, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040076-abs.
html.

Ronzon:2016:SRH

[Ron16] Thomas Ronzon. Software retrofit in high-availability systems: When uptime matters. IEEE Software, 33(2):11–17, March/ April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http:

                                        //csdl.computer.org/
                                        csdl/mags/so/2016/02/
                                        mso2016020011.html.

Rousseau:2012:SBH

[Rou12] David Rousseau. The software behind the Higgs boson discovery. IEEE Software, 29(5):11–15, September/October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rozanski:2015:FPS

[Roz15]  Nick Rozanski.  The

five properties of successful architectural oversight. IEEE Software, 32(1):110–112, January/ February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        01/mso2015010110.html.

Rodriguez:2019:SVV

[RPE19] M. Rodriguez, M. Piattini, and C. Ebert. Software verification and validation technologies and tools. IEEE Software, 36(2):13–24, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Reales:2014:MT

[RPFA^{+}14] Pedro Reales, Macario Polo, Jose Luis Fernandez-Aleman, Ambrosio Toval, and Mario Piattini. Mutation testing. IEEE Software, 31(3):30–35, May/ June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ryu:2019:TAB

[RPP19] S. Ryu, J. Park, and J. Park. Toward analysis and bug finding in JavaScript Web applications in the wild. IEEE Software, 36(3):74– 82, May/June 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rahman:2015:RSL

[RR15] Md Tajmilur Rahman and Peter C. Rigby. Release stabilization on Linux and Chrome. IEEE Software, 32(2):81–88, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020081-abs.
html.

Rein:2017:GBB

[RRL^{+}17] Patrick Rein, Stefan Ramson, Jens Lincke, Tim Felgentreff, and Robert Hirschfeld. Group-based behavior adaptation mechanisms in object-oriented systems. IEEE Software, 34(6):78–82, November/ December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        06/mso2017060078-abs.
                                        html.

Robles:2019:TYO

[RSAT19] G. Robles, I. Steinmacher, P. Adams, and C. Treude. Twenty years of open source software: From skepticism to mainstream. IEEE Software, 36(6):12– 15, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rudorfer:2012:BCF

[RSH12] Arnold Rudorfer, Tobias Stenzel, and Gerold Herold. A business case for feature-oriented requirements engineering. IEEE Software, 29(5):54– 59, September/October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rutkowski:2013:GIS

[RSH13] Anne-Francoise Rutkowski, Carol Saunders, and Les Hatton. The generational impact of software. IEEE Software, 30(3):87– 89, May/June 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rademacher:2018:CDD

[RSS18] F. Rademacher, J. Sorgalla, and S. Sachweh. Challenges of domaindriven microservice design: A model-driven perspective. IEEE Software, 35 (3):36–43, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ridi:2012:DSD

[RTV12] Lorenzo Ridi, Jacopo Torrini, and Enrico Vicario. Developing a scheduler with differencebound matrices and the Floyd–Warshall algorithm. IEEE Software, 29(1): 76–83, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rutkowski:2016:WSN

[Rut16] Anne-Francoise Rutkowski.

Work substitution: A neo-Luddite look at software growth. IEEE Software, 33(3):101–104, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030101-abs.
html.

Robbes:2013:SAE

[RVB13] Romain Robbes, Rene Vidal, and Maria Cecilia Bastarrica. Are software analytics efforts worthwhile for small companies? The case of Amisoft. IEEE Software, 30(5):46– 53, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rutkowski:2017:NFL

[RvGH17] Anne-Francoise Rutkowski, Michiel van Genuchten, and Les Hatton. No free lunch for software after all. IEEE Software, 34(5):13–15, September/ October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        05/mso2017050013-abs.
                                        html.

Ramos:2018:APS

[RVT18] M. Ramos, M. T. Valente, and R. Terra. AngularJS performance: A survey study. IEEE Software, 35(2):72–79, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Robillard:2010:RSS

[RWZ10] Martin Robillard, Robert Walker, and Thomas Zimmermann. Recommendation systems for software engineering. IEEE Software, 27(4):80–86, July/ August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Schaminee:2011:SWR

[SA11] Han Schaminee and Hans Aerts. Short and winding road: Software in car navigation systems. IEEE Software, 28(4):19– 21, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Scacchi:2019:SSE

[SA19] W. Scacchi and T. A.

Alspaugh. Securing software ecosystem architectures: Challenges and opportunities. IEEE Software, 36(3):33–38, May/ June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sadafule:2014:MAD

[Sad14] Rahul D. Sadafule. Mobile app development for the Indian market. IEEE Software, 31(3):17–20, May/ June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Salinas:2017:KMM

[Sal17] Edaena Salinas. Katie Malone on machine learning. IEEE Software, 34(4): 92–96, July/August 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040092.html.

Salinas:2018:TBC

[Sal18]  E. Salinas.  Tammy

Bütow on chaos engineering. IEEE Software, 35 (5):125–128, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Salinas:2019:PHF

[Sal19]  E. Salinas.  Pat Hel-

land on failure and resilience in distributed systems. IEEE Software, 36 (2):140–143, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Strandberg:2017:ASL

[SAO^{+}17] Per Erik Strandberg, Wasif Afzal, Thomas J. Ostrand, Elaine J. Weyuker, and Daniel Sundmark. Automated system-level regression test prioritization in a nutshell. IEEE Software, 34(4):30–37, July/ August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040030-abs.
html.

Spinellis:2014:SDT

[SAT14] Diomidis Spinellis and Stephanos Androutsellis-Theotokis. Software development tooling: Information, opinion, guidelines, and tools. IEEE Software, 31(6):21–23, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060021.html.

Srba:2016:WSO

[SB16] Ivan Srba and Maria Bielikova. Why is Stack Overflow failing? Preserving sustainability in community question answering. IEEE Software, 33(4): 80–89, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040080-abs.
html.

Snyder:2018:UAG

[SC18] Barry Snyder and Bill Curtis. Using analytics to guide improvement during an Agile–DevOps transformation. IEEE Software, 35(1):78–83, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010078-abs.
html.

Scacchi:2017:PTC

[Sca17] Walt Scacchi. Practices and technologies in computer game software engineering. IEEE Software, 34(1):110–116, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010110-abs.
html.

Scanlon:2019:CFO

[Sca19] T. P. Scanlon. Critical factors for open source advancement in the U.S. Department of Defense. IEEE Software, 36(6):29– 33, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Souza:2015:RRP

[SCB15] Rodrigo Souza, Christina Chavez, and Roberto A. Bittencourt. Rapid releases and patch backouts: A software analytics approach. IEEE Software, 32(2):89–96, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020089-abs.
html.

Serrano:2011:DRB

[SCC11] Nicolas Serrano and Joseba Campos-Capelastegui. Don’t read books. IEEE Software, 28(2):92–94, March/ April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shull:2016:FSE

[SCC^{+}16] Forrest Shull, Anita Carleton, Jeromy Carriere, Rafael Prikladnicki, and Dongmei Zhang. The future of software engineering. IEEE Software, 33(1):32–35, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010032.html.

Schieferdecker:2012:MBT

[Sch12] Ina Schieferdecker. Modelbased testing. IEEE Software, 29(1):14–18, January/February 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Schermann:2018:CEC

[SCL18] G. Schermann, J. Cito, and P. Leitner. Continuous experimentation: Challenges, implementation techniques, and current research. IEEE Software, 35(2):26–31, March/ April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Scriber:2018:FDB

[Scr18]  B. A. Scriber.  A

framework for determining blockchain applicability. IEEE Software, 35(4): 70–77, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Smite:2015:CSG

[SCW15] Darja Smite, Fabio Calefato, and Claes Wohlin. Cost savings in global software engineering: Where’s the evidence? IEEE Software, 32(4):26–32, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2015/
                                        04/mso2015040026-abs.
                                        html.

Siqueira:2018:CDB

[SCW^{+}18] R. Siqueira, D. Camarinha, M. Wen, P. Meirelles, and F. Kon. Continuous delivery: Building trust in a large-scale, complex government organization. IEEE Software, 35(2):38–43, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Schmid:2013:PLE

[SdA13] Klaus Schmid and Eduardo Santana de Almeida. Product line engineering. IEEE Software, 30(4):24– 30, July/August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Santos:2018:SRS

[Sdd^{+}18] B. M. Santos, I. G. de Guzmán, V. V. de Camargo, M. Piattini, and C. Ebert. Software refactoring for system modernization. IEEE Software, 35(6):62–67, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2017:VEL

[SDSP17] Forrest Shull, Tore Dyba, Helen Sharp, and Rafael Prikladnicki. Voice of evidence: A look back. IEEE Software, 34(4):23– 25, July/August 2017.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040023-abs.
html.

Seetharam:2012:DSM

[See12] Balachandran Seetharam. Developing software for a mobile service in India. IEEE Software, 29(4): 34–39, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Staples:2019:SAB

[SEK^{+}19] J. Staples, C. Endicott, L. Krause, P. Pal, P. Samouelian, R. Schantz, and A. Wellman. A semiautonomic bytecode repair framework. IEEE Software, 36(2):97–102, March/April 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Salvaneschi:2014:PIF

[SEM14] Guido Salvaneschi, Patrick Eugster, and Mira Mezini. Programming with implicit flows. IEEE Software, 31(5):52–59, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/

05/mso2014050052-abs.
html.

Stol:2015:ISA

[SF15] Klaas-Jan Stol and Brian Fitzgerald. Inner source — adopting open source development practices in organizations: A tutorial. IEEE Software, 32(4): 60–67, July/August 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040060-abs.
html.

Serrano:2015:ISC

[SGH15] Nicolas Serrano, Gorka Gallardo, and Josune Hernantes. Infrastructure as a service and cloud technologies. IEEE Software, 32(2):30–36, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020030.html.

Soares:2010:MPR

[SGSM10] Gustavo Soares, Rohit Gheyi, Dalton Serey, and Tiago Massoni. Making program refactoring safer. IEEE Software, 27(4):52–57, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shaw:2012:RDS

[Sha12] Mary Shaw. The role of design spaces. IEEE Software, 29(1):46–50, January/February 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shepperd:2015:HDK

[She15] Martin Shepperd. How do I know whether to trust a research result? IEEE Software, 32(1):106–109, January/February 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        01/mso2015010106-abs.
                                        html.

Serrano:2013:MWA

[SHG13] Nicolas Serrano, Josune Hernantes, and Gorka Gallardo. Mobile Web apps. IEEE Software, 30(5):22– 27, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Serrano:2014:SOA

[SHG14] Nicolas Serrano, Josune Hernantes, and Gorka Gallardo. Service-oriented architecture and legacy systems. IEEE Software, 31(5):15–19, September/ October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050015-abs.
html.

Song:2012:LDN

[SHR12] Xiping Song, Beatrice Hwong, and Johannes Ros. Lessons from developing nonfunctional requirements for a software platform. IEEE Software, 29(2):74–80, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2011:AFL

[Shu11a] Forrest Shull. Assuring the future? A look at validating climate model software. IEEE Software, 28(6):4–8, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shull:2011:MMH

[Shu11b] Forrest Shull. Managing Montezuma: Handling all the usual challenges of software development, and making it fun: An interview with Ed Beach. IEEE Software, 28(5):4–7, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shull:2011:PWF

[Shu11c] Forrest Shull. Perfectionists in a world of finite resources. IEEE Software, 28(2):4–6, March/ April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shull:2011:PWT

[Shu11d] Forrest Shull. Protection from wishful thinking. IEEE Software, 28(4):3–6, July/August 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shull:2011:RA

[Shu11e] Forrest Shull. The road ahead. IEEE Software, 28 (1):2–4, January/February 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shull:2011:WHJ

[Shu11f]  Forrest Shull.  Watts

Humphrey: 4 July 1927– 28 October 2010. IEEE Software, 28(1):5, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Shull:2012:BNW

[Shu12a] Forrest Shull. A brave new world of testing? An interview with Google’s James Whittaker. IEEE Software, 29(2):4–7, March/ April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2012:DWY

[Shu12b] Forrest Shull. Designing a world at your fingertips: a look at mobile user interfaces. IEEE Software, 29(4):4–7, July/ August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2012:DPP

[Shu12c] Forrest Shull. Disbanding the “process police”: New visions for assuring compliance. IEEE Software, 29(3):3–6, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2012:B

[Shu12d] Forrest Shull. I believe!

IEEE Software, 29(1):4– 7, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2012:LF

[Shu12e]  Forrest Shull.  Look-

ing forward. IEEE Software, 29(5):2–5, September/October 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2012:R

[Shu12f] Forrest Shull. Research 2.0? IEEE Software, 29(6):4–8, November/December 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2013:EVA

[Shu13a] Forrest Shull. Engineering values: From architecture games to agile requirements. IEEE Software, 30(2):2–6, March/ April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2013:ELG

[Shu13b] Forrest Shull. From the Editor: A lifetime guarantee. IEEE Software, 30(6): 4–8, November/December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2013:GIB

[Shu13c] Forrest Shull. Getting an intuition for big data. IEEE Software, 30(4):3–6, July/August 2013. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2013:HES

[Shu13d] Forrest Shull. The human element in social networking. IEEE Software, 30 (1):2–6, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2013:OCC

[Shu13e]  Forrest Shull.  The

only constant is change. IEEE Software, 30(5):4–9, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2013:SYS

[Shu13f] Forrest Shull. Sharing your story. IEEE Software, 30 (3):4–7, May/June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2014:CT

[Shu14a] Forrest Shull. The clock is ticking. IEEE Software, 31(3):4–8, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2014:DDE

[Shu14b] Forrest Shull. Data, data everywhere . . .. IEEE Software, 31(5):4–7, September/October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050004.html.

Shull:2014:EPR

[Shu14c] Forrest Shull. From the Editor: Progression, regression, or stasis? IEEE Software, 31(1):4–8, January/February 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2014:OBH

[Shu14d] Forrest Shull. Our best hope. IEEE Software, 31(4):4–8, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2014:PB

[Shu14e] Forrest Shull. Passing the baton. IEEE Software, 31(6):4–8, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060004.html.

Shull:2014:TCM

[Shu14f]  Forrest Shull.  The

true cost of mobility? IEEE Software, 31(2):5–9, March/April 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sinha:2013:ERE

[SJ13] Atish P. Sinha and Hemant Jain. Ease of reuse: An empirical comparison of components and objects. IEEE Software, 30(5):70–75, September/ October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Seetharam:2015:MMI

[SJ15] Balachandran Seetharam and Drew Johnson. Mobile money’s impact on Tanzanian agriculture. IEEE Software, 32(1):29–34, January/February 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010029-abs.
html.

Sagesser:2013:IIL

[SJG13] Katharina Sagesser, Bonney Joseph, and Rainer Grau. Introducing an iterative life-cycle model at Credit Suisse IT Switzerland. IEEE Software, 30(2):68–73, March/April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sjoberg:2012:QEU

[SJS12] Dag I. K. Sjoberg, Anders Johnsen, and Jorgen Solberg. Quantifying the effect of using Kanban versus Scrum: a case study. IEEE Software, 29(5):47–53, September/ October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Smite:2014:VTG

[SKK14] Darja Smite, Marco Kuhrmann, and Patrick Keil. Virtual teams [Guest Editors’ introduction]. IEEE Software, 31(6):41–46, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060041.html.

Stol:2017:CSE

[SLB17] Klaas-Jan Stol, Thomas D. LaToza, and Christian Bird. Crowdsourcing for software engineering. IEEE Software, 34(2):30– 36, March/April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
02/mso2017020030.html.

Savolainen:2010:CCS

[SM10a] Juha Savolainen and Tomi Mannisto. Conflictcentric software architectural views: Exposing trade-offs in quality requirements. IEEE Software, 27(6):33–37, November/December 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sutherland:2010:SR

[SM10b] Malcolm Sutherland and Neil Maiden. Storyboarding requirements. IEEE Software, 27(6):9– 11, November/December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Smart:2018:TAD

[Sma18]  J. Smart.  To trans-

form to have agility, don’t do a capital A, capital T agile transformation. IEEE Software, 35(6):56– 60, November/December 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sun:2016:EPP

[SMAU16] Wenying Sun, George Marakas, and Miguel Aguirre-Urreta. The effectiveness of pair programming: Software professionals’ perceptions. IEEE Software, 33(4):72– 79, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/

04/mso2016040072-abs.
html.

Smite:2019:SGH

[SMLF19] D. Smite, N. B. Moe, G. Levinta, and M. Floryan. Spotify guilds: How to succeed with knowledge sharing in large-scale agile organizations. IEEE Software, 36(2):51–57, March/ April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Schmitz:2019:UEE

[SMN19] K. Schmitz, R. Mahapatra, and S. Nerur. User engagement in the era of hybrid agile methodology. IEEE Software, 36(4): 32–40, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Silva:2015:SPM

[SMP15] Pedro Silva, Ana M. Moreno, and Lawrence Peters. Software project management: Learning from our mistakes [voice of evidence]. IEEE Software, 32(3):40–43, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030040-abs.
html.

Sierszecki:2014:GSG

[SMS^{+}14] Krzysztof Sierszecki, Tommi Mikkonen, Michaela Steffens, Thomas Fogdal, and Juha Savolainen. Green software: Greening what and how much? IEEE Software, 31(3):64–68, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shull:2010:WDW

[SMT^{+}10] Forrest Shull, Grigori Melnik, Burak Turhan, Lucas Layman, Madeline Diep, and Hakan Erdogmus. What do we know about test-driven development? IEEE Software, 27(6):16–19, November/ December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Savolainen:2013:LTP

[SNMF13] Juha Savolainen, Nan Niu, Tommi Mikkonen, and Thomas Fogdal. Longterm product line sustainability with planned staged investments. IEEE Software, 30(6):63–69, November/December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sommerville:2016:ISP

[Som16] Ian Sommerville. IEEE Software and professional development. IEEE Software, 33(2):90–92, March/ April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http:

//csdl.computer.org/
csdl/mags/so/2016/02/
mso2016020090.html.

Sandberg:2011:ACR

[SPA11] Anna Sandberg, Lars Pareto, and Thomas Arts. Agile collaborative research: Action principles for industry-academia collaboration. IEEE Software, 28(4):74–83, July/ August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Szyperski:2016:TEB

[SPB16] Clemens Szyperski, Martin Petitclerc, and Roger Barga. Three experts on big data engineering. IEEE Software, 33(2):68– 72, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
02/mso2016020068.html.

Spinellis:2010:CD

[Spi10a] Diomidis Spinellis. Code documentation. IEEE Software, 27(4):18–19, July/August 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2010:FD

[Spi10b] Diomidis Spinellis. Farewell to disks. IEEE Software, 27(6):82–83, November/ December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2010:ST

[Spi10c] Diomidis Spinellis. Software tracks. IEEE Software, 27(2):10–11, March/ April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2010:UE

[Spi10d] Diomidis Spinellis. UML everywhere. IEEE Software, 27(5):90–91, September/October 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2011:AD

[Spi11a] Diomidis Spinellis. Agility drivers. IEEE Software, 28(4):96, 95, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2011:CUO

[Spi11b] Diomidis Spinellis. Choosing and using open source components. IEEE Software, 28(3):96, 95, May/ June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2011:EE

[Spi11c] Diomidis Spinellis. elytS edoC. IEEE Software, 28 (2):104, 103, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2011:FI

[Spi11d] Diomidis Spinellis. Faking it. IEEE Software, 28(5):96, 95, September/ October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2012:ALC

[Spi12a] Diomidis Spinellis. APIs, libraries, and code. IEEE Software, 29(6):85–86, November/December 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2012:DIS

[Spi12b] Diomidis Spinellis. Don’t install software by hand. IEEE Software, 29(4):86– 87, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2012:G

[Spi12c] Diomidis Spinellis. Git. IEEE Software, 29(3):100– 101, May/June 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2012:PMS

[Spi12d] Diomidis Spinellis. Package management systems. IEEE Software, 29(2):84– 86, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2012:RC

[Spi12e] Diomidis Spinellis. Refactoring on the cheap. IEEE Software, 29(1):96, 94–95, January/February 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2012:VM

[Spi12f] Diomidis Spinellis. Virtualize me. IEEE Software, 29(5):91–93, September/ October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2013:DD

[Spi13a] Diomidis Spinellis. Differential debugging. IEEE Software, 30(5):19–21, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2013:FDE

[Spi13b] Diomidis Spinellis. The frictionless development environment scorecard. IEEE Software, 30(6):30– 32, November/December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2013:IBD

[Spi13c] Diomidis Spinellis. The importance of being declarative. IEEE Software, 30(1):90–91, January/ February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2013:PGV

[Spi13d] Diomidis Spinellis. Portability: Goodies vs. the hair shirt. IEEE Software, 30(4):22–23, July/ August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2013:SS

[Spi13e] Diomidis Spinellis. Systems software. IEEE Software, 30(3):18–19, May/ June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2014:BI

[Spi14a] Diomidis Spinellis. Bespoke infrastructures. IEEE Software, 31(1):23–25, January/February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2014:DC

[Spi14b] Diomidis Spinellis. Developing in the cloud. IEEE Software, 31(2):41– 43, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2014:FDN

[Spi14c] Diomidis Spinellis. First, do no harm. IEEE Software, 31(5):12–14, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050012.html.

Spinellis:2014:SOR

[Spi14d] Diomidis Spinellis. Service orchestration with Rundeck. IEEE Software, 31(4):16–18, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2015:ADP

[Spi15a] Diomidis Spinellis. Architecture from a developer’s perspective. IEEE Software, 32(5):4–7, September/October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050004.html.

Spinellis:2015:EOF

[Spi15b] Diomidis Spinellis. Extending our field’s reach. IEEE Software, 32(6):4– 6, November/December 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060004.html.

Spinellis:2015:FDT

[Spi15c] Diomidis Spinellis. Fewer dead trees, more engagement. IEEE Software, 32(1):4–6, January/ February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010004.html.

Spinellis:2015:I

[Spi15d] Diomidis Spinellis. Introductions. IEEE Software, 32(4):3–5, July/ August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040003.html.

Spinellis:2015:RST

[Spi15e] Diomidis Spinellis. Recruiting a star team. IEEE Software, 32(3):3– 5, May/June 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030003.html.

Spinellis:2015:SIR

[Spi15f] Diomidis Spinellis. The strategic importance of release engineering. IEEE Software, 32(2):3–5, March/ April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http://

www.computer.org/csdl/
mags/so/2015/02/mso2015020003.
html.

Spinellis:2016:BDD

[Spi16a] Diomidis Spinellis. Being a DevOps developer. IEEE Software, 33(3):4– 5, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030004.html.

Spinellis:2016:CRS

[Spi16b] Diomidis Spinellis. The changing role of the software architect. IEEE Software, 33(6):4–6, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060004.html.

Spinellis:2016:DDT

[Spi16c] Diomidis Spinellis. Developer, debug thyself. IEEE Software, 33(1):3–5, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010003.html.

Spinellis:2016:MSB

[Spi16d] Diomidis Spinellis. Managing a software business. IEEE Software, 33(5):4–7, September/October 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050004.html.

Spinellis:2016:RQ

[Spi16e] Diomidis Spinellis. Reflecting on quality. IEEE Software, 33(4):4–5, July/ August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040004.html.

Spinellis:2016:SP

[Spi16f] Diomidis Spinellis. Serving professionals. IEEE Software, 33(2):4–6, March/ April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http:

//csdl.computer.org/
csdl/mags/so/2016/02/
mso2016020004.html.

Spinellis:2017:ESH

[Spi17a] Diomidis Spinellis. The elusiveness of smart healthcare. IEEE Software, 34 (6):4–6, November/December 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

URL https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060004.html.

Spinellis:2017:HAC

[Spi17b] Diomidis Spinellis. How abundance changes software engineering. IEEE Software, 34(3):4–7, May/ June 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL https://

www.computer.org/csdl/
mags/so/2017/03/mso2017030004.
html.

Spinellis:2017:SRS

[Spi17c] Diomidis Spinellis. The social responsibility of software development. IEEE Software, 34(2):4–6, March/ April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL https://

www.computer.org/csdl/
mags/so/2017/02/mso2017020004.
html.

Spinellis:2017:SEI

[Spi17d] Diomidis Spinellis. Softwareengineering the Internet of Things. IEEE Software, 34(1):4–6, January/ February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010004.html.

Spinellis:2017:SRR

[Spi17e] Diomidis Spinellis. Software reliability redux. IEEE Software, 34(4):4–7, July/August 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
04/mso2017040004.html.

Spinellis:2017:SAS

[Spi17f] Diomidis Spinellis. Stateof-the-art software testing. IEEE Software, 34(5):4–6, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050004.html.

Spinellis:2018:BSD

[Spi18a] D. Spinellis. Being a software developer. IEEE Software, 35(4):4–7, July/ August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2018:CPR

[Spi18b] D. Spinellis. The challenges and practices of release engineering. IEEE Software, 35(2):4–7, March/ April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2018:ER

[Spi18c] D. Spinellis. The Editor’s retrospective. IEEE Software, 35(6):4–7, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2018:SES

[Spi18d] D. Spinellis. Self-evolving software architectures. IEEE Software, 35(3):4–7, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2018:SHM

[Spi18e] D. Spinellis. The success of a heavenly marriage. IEEE Software, 35(5):3–6, September/October 2018.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2018:UCI

[Spi18f] Diomidis Spinellis. Under the covers of IEEE Software. IEEE Software, 35(1):4–7, January/ February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010004.html.

Schulenklopper:2016:WTJ

[SR16] Jochem Schulenklopper and Eelco Rommes. Why they just don’t get it: Communicating about architecture with business stakeholders. IEEE Software, 33(3):13–19, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030013-abs.
html.

Skyrme:2014:SMC

[SRI14] Alexandre Skyrme, Noemi Rodriguez, and Roberto Ierusalimschy. Scripting multiple CPUs with safe data sharing. IEEE Software, 31(5):44–51, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050044-abs.
html.

Spinellis:2011:LS

[SS11] Diomidis Spinellis and Henry Spencer. Lessons from space. IEEE Software, 28(6):26–28, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Srinivasan:2017:WAS

[SS17] Satish M. Srinivasan and Raghvinder S. Sangwan. Web app security: A comparison and categorization of testing frameworks. IEEE Software, 34(1):99–102, January/ February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010099-abs.
html.

Sun:2018:PAM

[SS18] W. Sun and C. Schmidt.

Practitioners agile-methodology use and job perceptions. IEEE Software, 35(2):52– 61, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shahbaz:2012:SIU

[SSE12] Muzammil Shahbaz, K. C. Shashidhar, and Robert Eschbach. Specification inference using systematic reverse-engineering methodologies: An automotive industry application. IEEE Software, 29(6):62–69, November/ December 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sharma:2015:CSR

[SSS15a] Tushar Sharma, Girish Suryanarayana, and Ganesh Samarthyam. Challenges to and solutions for refactoring adoption: An industrial perspective. IEEE Software, 32(6):44–51, November/ December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060044-abs.
html.

Suryanarayana:2015:SPV

[SSS15b]  Girish  Suryanarayana,

Tushar Sharma, and Ganesh Samarthyam. Software process versus design quality: Tug of war? IEEE Software, 32(4):7– 11, July/August 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040007.html.

Sharif:2016:TDE

[SSWM16] Bonita Sharif, Timothy Shaffer, Jenna Wise, and Jonathan I. Maletic. Tracking developers’ eyes in the IDE. IEEE Software, 33(3):105–108, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030105-abs.
html.

Serrano:2010:WP

[ST10] Nicolas Serrano and Jose Manuel Torres. Web 2.0 for practitioners. IEEE Software, 27 (3):11–15, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Steinmacher:2019:LMG

[STG19] I. Steinmacher, C. Treude, and M. A. Gerosa. Let me in: Guidelines for the successful onboarding of newcomers to open source projects. IEEE Software, 36(4):41–49, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sturtevant:2018:MAM

[Stu18] Dan Sturtevant. Modular architectures make you agile in the long run. IEEE Software, 35(1):104–108, January/February 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2018/
                                        01/mso2018010104-abs.
                                        html.

Suneja:2019:SOD

[Sun19] N. Suneja. ScyllaDB optimizes database architecture to maximize hardware performance. IEEE Software, 36(4):96–100, July/ August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sneed:2016:SDS

[SV16] Harry M. Sneed and Chris Verhoef. From software development to software assembly. IEEE Software, 33(5):80–85, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        05/mso2016050080-abs.
                                        html.

Schuwer:2015:IBO

[SvGH15] Robert Schuwer, Michiel van Genuchten, and Les Hatton. On the impact of being open. IEEE Software, 32(5):81–83, September/October 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050081.html.

Smite:2016:WTH

[SvS16] Darja Smite and Rini van Solingen. What’s the true hourly cost of offshoring? IEEE Software, 33(5):60–70, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050060-abs.
html.

Smite:2011:SFS

[SW11a] Darja Smite and Claes Wohlin. Strategies facilitating software product transfers. IEEE Software, 28(5):60–66, September/ October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Smite:2011:WEG

[SW11b] Darja Smite and Claes Wohlin. A whisper of evidence in global software engineering. IEEE Software, 28(4):15–18, July/ August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stopford:2017:TDC

[SWA17] Ben Stopford, Ken Wallace, and John Allspaw. Technical debt: Challenges and perspectives. IEEE Software, 34(4):79– 81, July/August 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        04/mso2017040079-abs.
                                        html.

Solanas:2017:RAH

[SWB^{+}17] Agusti Solanas, Jens H. Weber, Ayse Basar Bener, Frank van der Linden, and Rafael Capilla. Recent advances in healthcare software: Toward context-aware and smart solutions. IEEE Software, 34(6):36–40, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        06/mso2017060036.html.

Symons:2010:SIP

[Sym10] Charles Symons. Software industry performance: What you measure is what you get. IEEE Software, 27(6):66– 72, November/December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Symons:2012:ESP

[Sym12] Charles Symons. Exploring software project effort versus duration tradeoffs. IEEE Software, 29(4): 67–74, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Temple:2017:LCV

[TAJB17] Paul Temple, Mathieu Acher, Jean-Marc Jezequel, and Olivier Barais. Learning contextual-variability models. IEEE Software, 34(6):64–70, November/ December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060064-abs.
html.

Taylor:2018:R

[Tay18] D. Taylor. 2017 reviewers. IEEE Software, 35(2): 8–10, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Teh:2012:SPS

[TBvRB12] Alvin Teh, Elisa Baniassad, Dirk van Rooy, and Clive Boughton. Social psychology and software teams: Establishing task-effective group norms. IEEE Software, 29(4):53–58, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tarhan:2012:AQM

[TD12] Ayca Tarhan and Onur Demirors. Apply quantitative management now. IEEE Software, 29(3):77– 85, May/June 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tenbergen:2019:RER

[TD19] B. Tenbergen and M. Daun. Is requirements-engineering research delivering what it promised?: A review of its accomplishments and opportunities after 10 years. IEEE Software, 36(4):6– 11, July/August 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Terzakis:2011:VRG

[Ter11] John Terzakis. Virtual retrospectives for geographically dispersed software teams. IEEE Software, 28 (3):12–15, May/June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Turetken:2012:CCR

[TEvdHP12] Oktay Turetken, Amal Elgammal, Willem-Jan van den Heuvel, and Michael P. Papazoglou. Capturing compliance requirements: a patternbased approach. IEEE Software, 29(3):28–36, May/June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Thones:2015:BOL

[Tho15a] Johannes Thones. Barry O’Reilly on lean enterprises. IEEE Software, 32(6):101–104, November/ December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060101.html.

Thones:2015:M

[Tho15b] Johannes Thones. Microservices. IEEE Software, 32(1):116, January/ February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
01/mso2015010116.html.

Tomita:2011:SET

[TI11] Koji Tomita and Keiichi Ito. Software in an evolving train traffic control system. IEEE Software, 28(2):19–21, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tilkov:2015:MCB

[Til15] Stefan Tilkov. The modern cloud-based platform. IEEE Software, 32(2):116, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020116.html.

Tilkov:2016:JFW

[Til16a] Stefan Tilkov. Jay Fields on working with unit tests. IEEE Software, 33(5):117–120, September/October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050117-abs.
html.

Tilkov:2016:VVR

[Til16b] Stefan Tilkov. Vaughn Vernon on reactive programming with the actor model. IEEE Software, 33(3):109–112, May/ June 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http:

//csdl.computer.org/

csdl/mags/so/2016/03/
mso2016030109.html.

Trubiani:2019:PIH

[TJC^{+}19] C. Trubiani, P. Jamshidi, J. Cito, W. Shang, Z. M. Jiang, and M. Borg. Performance issues? Hey DevOps, mind the uncertainty. IEEE Software, 36 (2):110–117, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tamburri:2016:ARC

[TKF16] Damian A. Tamburri, Rick Kazman, and Hamed Fahimi. The architect’s role in community shepherding. IEEE Software, 33(6):70–79, November/ December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060070-abs.
html.

Taibi:2018:DMB

[TL18] D. Taibi and V. Lenarduzzi. On the definition of microservice bad smells. IEEE Software, 35 (3):56–62, May/June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tamburri:2013:ULS

[TLvV13] Damian A. Tamburri, Patricia Lago, and Hans van Vliet. Uncovering latent social communities in software development. IEEE Software, 30(1): 29–36, January/February 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Taivalsaari:2017:RPW

[TM17] Antero Taivalsaari and Tommi Mikkonen. A roadmap to the programmable world: Software challenges in the IoT era. IEEE Software, 34(1):72–80, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010072-abs.
html.

Taivalsaari:2018:TIC

[TM18] A. Taivalsaari and T. Mikkonen. A taxonomy of IoT client architectures. IEEE Software, 35(3):83– 88, May/June 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Traub:2017:FAA

[TMB17] Matthias Traub, Alexander Maier, and Kai L. Barbehon. Future automotive architecture and the impact of IT trends. IEEE Software, 34(3):27– 32, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
03/mso2017030027.html.

Torres:2015:CKW

[Tor15] Francisco Torres. Context is king: What’s your software’s operating range? IEEE Software, 32(5):9–12, September/October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050009-abs.
html.

Toupin:2011:UTD

[Tou11] Dominique Toupin. Using tracing to diagnose or monitor systems. IEEE Software, 28(1):87–91, January/February 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tsuchitoi:2011:MYO

[TS11] Yuki Tsuchitoi and Hideki Sugiura. 10 MLOC in your office copier. IEEE Software, 28(6):93–95, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Taivalsaari:2012:CHC

[TS12] Antero Taivalsaari and Kari Systa. Cloudberry: An HTML5 cloud phone platform for mobile devices. IEEE Software, 29(4):40–45, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tsakiris:2011:MSI

[Tsa11] Anthony Tsakiris. Managing software interfaces of on-board automotive controllers. IEEE Software, 28(1):73–76, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Thurimella:2017:GMR

[TSPB17] Anil Kumar Thurimella, Mathias Schubanz, Andreas Pleuss, and Goetz Botterweck. Guidelines for managing requirements rationales. IEEE Software, 34(1):82–90, January/February 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        01/mso2017010082-abs.
                                        html.

Toledo:2011:ACJ

[TT11] Rodolfo Toledo and Eric Tanter. Access control in JavaScript. IEEE Software, 28(5):76–84, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Telea:2010:VTS

[TVS10] Alexandru Telea, Lucian Voinea, and Hans Sassenburg. Visual tools for software architecture understanding: a stakeholder perspective. IEEE Software, 27(6):46–53, November/December 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Tang:2012:DSS

[TvV12] Antony Tang and Hans van Vliet. Design strategy and software design effectiveness. IEEE Software, 29(1):51–55, January/February 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tratt:2014:IGB

[TW14a] Laurence Tratt and Adam Welc. An interview with Gilad Bracha. IEEE Software, 31(5):76–79, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050076-abs.
html.

Tratt:2014:PL

[TW14b] Laurence Tratt and Adam Welc. Programming languages. IEEE Software, 31(5):33–34, September/ October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2014/
                                        05/mso2014050033.html.

Tang:2016:FCS

[TYXL16] Zhengrong Tang, Melissa Yang, Joshua Xiang, and John Liu. The future of Chinese software development. IEEE Software, 33(1):40–44, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://www.computer.
                                        org/csdl/mags/so/2016/
                                        01/mso2016010040.html.

Unterkalmsteiner:2018:PIA

[UG18] M. Unterkalmsteiner and T. Gorschek. Process improvement archaeology: What led us here, and what’s next? IEEE Software, 35(4):53–61, July/ August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Uittenbogaard:2013:SSP

[Uit13] Arjen Uittenbogaard. Storytelling for software professionals. IEEE Software, 30(3):9–12, May/ June 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Usaola:2010:MTC

[UM10] Macario Polo Usaola and Pedro Reales Mateo. Mutation testing cost reduction techniques: a survey. IEEE Software, 27 (3):80–86, May/June 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Uddin:2015:HAD

[UR15] Gias Uddin and Martin P. Robillard. How API documentation fails. IEEE Software, 32(4):68– 75, July/August 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040068-abs.
html.

Venkatesh:2017:SAD

[VAC^{+}17] Jagannathan Venkatesh, Baris Aksanli, Christine S. Chan, Alper S. Akyurek, and Tajana S. Rosing. Scalable-application design for the IoT. IEEE Software, 34(1):62–70, January/February 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        01/mso2017010062-abs.
                                        html.

Valerdi:2017:WSL

[Val17] Ricardo Valerdi. Why software is like baseball. IEEE Software, 34(5):7–9, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        05/mso2017050007-abs.
                                        html.

Vijayasarathy:2016:CSD

[VB16] Leo R. Vijayasarathy and Charles W. Butler. Choice of software development methodologies: Do organizational, project, and team characteristics matter? IEEE Software, 33(5):86–94, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        05/mso2016050086-abs.
                                        html.

Vianna:2017:PHS

[VBP17] Henrique Damasceno Vianna, Jorge Luis Victoria Barbosa, and Fabio Pittoli. In the pursuit of Hygge software. IEEE Software, 34(6):48–52, November/December 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
06/mso2017060048-abs.
html.

VanGeet:2010:REM

[VD10] Joris Van Geet and Serge Demeyer. Reverse engineering on the mainframe: Lessons learned from “in vivo” research. IEEE Software, 27(4):30–36, July/ August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Verborgh:2015:DCL

[VD15] Ruben Verborgh and Jos De Roo. Drawing conclusions from linked data on the Web: The EYE reasoner. IEEE Software, 32 (3):23–27, May/June 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
03/mso2015030023-abs.
html.

vanderBurgt:2017:ITS

[vdBB17] Therry van der Burgt and Allard Baronner. An information and tracking system for inland shipping. IEEE Software, 34(3):105– 110, May/June 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        03/mso2017030105-abs.
                                        html.

Vilkki:2012:PCV

[VE12] Kati Vilkki and Hakan Erdogmus. Point/counterpoint: The value of lean vs. lean is a fad. IEEE Software, 29(5):60–63, September/ October 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Villela:2019:URE

[VGD19] K. Villela, E. C. Groen, and J. Doerr. Ubiquitous requirements engineering: A paradigm shift that affects everyone. IEEE Software, 36(2):8–12, March/ April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

vanGenuchten:2010:SWI

[vGH10] Michiel van Genuchten and Les Hatton. Software: What’s in it and what’s it in? IEEE Software, 27(1): 14–16, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

vanGenuchten:2011:SM

[vGH11] Michiel van Genuchten and Les Hatton. Software mileage. IEEE Software, 28(5):24–26, September/ October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

vanGenuchten:2012:CAG

[vGH12] Michiel van Genuchten and Les Hatton. Compound annual growth rate for software. IEEE Software, 29(4):19–21, July/ August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

vanGenuchten:2013:MI

[vGH13] Michiel van Genuchten and Les Hatton. Metrics with impact. IEEE Software, 30(4):99–101, July/ August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

vanGenuchten:2015:MSI

[vGH15] Michiel van Genuchten and Les Hatton. Managing software’s impact. IEEE Software, 32(4):15– 17, July/August 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040015-abs.
html.

vanGenuchten:2019:TYI

[vGH19] M. van Genuchten and L. Hatton. Ten years of “impact” columns: The good, the bad, and the ugly. IEEE Software, 36(6):57–60, November/ December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

vanGenuchten:2014:YUW

[vGMRW14] Michiel van Genuchten, Ronny Mans, Hajo Reijers, and Daniel Wismeijer. Is your upgrade worth it? Process mining can tell. IEEE Software, 31(5):94–100, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050094-abs.
html.

vanHeesch:2014:DCA

[vHEA^{+}14] Uwe van Heesch, Veli-Pekka Eloranta, Paris Avgeriou, Kai Koskimies, and Neil Harrison. Decisioncentric architecture reviews. IEEE Software, 31(1):69–76, January/ February 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Vial:2015:DRL

[Via15] Gregory Vial. Database refactoring: Lessons from the trenches. IEEE Software, 32(6):71–79, November/December 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2015/
06/mso2015060071-abs.
html.

Vial:2018:DDD

[Via18] G. Vial. Different databases for different strokes. IEEE Software, 35(2):80–85, March/April 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Vial:2019:LPO

[Via19] G. Vial. Lessons in persisting object data using object-relational mapping. IEEE Software, 36(6):43– 52, November/December 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Verbrugge:2011:GEI

[VK11] Clark Verbrugge and Paul Kruszewski. Guest Editors’ introduction: Engineering fun. IEEE Software, 28(5):27–29, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Vandierendonck:2012:TTP

[VM12] Hans Vandierendonck and Tom Mens. Techniques and tools for parallelizing software. IEEE Software, 29(2):22–25, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

vanMalkenhorst:2012:GU

[vMM12] Michiel van Malkenhorst and Lex Mollinger. Going underground. IEEE Software, 29(3):17–20, May/ June 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Voas:2012:MSA

[VMvG12] Jeffrey Voas, J. Bret Michael, and Michiel van Genuchten. The mobile software app takeover. IEEE Software, 29(4):25– 27, July/August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Vorraber:2019:AMC

[VMVS19] W. Vorraber, M. Mueller, S. Voessner, and W. Slany. Analyzing and managing complex software ecosystems: A framework to understand value in information systems. IEEE Software, 36(3):55–60, May/ June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Varadharajan:2018:AUR

[VOGB18] V. S. Varadharajan, D. S.

Onge, C. Guß, and G. Beltrame. Over-the-air updates for robotic swarms. IEEE Software, 35(2):44– 50, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Volter:2010:AL

[Vol10]  Markus Volter.  Ar-

chitecture as language. IEEE Software, 27(2):56– 64, March/April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Volter:2011:PMB

[Vol11]  Markus Volter.  From

programming to modeling — and back again. IEEE Software, 28(6):20– 25, November/December 2011. ISSN 0740-7459 (print), 0740-7459 (electronic).

Vierhauser:2016:MRS

[VRG16] Michael Vierhauser, Rick Rabiser, and Paul Grunbacher. Monitoring requirements in systems of systems. IEEE Software, 33(5):22–24, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050022-abs.
html.

Vallon:2016:AFA

[VSB^{+}16] Raoul Vallon, Stefan Strobl, Mario Bernhart, Rafael Prikladnicki, and Thomas Grechenig. ADAPT: A framework for agile distributed software development. IEEE Software, 33(6):106–111, November/ December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060106-abs.
html.

vonWangenheim:2010:CSP

[vWHZ^{+}10] Christiane Gresse von Wangenheim, Jean Carlo R. Hauck, Alessandra Zoucas, Clenio F. Salviano, Fergal McCaffery, and Forrest Shull. Creating software process capability/maturity models. IEEE Software, 27(4):92– 94, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Voelter:2015:PMF

[VWK15] Markus Voelter, Jos Warmer, and Bernd Kolb. Projecting a modular future. IEEE Software, 32(5):46–52, September/ October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
05/mso2015050046-abs.
html.

Wohlin:2012:SFP

[WAA^{+}12] Claes Wohlin, Aybuke Aurum, Lefteris Angelis, Laura Phillips, Yvonne Dittrich, Tony Gorschek, Hakan Grahn, Kennet Henningsson, Simon Kagstrom, Graham Low, Per Rovegard, Piotr Tomaszewski, Christine van Toorn, and Jeff Winter. The success factors powering industry– academia collaboration. IEEE Software, 29(2):67– 73, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Waterman:2018:ARUa

[Wat18a] M. Waterman. Agility, risk, and uncertainty, part 1: Designing an agile architecture. IEEE Software, 35(2):99–101, March/April 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Waterman:2018:ARUb

[Wat18b] M. Waterman. Agility, risk, and uncertainty, part 2: How risk impacts agile architecture. IEEE Software, 35(3):18–19, May/ June 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wampler:2010:GEI

[WC10] Dean Wampler and Tony Clark. Guest Editors’ introduction: Multiparadigm programming. IEEE Software, 27(5):20– 24, September/October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wampler:2010:MPI

[WCFG10] Dean Wampler, Tony Clark, Neal Ford, and Brian Goetz. Multiparadigm programming in industry: a discussion with Neal Ford and Brian Goetz. IEEE Software, 27(5):61–64, September/ October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Whalen:2016:RAS

[WCG16] Michael W. Whalen, Darren Cofer, and Andrew Gacek. Requirements and architectures for secure vehicles. IEEE Software, 33(4):22–25, July/ August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040022-abs.
html.

Weyrich:2016:RAI

[WE16] Michael Weyrich and Christof Ebert. Reference architectures for the Internet of Things. IEEE Software, 33(1):112–116, January/February 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010112.html.

Woods:2010:PC

[WES10] Eoin Woods, David Emery, and Bran Selic. Point/ counterpoint. IEEE Software, 27(6):54–57, November/December 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Woods:2018:PAE

[WF18] E. Woods and G. Fairbanks. The pragmatic architect evolves. IEEE Software, 35(6):12–15, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Weyns:2010:SAU

[WG10] Danny Weyns and Michael Georgeff. Self-adaptation using multiagent systems. IEEE Software, 27(1): 86–91, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Weinreich:2016:ARP

[WG16] Rainer Weinreich and Iris Groher. The architect’s role in practice: From decision maker to knowledge manager? IEEE Software, 33(6):63–69, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060063-abs.
html.

Whalen:2013:YWM

[WGC^{+}13] Michael W. Whalen, Andrew Gacek, Darren Cofer, Anitha Murugesan, Mats P. E. Heimdahl, and Sanjai Rayadurgam. Your “what” is my “how”: Iteration and hierarchy in system design. IEEE Software, 30(2):54–60, March/ April 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Whittaker:2012:MTP

[Whi12] James A. Whittaker. The 10-minute test plan. IEEE Software, 29(6):70– 77, November/December 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Whittle:2019:YSV

[Whi19] J. Whittle. Is your software valueless? IEEE Software, 36(3):112–115, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wloka:2010:TSC

[WHR10] Jan Wloka, Einar Hoest, and Barbara G. Ryder. Tool support for changecentric test development. IEEE Software, 27(3):66– 71, May/June 2010. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Whittle:2014:SPM

[WHR14] Jon Whittle, John Hutchinson, and Mark Rouncefield. The state of practice in model-driven engineering. IEEE Software, 31 (3):79–85, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wolff:2015:TD

[WJ15] Eberhard Wolff and Sven Johann. Technical debt. IEEE Software, 32(4):94– c3, July/August 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040094.html.

Wallnau:2011:PC

[WK11] Kurt Wallnau and Philippe Kruchten. Point/counterpoint. IEEE Software, 28(3):56– 59, May/June 2011. CO-DEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wester:2015:SBM

[WK15] Rogier Wester and John Koster. The software behind Moore’s Law. IEEE Software, 32(2):37–40, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020037-abs.
html.

Wang:2014:MOS

[WKS^{+}14] Xiaofeng Wang, Ilona Kuzmickaja, Klaas-Jan Stol, Pekka Abrahamsson, and Brian Fitzgerald. Microblogging in open source software development: The case of Drupal and Twitter. IEEE Software, 31(4):72–80, July/ August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wachsmuth:2014:LDS

[WKV14] Guido H. Wachsmuth, Gabriël D. P. Konat, and Eelco Visser. Language design with the Spoofax language workbench. IEEE Software, 31(5):35–43, September/ October 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
05/mso2014050035-abs.
html.

Wesenberg:2012:MAM

[WL12] Harald Wesenberg and Einar Landre. Making architecture matter. IEEE Software, 29(3):21– 23, May/June 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Werner:2019:CML

[WLD19] C. Werner, Z. S. Li, and D. Damian. Can a machine learn through customer sentiment?: A costaware approach to predict support ticket escalations. IEEE Software, 36(5):38–45, September/ October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Williams:2018:ESV

[WMM18] L. Williams, G. McGraw, and S. Migues. Engineering security vulnerability prevention, detection, and response. IEEE Software, 35(5):76–80, September/ October 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Woods:2014:HUA

[Woo14a] Eoin Woods. Harnessing UML for architectural description — the context view. IEEE Software, 31(6):30–33, November/December 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2014/
06/mso2014060030.html.

Woods:2014:RPA

[Woo14b] Eoin Woods. Return of the pragmatic architect. IEEE Software, 31(3):10– 13, May/June 2014. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Woods:2015:AAW

[Woo15a] Eoin Woods. Aligning architecture work with agile teams. IEEE Software, 32(5):24–26, September/ October 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.

org/csdl/mags/so/2015/
05/mso2015050024.html.

Woods:2015:AGM

[Woo15b] Eoin Woods. Architecting in the gaps: A metaphor for architecture work. IEEE Software, 32(4):33– 35, July/August 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
04/mso2015040033.html.

Woods:2016:HPA

[Woo16a] Eoin Woods. Harnessing the power of architectural design principles. IEEE Software, 33(4):15– 17, July/August 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040015-abs.
html.

Woods:2016:OFA

[Woo16b]  Eoin Woods.  Opera-

tional: The forgotten architectural view. IEEE Software, 33(3):20–23, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030020-abs.
html.

Woods:2016:SAC

[Woo16c] Eoin Woods. Software architecture in a changing world. IEEE Software, 33(6):94–97, November/December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2016/
                                        06/mso2016060094-abs.
                                        html.

Woods:2017:SAC

[Woo17] Eoin Woods. Should architects code? IEEE Software, 34(5):20–21, September/October 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        https://www.computer.
                                        org/csdl/mags/so/2017/
                                        05/mso2017050020-abs.
                                        html.

Wray:2010:HPP

[Wra10] Stuart Wray. How pair programming really works. IEEE Software, 27(1): 50–55, January/February 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wright:2011:LLA

[Wri11] Arthur Wright. Lessons learned: Architects are facilitators, too! IEEE Software, 28(1):70–72, January/February 2011. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Weyrich:2014:MMC

[WSE14] Michael Weyrich, Jan-Philipp Schmidt, and Christof Ebert. Machineto-machine communication. IEEE Software, 31(4):19–23, July/August 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wu:2016:BPH

[WZX^{+}16] Dongyao Wu, Liming Zhu, Xiwei Xu, Sherif Sakr, Daniel Sun, and Qinghua Lu. Building pipelines for heterogeneous execution environments for big data processing. IEEE Software, 33(2): 60–67, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
02/mso2016020060-abs.
html.

Xiao:2018:HRY

[XYCW18] Lu Xiao, Zhongyuan Yu, Bohong Chen, and Xiao Wang. How robust is your development team? IEEE Software, 35(1):64–71, January/February 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.

org/csdl/mags/so/2018/
01/mso2018010064-abs.
html.

Yang:2018:AAS

[YFMH18] Ye Yang, Davide Falessi, Tim Menzies, and Jairus Hihn. Actionable analytics for software engineering. IEEE Software, 35(1):51–53, January/February 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2018/
01/mso2018010051.html.

Yoshida:2017:KAT

[YLK^{+}17] Hiroaki Yoshida, Guodong Li, Takuki Kamiya, Indradeep Ghosh, Sreeranga Rajan, Susumu Tokumoto, Kazuki Munakata, and Tadahiro Uehara. KLOVER: Automatic test generation for C and C++ programs, using symbolic execution. IEEE Software, 34(5):30–37, September/ October 2017. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
05/mso2017050030-abs.
html.

Zambonelli:2017:KAI

[Zam17] Franco Zambonelli. Key abstractions for IoT-oriented software engineering. IEEE Software, 34(1):38–45, January/February 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2017/
01/mso2017010038-abs.
html.

Zwart:2016:CVU

[ZB16] Simon Portegies Zwart and Jeroen Bedorf. Creating the virtual universe. IEEE Software, 33(5):25–29, September/ October 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
05/mso2016050025-abs.
html.

Zhu:2016:DP

[ZBCS16] Liming Zhu, Len Bass, and George Champlin-Scharff. DevOps and its practices. IEEE Software, 33(3):32– 34, May/June 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
03/mso2016030032.html.

Zdun:2012:GEI

[ZBOC12] Uwe Zdun, Ayse Bener, and Erlinda L. Olalia-Carin. Guest Editors’ introduction: Software engineering for compliance.

IEEE Software, 29(3):24– 27, May/June 2012. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zhang:2019:BRD

[ZCLL19] T. Zhang, J. Chen, X. Luo, and T. Li. Bug reports for desktop software and mobile apps in GitHub: What’s the difference? IEEE Software, 36(1):63–71, January/February 2019. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zdun:2013:SAD

[ZCTZ13] Uwe Zdun, Rafael Capilla, Huy Tran, and Olaf Zimmermann. Sustainable architectural design decisions. IEEE Software, 30(6):46–53, November/ December 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zhu:2013:DDA

[ZF13] Yang-Ming Zhu and David Faller. Defect-density assessment in evolutionary product development: A case study in medical imaging. IEEE Software, 30(4):81–87, July/August 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zhang:2013:SAP

[ZHD^{+}13] Dongmei Zhang, Shi Han, Yingnong Dang, Jian-Guang Lou, Haidong Zhang, and Tao Xie. Software analytics in practice. IEEE Software, 30(5):30– 37, September/October 2013. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zhang:2014:IUC

[ZHG14] Chenlei Zhang, Abram Hindle, and Daniel M. German. The impact of user choice on energy consumption. IEEE Software, 31(3):69–75, May/June 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zimmermann:2011:ADR

[Zim11] Olaf Zimmermann. Architectural decisions as reusable design assets. IEEE Software, 28(1): 64–69, January/February 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zimmermann:2015:ART

[Zim15] Olaf Zimmermann. Architectural refactoring: A task-centric view on software evolution. IEEE Software, 32(2):26–29, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020026-abs.
html.

Zhang:2010:MSQ

[ZK10] Hongyu Zhang and Sunghun Kim. Monitoring software quality evolution for defects. IEEE Software, 27(4):58–64, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zimmermann:2014:PHA

[ZK14] Olaf Zimmermann and Heiko Koziolek. Program highlights and articles from SEI SATURN 2013. IEEE Software, 31(2):87–90, March/April 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zheng:2016:CMD

[ZKB^{+}16] Yudi Zheng, Stephen Kell, Lubomir Bulej, Haiyang Sun, and Walter Binder. Comprehensive multiplatform dynamic program analysis for Java and Android. IEEE Software, 33(4):55–63, July/ August 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
04/mso2016040055-abs.
html.

Zhang:2012:EMA

[ZML^{+}12] Linghao Zhang, Xiaoxing Ma, Jian Lu, Tao Xie, Nikolai Tillmann, and Peli de Halleux. Environmental modeling for automated Cloud application testing. IEEE Software, 29(2):30–35, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zanatta:2016:SCP

[ZMP^{+}16] Alexandre Lazaretti Zanatta, Leticia Santos Machado, Graziela Basilio Pereira, Rafael Prikladnicki, and Erran Carmel. Software crowdsourcing platforms. IEEE Software, 33(6):112–116, November/ December 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

https://www.computer.
org/csdl/mags/so/2016/
06/mso2016060112-abs.
html.

Zhang:2011:AMD

[ZP11] Yuefeng Zhang and Shailesh Patel. Agile model-driven development in practice. IEEE Software, 28(2):84– 91, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zimmermann:2016:DEI

[ZPHW16] Olaf Zimmermann, Cesare Pautasso, Gregor Hohpe, and Bobby Woolf. A decade of enterprise integration patterns: A conversation with the authors. IEEE Software, 33(1):13–19, January/February 2016. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2016/
01/mso2016010013.html.

Zanatta:2017:BFN

[ZSMP17] Alexandre Lazaretti Zanatta, Igor Steinmacher, Leticia Santos Machado, and Rafael Prikladnicki. Barriers faced by newcomers to software-crowdsourcing projects. IEEE Software, 34(2):37–43, March/ April 2017. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL https://

www.computer.org/csdl/
mags/so/2017/02/mso2017020037-
abs.html.

Zhang:2012:SCI

[ZWGY12] Daqing Zhang, Zhu Wang, Bin Guo, and Zhiwen Yu. Social and community intelligence: Technologies and trends. IEEE Software, 29(4):88–92, July/ August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zhang:2016:DIF

[ZXL^{+}16] Weishan Zhang, Liang Xu, Zhongwei Li, Qinghua Lu, and Yan Liu. A deepintelligence framework for online video processing. IEEE Software, 33(2):44– 51, March/April 2016. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://csdl.computer.
org/csdl/mags/so/2016/
02/mso2016020044-abs.
html.

Zhu:2015:ARH

[ZXT^{+}15] Liming Zhu, Donna Xu, An Binh Tran, Xiwei Xu, Len Bass, Ingo Weber, and Srini Dwarakanathan. Achieving reliable highfrequency releases in cloud environments. IEEE Software, 32(2):73–80, March/April 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

http://www.computer.
org/csdl/mags/so/2015/
02/mso2015020073-abs.
html.

Zhang:2015:SBC

[ZZJJ15] Wei Zhang, Haiyan Zhao, Yi Jiang, and Zhi Jin. Stigmergy-based construction of Internetware artifacts. IEEE Software, 32(1):58–66, January/February 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL

                                        http://csdl.computer.
                                        org/csdl/mags/so/2015/
                                        01/mso2015010058-abs.
                                        html.